A Complete Bibliography of ACM Transactions on Algorithms

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 October 2025 Version 1.62

Title word cross-reference

(2 + ϵ) [PS19]. (α, β) [BKMP10]. (Δ + 1) [BGK^{+}22, HP22]. (h, k) [BEJK19]. (k, l) [BDR23]. (k, r) [DFHT05]. (min, +) [CMWW19]. (n − 1) [RW10a]. (s, t) [BBP23]. 1 [KN16]. 1 − 1/e [HTWZ19]. 1.5 [KN16]. 1/3 [DFM23]. 2 [ERV16, Fuj12, GS17, GILP16, HCT^{+}11, HVV19, KN16, LMM^{+}21, SS18]. 2 + ϵ [AGLW18]. 2.5545 [HCT^{+}11]. 3 [CLL^{+}12, FLL^{+}19, Heg06]. 4/3 [HVV19]. 4k^{2} [Tho10a, Tho10b]. β [ADGH21]. ℓ_{1} [WW22]. ℓ_{p} [Wei22]. F [ALM^{+}20, Gab25]. H [BCSV20, VWY10]. K [DM09, BBB20, ABF^{+}18, AMS06, BHW20, BEKN23, BPR^{+}17, CGK20, Cha10b, CV20, CMVZ16, CD24, DKR16, DKN17, FHR07, FN10, FPZ23, FGK^{+}24c, GIN^{+}17, GHNR10a, GHNR10b, GWZ21, HHL^{+}16, HMS07, JN24, Laf23, Lev09, Li17, LN22, LBSZ21, PT16, RT22, RRS07, RZ12, SST22, WZ16]. k/r [GWZ21]. L [GGI^{+}21]. l_{1} [Fil25, BDW19]. M [HIMŽ19]. n [RW10a]. n^{0.4} [BCAD23]. n^{4/3} [Bon22]. O(1) [BGK^{+}22]. O(log, k) [FM25]. O(log k) [ACER19]. O(log n) [KK13]. o(mn) [Cha12]. O(n^{2}) [BS06a]. O(n^{2.75}) [AFM08]. O(n^{2/3}m) [Gab25]. O(n^{3}) [GT08]. O(n log^{2} n) [KMW10].

O(n log n) [BKM09b, BKM09a]. O(nm) [CMA^{+}19, KMMP07]. O(\sqrt{log n}) [Kol25, PS23]. O(V E) [DC05]. O^{⋆}(2.7k) [LN22]. P_{3} [LPS24]. P_{5} [ALL^{+}25]. P_{6} [GKPP22, LPV18]. r [GWZ21]. s − t [CLS^{+}22]. st [BSWN15, KW16b]. t [DP06].

-apices [SST22]. -Approximate [DFM23]. -Approximating [LMM^{+}21]. -Approximation [BCAD23, HVV19, KN16, PS19, Fuj12, FM25, KK13, Kol25]. -ary [DP06, RRS07]. -Center [CGK20, BHW20, DFHT05]. -clustering [Lev09]. -Coloring [BGK^{+}22, HP22]. -Colorings [BCSV20]. -Competitive [PS23, ACER19]. -Convex [HIMŽ19]. -Convolution [CMWW19, Cha20]. -CSP [GS17]. -Cut [BSWN15, FPZ23, KW16b]. -cuts [BBP23]. -d [DM09]. -deletion [ALM^{+}20]. -Dyck [FGK^{+}24c]. -Edge [GILP16, HVV19]. -Facility [ABF^{+}18]. -Forest [DKN17, GHNR10a, GHNR10b]. -free [ALL^{+}25, GKPP22, LPV18]. -Heavy [BDW19]. -Hitting [FLL^{+}19]. -Independence [PT16]. -Leaf [Laf23]. -Matching [Gab25]. -Means [BBB20]. -Median [BDR23, BPR^{+}17, HHL^{+}16, Li17]. -mismatch [JN24]. -Opt [ERV16]. -packings [GGI^{+}21, LPS24]. -Path [GWZ21]. -Paths [LBSZ21]. -permutations [RW10a]. -Plurality [ADGH21]. -restrictions [AMS06]. -Route [CMVZ16]. -Server [BEJK19, BEKN23, CV20, FN10]. -Set [FLL^{+}19, Cha10b, RW10a]. -Shortcut [CD24]. -shredders [Heg06]. -Simple [GWZ21]. -Spanner [DKR16]. -spanners [BKMP10]. -subgraphs [VWY10]. -Submodular [WZ16]. -time [KMW10]. -traveling [FHR07]. -Walks [SS18].

0.8776-Approximation [ABG16].

1-center [YLW08]. 1-median [YLW08].

2-gathering [SZ10]. 2002 [Epp07]. 2007 [Gab09]. 2008 [HT10]. 2016 [BGN^{+}18]. 2017 [MWY19]. 2019 [HK22]. 2020 [KADK22]. 2021 [ELMS25]. 2022 [DMT24]. 2023 [BKN^{+}25]. 2D [KLM^{+}25].

3-approximation [BPGN09]. 3SUM [Cha20]. 3SUM-hard [Cha20].

4-leaf [BLS08].

above [LPS24]. Access [ARS^{+}14, BJKK18, CKR12, ELMR21]. accesses [CFLM07]. Achieving [APF^{+}10]. ACM [BKN^{+}25, BGN^{+}18, DMT24, ELMS25, HK22, KADK22]. ACM-SIAM [BGN^{+}18, DMT24, HK22, KADK22]. Activation [Fuk17]. Acyclic [HHM^{+}18]. Ad [KMPS16]. Adaptive [BPL18, CW16, GM12, Jug24, MPV10]. Addendum [GT16a]. Adder [BH19, HS18]. Adding [GW22]. Additive [BKMP10, BW21, PZ24, RS11a].

Address [BBC^{+}25]. Address-Translation [BBC^{+}25]. Adjustable [DEK21]. Adjusting [Elm17, ST25]. Admission [AAG09]. Admit [LT24]. Admits [PS23]. Adversarial [CKR12, NE19, CDHW09]. adversary [AC10]. Advertising [AFH^{+}16]. Advice [BFKR21, BFKR25, GP19]. Affine [VZ21, GMT11]. Agents [DP14]. aggregation [BMSV^{+}09]. Agnostic [Wim16]. agreeable [JLSS12]. agreement [KKK^{+}10]. Algebraic [AAE^{+}25, AK18, CLL14, Vig14]. Algebras [KW16a]. Algorithm [ACER19, AFS18, AKS17, AKLR20, AER15, BKMV20, BDW19, BLS23, BFPP18, BKN21, CMA^{+}19, CLL^{+}12, CMV16, CJL17, CI17, CGMY22, DFM23, DKN17, DS19, ERV16, FPS22, FGK^{+}16, FGK^{+}24c, GKPP22, HHL^{+}16, Hir19, HU20, HVV19, Jug24, KKK18, KK13, KK16, KKR^{+}20, Kol25, KN16, MMS14, PRS20, PS23, AFM08, And10, AMM07, AR06, BB09, Blä08, BS10, BD11, CCKR11, DC05, CCM10, CKS05, Cla10, CNP^{+}11, DMRW09, DS08, Dji10, DV10, EK07, EPR10, EFKN09, Fuj12, HMS07, IM12, JZ06, KY13, Kau07, KMW10, RW09, VH05]. Algorithmic [AMS06, BCFN07, FMM^{+}25, GIKW19, HS17, PSZ24]. Algorithms [ASW08, ALM^{+}20, ABF^{+}18, AMW20, AMNS17, AKS21, ACE^{+}23, AFK^{+}18, ADKS25, BGGN16, BBHT17, BKN14, BEKN23, BKN^{+}25, BODD^{+}20, BGH19, BR14, BGN^{+}18, BBM19, BF24, BBT12, BCMSM12, Bre23, BSSX20, BF18, CHA18, Cab19, CFRY22, CLL14, CV20, CFM21, CMVZ16, CDHW09, CMY11, CDP19, CJKV24, DMT24, DGM18, DHK16, DH18, DPS18, EFM^{+}16, EW20, EN19, ELMS25, EHL^{+}18, ELR^{+}08, FLN14, FL25, FLK^{+}20, FLPR12, GMP23a, GMP23b, GS17, GKM16, GHH22, GLLZ21, HH17, HKKK16, HT21, Har18, Har21, HL13, HK22, ISG07, JPS22, JCCW22, KADK22, KLM^{+}25, KLP^{+}16a, KBNvL20, KX19, KLP16b, KMPS16, Lac13, LNR^{+}14, LMS18, LRS18, LPS^{+}20, MV15, MP22, NS16, RS17, SST22, SHHA16, Swa16, Ter25, AAK06, AMR09, AF07, AA14, AR09, AGvS13, AKL10, AKR12, BCD12, BKS12]. algorithms [BAT11, BFK^{+}12, BCM11, BF07, CPR^{+}11, CMO^{+}08, CMM09, CKP12, CJST07, DFHT05, DI06, DJP^{+}12, EF12, Elk11, Epp06, FP10, FZ07, GS09, GKLT09, GKK^{+}09, GHPT05, HSS07, Iba08, IMY10, JR05, KNS^{+}07, MV08, MZ12, PU07, PR08, RSS06, SZ10, YLW08, CEGK11]. All-Or-Nothing [AFH^{+}16]. All-Pairs [KT18, Cha12, RS11a, MTZ10]. Allocation [AKS17, PS16, CCKR11, GN14]. Allocations [AMNS17, LS24]. Almost [AL13, BLL^{+}24, BDLP23, DJO^{+}25, DH18, CPL12, Elk05, FKW11]. Almost-Optimal [BDLP23]. Alphabet [BN14]. Alphabet-Independent [BN14]. Alternating [NRS18]. Alternation [BK08]. Alternative [Jug24]. amid [AFS18]. amnesic [GLPP08]. among [CW15, FKS08]. Amortized [GHT18]. Analyses [BF24]. Analysis [BBHT17, BCKM20, BKK^{+}19, ERV16, ER17, FN20, GHPT05, SCRS17, WNN15, dBBJW21, AAY10, AR09, BK08, BAT11, DMM^{+}12, DI06, DK12, EP05, Epp06, FBV09, GN14, GR10]. analytic [SSS^{+}11]. Analyzing [CCW18]. Anarchy [GHLL16, DHMZ12]. ancestor [GRR06]. Ancestors [Gab17]. annotated [GGN06]. Annotations [CCMT14]. anonymity [APF^{+}10]. Anonymous [DP14, GMP17, MPY22]. Answers [FGRZ25]. any [FKW11]. apices [SST22]. Application [AFH^{+}16, ARS^{+}14, Coh18]. Applications [AZBG^{+}22, ACE^{+}20, BR16, BFR23, CCLS25, CdM22, DEK21, DHK16, EK20, Fil24, GIKW19, Hir19, HJT17, KMNS17, KW16a, LPS^{+}20, PSZ24, Pre21, Swa16, AAY10, AG10, AZ08, BB08, DMM^{+}12, FGPS08, FGGV06, FSP08, NW07, RRS07, VWY10]. applied [BM08]. Approach [BFGT16, BFH21, BFV22, Gab16, LMMW16, AAA^{+}06, NW07, VB08]. Approximability [BG20, CGNS08]. Approximate [AEP18, AC10, AFK^{+}15, BS06a, BCHR20, BSSX20, CW16, DFM23, FJS14, GMP05, GLNS08, HLS07, LBSZ21, PP18, WY18, Wei22, AKR12, BFG09, CSTW12, Vis08]. Approximately [FGZ21, FGRZ25]. Approximating [AAdFM22, BDR23, CKM^{+}24, CLNV14, CDKL20, Das13, DKR16, FR10, GGI^{+}21, GKK23, GJL12, GGG10, GJNW23, HLS09, KR16, LMM^{+}21, LMMW16, MR09, Mar10, Nut09, Nut12, Oum08, RT13, WY16, Man12]. Approximation [Adj19, ALM^{+}20, ABF^{+}18, AMNS17, AKS21, AGLW18, ABG16, BR14, BHPR19, BKM15, BCAD23, BPR^{+}17, CFM21, CMVZ16, DHK16, DKN17, EFM^{+}16, FGL^{+}20, FKRS19, FM25, GKK^{+}09, HKKK16, HHL^{+}16, HW19, HL13, HVV19, IMY10, JMR19, JS23, JR05, KK13, KK16, KLM^{+}25, Kol25, KN16, KW16b, KNS^{+}07, MMS14, MV08, NHK08, PS19, Swa16, TY18, BFKS14, BKLP07, BPGN09, Blä08, BKM09a, BKM09b, CCKR11, CGR08, Cla10, DJP^{+}12, DV10, EFKN09, Fuj12, HIMY07, JZ06, Joh06, Kar08, Kar09, LDX09, RS09, SS08a, VH05, CPR^{+}11]. Approximations [ASS19, FIM^{+}06, Jac11, LM11]. APSP [AGW23, BGMW20, CW21]. arbitrarily [And09]. Arbitrary [BCP13, EP12]. Arborescence [FLK^{+}20, DV10]. Arborescences [KMSY25]. Arboricity [ELR20]. arc [GGN06]. arc-annotated [GGN06]. Archipelagos [GRS17]. architectures [NW07]. Arrangement [TY18, NS10]. Array [OOB24]. arrays [FGGV06, LK08]. Arrival [BH19, LS20]. Arrivals [HTWZ19]. Art [BM20]. ary [DP06, RRS07]. Ascending [BGH19]. Ascending-Price [BGH19]. Aspects [HS17, BCFN07]. Assembly [AAGH25, CRTZ24]. Assignment [AFH^{+}16, BH12, MMS14, MS17, Soc16, CNP^{+}11, HLS09, Jac11, LM11]. Asymmetric [AKS21, Blä08]. Asymptotic [HJT17, SS08a]. Asymptotically [FV19, GIN^{+}17, HS18, LZ25]. asynchronous [KKK^{+}10, KS08]. asynchronously [CPL12]. Atomic [FKS08, CKK10a]. Attachment [ELMR21]. auctions [BLW09]. Augmentation [Adj19, MV15, KKPW24, Nut09, VB08]. Augmenting [KN16, EFKN09]. Automata [CP25, AKL10]. automaton [CFI^{+}08]. Automorphisms [iKMNZ25]. Average [AR08a, AR09, BF24, RKH20, WNN15, IM12]. Average-case [AR08a, AR09, BF24]. axis [CKS09]. axis-parallel [CKS09].

B [FV16]. B-Tree [FV16]. back [LY08]. Backbone [HT21]. backtracking [Epp06]. Balance [ABG16]. Balanced [AG10, HST15, LS24, LY08, SYZ20]. Balancing [BHS14, BKK^{+}19, EDKM07, GMP05]. Balls [DT16, AHPSW10].

Bandwidth [CP12, BCFN07, EK06]. bandwidths [PRV11]. Barrier [AEMV24, BGK^{+}25, GS18]. Based [BGK^{+}25, Wal25, ABHS22, AMM07, BBM19, Cha10a, GN14, HMS12, MMS14, PU07]. bases [MM09]. Basic [DKR16, VZ21]. basis [NS09]. Batch [LS20]. Batch-arrival [LS20]. Batched [AK18, AAY10]. batching [BNGK^{+}09]. Bayesian [GM12]. be [BGMW20]. Beating [Adj19, HTWZ19]. Beats [BBM19]. Behaved [FV19]. Being [ABG16, FLM^{+}12]. Bend [BRW16]. Berlekamp [KY13]. best [FFM12, PUW08]. best-response [FFM12]. Better [ABG16, BKN14, BT24, BW21, FG08, Kar09]. between [AGW23, CCDL16, MA16]. Beyond [ABZ19, BGGS24, EHL^{+}18, FGG^{+}25, FGK^{+}24b, Har16, Li17, WZ16, ASS19, FGYZ22, PRS12]. BFS [PP16, PP18]. Bi [KR19]. Bi-Monotonicity [KR19]. Bias [Kav24]. Biased [ABG16, HSS22]. Bichromatic [ABD^{+}25, AAK06, Cha10b]. Biclustering [HL13]. biconnected [PS10]. Bicriteria [ABHS22, RS09, FN10]. bidders [BCC^{+}10]. Bidirected [CFM21]. bifamilies [Nut12]. Bimatrix [DFM23]. Bin [ADKS25, EPR13, HCT^{+}11, BNLT07, LK08]. Binary [FGL^{+}20, HS18, HIMŽ19, STK16, BHMS11]. Bipartite [AKLR20, HTWZ19, Lau06, LM19, GKK10]. bipartiteness [Epp09c]. Bipartization [BLL^{+}24]. Bisection [ABG16, SS09]. Bit [BH19]. bits [PR08]. Bivariate [MA16]. Black [AHRT05]. block [FZ07]. Blocks [Wal23]. Bloom [NE19]. Boolean [DHK16, FGYZ22, KKR19, Mas21]. Bootstrapping [FCFM09]. Bottleneck [AKS21, DS08]. Bound [AAdFM22, CRSS22, Kra14, Soc16, HCT^{+}11]. Bounded [ACGP16, Adj19, BEJK19, BGGS24, Cab22, CDJS17, CDP19, ELR20, EFF^{+}15, FL25, FMR24, FMM^{+}25, GSV20, GNSW20, JS23, LMS18, PPSV18, SHHA16, BEH^{+}10, BHKK12, Dji10, KK06, MR09, MZ12, OGGW10, PRS12, Svi10]. Bounded-degree [BGGS24, MR09]. Bounded-Genus [PPSV18]. bounded-length [KK06]. Bounded-tree-width [GNSW20]. Bounded-Treewidth [FMM^{+}25]. Bounding [FGPS08]. Bounds [ACHKP21, ABHS22, AFCK25, BODD^{+}20, BN15, BHD^{+}21, BCK^{+}20, CHGG^{+}17, CC25, CDKL20, CGH17, DI25, DLS14, FMR24, FMM^{+}25, HH17, Har19, JW13, KT18, SCRS17, WW22, WY18, AR08a, AGvS13, BD11, CGK^{+}11, Cha10a, CV07, FGPS08, GHKS08, MSS11, PU07]. box [AHRT05]. branch [GT08]. branch-decomposition [GT08]. Branching [FPZ23]. Branchings [HHM^{+}18]. Breadth [HKN17]. Breadth-First [HKN17]. Breaking [AEMV24, GS18]. Broadcast [BKN14, CEGK11, EK07, IM12]. Brownian [AZBG^{+}22]. Bruijn [SSA25]. Budget [EFF^{+}15, RS09]. Budgeted [BPR^{+}17]. Buffer [AER15, CR18, AKM08, CJST07, GS09, RS11b]. buffering [LK08]. Building [RT22]. Burden [EMS23]. buy [BKLP07]. Bypassing [GRSW16]. Byzantine [KKK^{+}10].

Cache [CR18, FV16, FLPR12]. Cache-Efficient [CR18]. Cache-Oblivious [CR18, FV16, FLPR12]. Caching [ACER19]. Cake [CF20, EP11, SHHA16].

Campaigns [AFH^{+}16]. Can [BGMW20, BCM11]. Cannot [BGMW20]. Cap [Coh18]. Capacitated [JS23, Li17, MZ23, BH12, ELR^{+}08, HKRL07, Hir19, JR05]. Capacities [Wan22, MV08]. capacity [AGG10]. Cardinality [Gab25]. carpool [CNP^{+}11]. Carry [BH19, IP11]. Cascading [CZ24]. Case [CHL^{+}20, FGL^{+}19, WNN15, AR08a, AR09, ADHY08, BF24, CFK^{+}07, CV07]. Center [CN19, CGK20, HPJ21, HPST19, ASW08, BHW20, DFHT05, YLW08]. Center-Type [HPST19]. Centrality [AGW23]. centralized [Elk11]. Certification [GT16a, GT16b]. chairman [CNP^{+}11]. changes [CFH07]. Channel [BJKK18, CKR12, Soc16]. channels [AZ07]. Checkpointing [BODD^{+}20]. Cheeger [KW16b]. Cheeger-Type [KW16b]. Chordal [ALM^{+}19, HSS07, Iba08]. CIOQ [AR06]. Circuits [BH19, FGLS19, HS18]. circular [NS10]. Class [HIMŽ19]. Classes [CDP19, KRS19, SST22, FP10, GKLT09]. Classic [dBBJW21]. Classical [BBHT17]. Classification [CY25, GLS10]. Claus [AFS12]. Clique [CDP19, FGL^{+}19, GHK^{+}24, HKP^{+}18, Oum08]. Clique-Width [GHK^{+}24, CDP19, FGL^{+}19, Oum08]. cliques [II09]. Clocks [ANFS17]. Clones [BG20]. Close [Wal25]. Closed [CMA^{+}19, CEV21, SST22]. closest [AKS08]. Closure [Epp18, Rod08]. Cluster [LPS24]. Clustering [ABS10, BHW20, BDR23, FKRS19, GMP23b, GLS10, RKH20, ZO08, APF^{+}10, CGK^{+}11, Epp09b, EV10, Lev09]. CNF [CDL^{+}16]. CNF-SAT [CDL^{+}16]. Co [Kra14]. Co-Nondeterminism [Kra14]. Coalescence [KMTS23]. Coalescing [KMTS23]. coalitions [FKS08]. Code [LZ25, KL06]. Codes [LZ25, MN18]. Coding [CLL^{+}12]. Coin [BBHT17]. Collapsing [KR24]. Collecting [CJJ20, Fuk17, HKKN12]. Collections [Nav25, CHLS07]. collective [GLPP08]. Color [CLL^{+}12]. Colored [EK20, FGK^{+}24b]. Coloring [BGK^{+}22, CHL^{+}20, FGL^{+}19, HP22, SW20b, BNCS08, CKS09, DKT11, HKS11, PRV11, SS08a]. Colorings [BCSV20]. Colors [DLS14]. Colouring [GHH22, ND25]. Column [BSSX20, Joh05, Joh06, Joh07, Khu05, Khu06, Khu07]. Column-sparse [BSSX20]. combination [GHPT05]. Combinatiorial [EF12]. Combinatorial [AKS17, AAdFM22, BST08, CMM25, CKM^{+}24, FGPS08, GM12, BCN12, FP10]. combinatorics [SSS^{+}11]. Combined [VZ21]. Commitment [EMS23]. Common [AW19, BCAD23, CMA^{+}19, Gab17, JN24, CLLJ08, CKS05]. Communication [KLPP23, KR16]. community [ZO08]. Compact [AGM^{+}08, BB08, KRX16, CSTW12]. Compacting [CKL^{+}09]. Comparison [Cha10a]. Comparison-based [Cha10a]. Comparisons [AFHN16, CGMY22, CY25]. Competitive [ACER19, AKM08, AER15, BEKN23, BAT11, BCI^{+}23, GR10, LMMW16, MRYY24, PS23]. Competitiveness [AEMV24]. Complete [CRTZ24]. Completeness [BG21, GIKW19, Joh05, Joh06, Joh07]. Completion [BFPP18, BCK^{+}20, GLP06]. Complexes [BS18]. Complexity [AAdFM22, BCC21, BCSV25, BCK^{+}20, BBB20, BFKR25, BK24, CKP^{+}19, CBFWW15, CDL^{+}25, CF20, Das13, DHM^{+}14a, DHPR16, EMTG23, FGZ21, FMR24, FMM^{+}25, GHK^{+}24, GJLS17, KKR19, KR24, LM19, Mas21, NZ24, PSZ24, dBBJW21, AK12a, BCMSM12, CEGK11, FLM^{+}12]. Component [HKM^{+}12]. Components [Lac13]. Composition [Neu22]. Compositions [Kra14]. Compressed [BN14, BGS23, BFV22, CHLS07, CEK^{+}21, FMMN07, FV16, Gaw13, Jez15, BFG09, FV10, MN08, RNO11]. compressing [FGGV06]. Compression [GMS19, KW14, FGGV06]. Computation [BPL18, BH19, CR18, CDP21, KKR^{+}20, MN18, FIM^{+}06, PT11]. Computations [FLS^{+}18, FMS^{+}10]. Computed [BGMW20]. Computing [AFN^{+}18, AFS18, AMNS17, BF09, BCKV06, BLS23, Cab22, CW15, CMV16, CD24, DGM18, Elk05, FGK^{+}16, MA16, Nav25, NZC11, RW10b, SS18, BB09, Dji10, EPR10]. Concentration [HH17, Har19]. Conditional [BCSV25, BKN21, DI25, KT18]. conditions [ST08]. configuration [HKS11]. confinement [AKPS10]. Conflict [ND25, BNCS08, CKS09]. Conflict-Free [ND25, BNCS08, CKS09]. conflicting [GHKS08]. Confusability [CCKN19]. Congestion [KK16, CKK10a, FFM12, FKS08, Swa12]. congruent [CKS09]. Conjunctive [FGRZ25, CFK^{+}07]. Connected [Bre23, FLST18, HVV19, Lac13, SYZ20]. Connectivity [AAGH25, BK16, CHGG^{+}17, CC25, CNP^{+}22, Gab16, GILP16, IIK^{+}23, KN16, MV15, CEGS11, CRV11, EFKN09, Kor10, Nut09, Nut12, VB08]. Conquer [GS17, HJT17, FGPS08]. Consensus [BBT12, CN14]. consequence [BGLZ09]. Conservative [BG20]. Consistent [DJO^{+}25]. Constant [AW19, ATG^{+}14, HHL^{+}16, HP22, Jac11, Laf23, LZ25, AHRT05, VP07]. Constant-time [HP22, AHRT05]. constants [IP11]. Constrained [ADD^{+}18, BR16, BFR23, CS11, EHR16, Mas21, BMSV^{+}09, DKK06]. Constraint [AZBG^{+}22, GRS17, GM14, CMM09, SS09]. Constraints [AAGH25, MS17, DH12, GLP06, JZ06]. Constructing [EN19, GW07, Elk11]. Construction [OOB24, AMS06]. Constructions [ES16, SSA25, SCRS17]. Constructive [CHL^{+}20]. contact [BGPV08]. content [BH12]. Continuous [CCW18, NW07]. continuous-discrete [NW07]. contractible [Cab10]. Contraction [BLL^{+}24, PSZ24]. Contractions [CKL^{+}21b, FPZ23]. control [AAG09, Hal12]. Convergence [CN14, EDKM07, GN14, FFM12]. Convex [BRW16, CDD^{+}15, DEK21, HIMŽ19, MW24, BB09, CAY10, GJL12, MSS11]. Convolution [CMWW19, Cha20]. convolutions [BF09]. Coordinate [BCSV25]. Coresets [Cla10]. Correction [BPL18, AEL^{+}12]. Correlation [BPR^{+}17]. Correlations [KKK18]. corridors [GGG10]. Corrigendum [GHKS13]. Cost [BFCF^{+}17, HJ15, MV15, AZ08, AKLR20, BKLP07, EKS05, Fuj12, IMM08, Lev09, Nut12, RS09]. Cost-Oblivious [BFCF^{+}17]. cost-sharing [IMM08]. cost-tree [Fuj12]. Costs [Adj19, BRW16, FGK^{+}24a, ELR^{+}08, ST10]. Counting [AFK^{+}15, BCN12, BKK17, CW16, FGZ21, FMR24, FGRZ25, FMM^{+}25, GRS17, GS17, GLSY23, LBSZ21, WY18, AC10, BCEG07]. Coupling [Wal25]. Cover [ACS^{+}22, DHK16, DKR16, ER16, FGG^{+}25, Fil25, INV16, KPR16, MMS14, AAG09, Fuj12, GKP08, HL06a, Kar09, Vis08]. covered [DC05]. Covering [GSV20, LPS^{+}20, SP24, CMA^{+}19, EKS05, HMS12]. CoveringLSH [Pag18]. Covers [Fuk17, GM14]. Creation [EFF^{+}15, GHLL16, DHMZ12]. Criteria [ADV^{+}16, GHKS08]. cross [IMM08]. cross-monotonic [IMM08]. cryptography [GHPT05]. CSP [GS17]. CSPs [CRZ20, KKR19, MWsZ23, PZ24]. Cubic [BLS23]. Cuckoo [Wal23, DK12]. Curse [CJ18]. Curved [CW15]. Curves [BDR23, BFL^{+}23, CdM22, IIK^{+}25, EP12]. Cut [BCC21, BSWN15, CKM^{+}24, CMVZ16, ER17, FPZ23, GHT18, Kol25, KW16b, Wah22, CKL^{+}09, CGR08]. Cuts [ACHKP21, BEH^{+}10, BBP23, CKL^{+}09, PT11]. Cutting [CF20, SHHA16, EP11]. Cycle [ALL^{+}25, AEL^{+}12, BFGT16, CRSS22, FGL^{+}19, KW14, CDEM10, GN08, KNS^{+}07, LDX09, MM09, PT11, RW10a, HKM^{+}12]. Cycle-Finding [CRSS22]. Cycles [FGK^{+}24a, GLSY23, Cab10]. cyclic [BS06b].

d [DM09, CLL^{+}12, HCT^{+}11]. D-Matching [CLL^{+}12]. DAGs [GR10, HH17, MTK^{+}19]. Data [BFV22, CCMT14, FF23, Gab17, GKS^{+}11, GIN^{+}17, HMZ16, MRYY24, SCRS17, Wal25, AKS08, ACY12, BCEG07, BB08, CFLM07, DIL07, GHKS06, GHKS13, GLS10, GKK^{+}09]. Data-Structure [MRYY24]. databases [CFK^{+}07]. dates [KKW12]. deadline [CCL^{+}09, GLP06]. deadlines [JLSS12]. Deamortization [BFH21]. Deciding [CCKN19, Kau07]. Decision [CPR^{+}11, CZ24, MTZ10]. Decisions [KR18]. Decoding [LZ25, AGKS07]. Decomposition [BLL^{+}24, PSZ24, AZ08, DMRW09, Epp09b, GT08]. Decompositions [CKS19, CKL^{+}21b, KLP^{+}16a, DC05]. Decremental [Lac13]. Degeneracy [CGH17, PRS12]. Degenerate [GLSY23]. Degree [HKP^{+}18, BGGS24, BHKK12, MR09]. Degree-4 [HKP^{+}18]. Delaunay [OGGW10]. Delay [AGGP21, EK06]. Delays [FP13]. Deletion [ALM^{+}19, ALM^{+}23, CM15, STK16, ALM^{+}20]. Deletions [ATG^{+}14]. Delta [KKR19]. Delta-Matroids [KKR19]. demand [CMS07, EK06]. Dense [KRS19, ST08, And10]. density [KP08, ST08]. Dependent [HSS22, KR18, CRR09, JMR19]. Depth [BEJK19, HS18]. Derandomized [Har19]. Derandomizing [CW21, CDP21]. Derivative [CDJS17]. Descent [Hir19, AKR12]. Design [BFV22, BK16, CLS^{+}22, CEV21, AZ08, CKK10b, CGNS08, EKS05, JR05, KSS09]. Despite [DPP14, AC10]. Detecting [LN22]. Detection [BFGT16, AEL^{+}12, HKM^{+}12]. Determining [AR08b]. Deterministic [BCEG07, BNCS08, BKK^{+}19, BDLP23, BF18, CHA18, CW21, CDJ23, CI17, DP14, FTK20, GP19, Har18, Lac13, LMPS18, LBSZ21, MPY22, NS16, SP24, TSZ14, MTZ10, Ruž08, Ruž09]. deterministically [KS08]. DFTs [HU20]. Diagrams [AK18, DHPR16, MP22]. Dial [GNR15, GHNR10a, GHNR10b]. Dial-a-Ride [GNR15]. Diameter [AGW23, Bon22, Cab19, CKM^{+}24, Sol13, WY16]. dictionaries [BB08, FGI09, RRS07, Ruž08]. Dictionary [CEK^{+}21].

Dictionary-Compressed [CEK^{+}21]. differ [EK06]. different [Jan05]. difficulty [HSB07]. Digraphs [CCLS25, CRSS22, HHM^{+}18, CRV11]. Dilation [GW22]. Dilworth [HHM^{+}18]. Dimension [ACGP16, JS23, KRX16, CGK^{+}11]. Dimensional [AEP18, BCSV25, CHLT14, HW19, MW24, WY18, CZ24, GLS10, NS09]. Dimensionality [CJ18]. Dimensions [AAE^{+}25, CHA18, FR10]. Directed [CCHM15, CDKL20, FM25, GILP16, LL25, BD11, CEGS11, EPR10, GN08, HKRL07, KMW10, NS10, RTZ08, RZ12, VB08]. Direction [ADD^{+}18]. Direction-Constrained [ADD^{+}18]. Discontinuity [CCW18]. Discounted [MTZ10]. Discovering [FKW11, GRS17]. Discrepancy [EPR13]. Discrete [AFK^{+}15, BKN^{+}25, BGN^{+}18, BKN21, DMT24, ELMS25, HK22, KADK22, NW07]. Discretization [KLPP23]. Discs [ACHM22]. Disequalities [FGRZ25]. Disjoint [AW19, IIK^{+}25, KK13, KK16, CK07, CS07, DS11, GW07, LPS24, ZO08]. Disk [BLL^{+}24, PSZ24]. disks [CKS09, GKK^{+}09]. displacements [Jan05, Vio05]. Disposal [CHJ^{+}18, SHHA16]. Dissections [FSP08]. Distance [ACGP16, AFN^{+}18, AFK^{+}15, BPL18, BW21, BGMW20, BKN21, BFL^{+}23, CMT22, CD24, EP16, FF23, FGK^{+}24c, GS18, GW20, GSW24, MA16, WY13, ABS10, AHPSW10, AK12a, BS06a, CSTW12, CW10, CM07, DMRW09, FR10, GLNS08, HPR14, MR09]. Distances [Cab19, GMV09]. Distinct [Bła20, KKW12]. distortion [Pet09]. Distributed [ASS19, AKR12, AKPS10, BHS14, BJKK18, CHL^{+}20, FGG^{+}25, GKP08, HT21, JW23, KMPS16, PRS20, SCRS17, CMY11]. distributing [FMS^{+}10]. Distribution [CL22, HS17, LCS^{+}19, BH12, CRV11, Vio05]. Distribution-free [LCS^{+}19]. Distribution-Sensitive [CL22]. Distributional [WNN15, CDHW09]. Distributions [CDJS17, HK25]. Divergent [GT16a, GT16b]. Diversification [BJLY17]. Divide [HJT17]. Divide-and-Conquer [HJT17]. Dividing [HJT17]. DNA [KSS09]. Do [SZ20]. Document [SSTV23]. Domain [FGYZ22, Mas21]. Domains [VZ21, Wim16, OGGW10]. dominance [BST08, Epp09a]. Dominating [ASS19, BDH^{+}20, FMM^{+}25, FLST18, GS17, FGPS08, GKLT09, PRS12]. Domination [HMVW19, LPV18]. Dominator [GT16a, GT16b]. Dominos [AARA23]. Doors [KR18]. Dotted [ALM^{+}12]. Doubling [ACGP16, CGMZ16, CJ18, CJJ20, JS23, KRX16]. Doubly [SYZ20]. Drawing [BRW16]. drawings [BLPS13]. Drift [LS24]. Drop [AEMV24]. Dual [AD16, BBP23, DH18, Hir19, WNN15, BCM11, VB08]. Dual-edge [BBP23]. Dual-Pivot [AD16, WNN15]. Dualities [CCLS25]. due [KKW12]. Dyck [FGK^{+}24c]. Dynamic [ACS^{+}22, AMW20, ALLS07, ANFS17, BCC^{+}10, BFH21, BGK^{+}22, BBM19, BJLY17, CN14, CL22, CKS19, DI25, DFL^{+}25, GS18, HKN17, HP22, KP08, MN08, MTK^{+}19, NS14, NS16, OSSW20, RST14, SW20b, Tao14, AKS08, AHTL05, BKS12, CHLS07, DI06, Elk11, Epp09a, GK09, Iba08, LK08, Rod08]. Dynamics [FV19, FFM12]. Dynamization [MRYY24]. ear [DC05]. Easy [KPR16]. Edge [BHPR^{+}20, BK16, CMA^{+}19, CHL^{+}20, CK07, GILP16, GM14, HVV19, IIK^{+}23, KK13, KK16, KN16, MV15, RKH20, BBP23, EFKN09, HKS11, MZ12, SS08a]. Edge-Connectivity [IIK^{+}23, KN16, MV15, EFKN09]. Edge-covering [CMA^{+}19]. Edge-Disjoint [KK13, KK16, CK07]. Edges [ADD^{+}18, GW22]. Edit [BGMW20, FGK^{+}24c, GS18, AK12a, CM07, DMRW09]. Editing [LPS24]. Editor [Gab05]. Editorial [Alb10a, Alb10b, APP17, BGN^{+}18, Coh21, RRSW16, Sri19, Buc08]. Effective [CLS^{+}22]. effectiveness [Swa12]. Efficiency [ST25]. Efficient [AAK06, AFS18, BFF^{+}22, BS18, Bre23, CFRY22, CR18, DH18, EN19, FGK^{+}16, LZ25, LPV18, MN18, SSA25, SZ10, AAY10, AF07, ADHY08, MZ12]. efficiently [AGKS07, BKLP07, CAY10, Heg06]. Eigenvalues [TY18]. Election [ARS^{+}14, CKP^{+}19, CDJ23, GMP17, MPY22, KKK^{+}10]. element [CS07]. element-disjoint [CS07]. Elements [Bła20]. Elimination [YB12, CMM25]. Embedded [ADF^{+}15, Cab10]. Embedding [ADD^{+}18, BR16, LDX09]. Embeddings [AFT19, AEP18, CGR08, Fil25, WW22, ABD^{+}08, CMO^{+}08, IN07]. Emulators [EN19]. encoding [FSP08, RRS07]. Encodings [GIN^{+}17]. End [KMPS16]. End-to-End [KMPS16]. Energy [AF07, CKP^{+}19, CDJ23, DH18, BCD12, CCL^{+}09]. Energy-Efficient [DH18, AF07]. entity [CPR^{+}11]. Entropy [CDI^{+}12, GMS19, PU07, AMM07, CCM10, GMV09, MN08]. Entropy-based [PU07, AMM07]. entropy-compressed [MN08]. Enumerating [BDH^{+}20, CMA^{+}19]. Enumeration [II09]. Enumerative [HS17]. Environments [NE19]. Envy [SHHA16]. Envy-Free [SHHA16]. equal [GP08, IMY10]. equal-length [GP08]. Equality [Pre21]. equals [FGGV06]. Equations [DJO^{+}25, Epp06]. Equilibria [DFM23, FGK^{+}16, BCKV06, CV07]. Equilibrium [Das13, EDKM07]. equitable [CAY10]. equivalence [Kau07]. Equivalences [AGW23]. Equivalent [CMWW19, GMW20]. erg [PUW08]. Error [JW13, AKPS10, RS11a]. Errors [KT19]. Essentially [KPR16]. Estimating [EHL^{+}18, CCM10]. Estimation [BHPR^{+}20, BT24, HK25, JW23, GMV09]. ETH [BLL^{+}24, KBNvL20]. ETH-Tight [BLL^{+}24, KBNvL20]. Euclidean [AFHS20, BT25, BKM15, CS08, LT24, Sol13]. Eulerian [FLM^{+}12, KK13]. Evaluation [DHK16]. Even [CP12, KKR19]. everywhere [CPL12]. evolution [FBV09]. Exact [AFK^{+}18, CMT22, CP12, GHT18, HJT17, Vio05, BFK^{+}12]. Excluded [FLST18, LR15]. Excluding [Neu24, RW09]. Expansion [CLL^{+}12, PR12]. Expected [DHPR16, BS06a]. Experimental [DI06]. Experiments [FGGV06]. explicit [RW10a]. Exploitation [GSV20]. Exploration [DP14, GP19, TSZ14, AGP^{+}11, CFI^{+}08, DKK06]. Exponential [ANFS17, CKP^{+}19, CNP^{+}22, DHM^{+}14a, FP13, GKLT09, JMR22, KLP^{+}16a]. exponentially [PR08]. expression [BFG09]. extended [HPR14]. extension [GN14, KMMP07]. extensions [BHZ13]. External [BF24, SSTV23, CFLM07, LK08]. Extremal [CC25]. extreme [GGM10].

Face [Fil25, DS11]. Faces [KBNvL20]. Facility [ABF^{+}18, ANFS17, KS16, MP22, ST10, BH12, Fot11, FS11, MV08, Svi10, SS08b]. Factor [Adj19, ABCH25, HHL^{+}16, Cha20, Jac11]. Factors [Neu22]. Failing [CMT22]. Failures [CC25]. Fair [AKS17, BK24, PS16, GMP05, GN14]. False [Pag18]. Families [CEV21, FLPS17, GKM16, AG10]. family [ASS08]. Fan [HS18]. Fan-Out [HS18]. Fast [AL13, BGGS24, BHK^{+}16, BCP24, CN14, CJL17, ES16, HU20, KMTS23, KKK^{+}10, PT11, WY13, YZ05, BS10, BD11, KSS09, RW10b]. Faster [AKLR20, BKK17, BH19, CMV16, CD17, CP12, DGM18, EW20, FLN14, FPS22, FZ07, GS17, GW20, HW19, JS25, Jez15, Kar08, KKK18, KX19, KLP16b, LNR^{+}14, RSS06, Ter25, BG11, Dji10, Rod08, TM08, MM09]. father [SSS^{+}11]. Fault [BCHR20, HHL^{+}16, PP16, PP18, SS08b]. Fault-Tolerant [HHL^{+}16, PP16, PP18, SS08b]. Faulty [KT19]. Feedback [ALM^{+}19, CCHM15, KR18, LN22, LRS18, LMM^{+}21, RSS06, Tho10a, Tho10b]. Few [BHK^{+}16, KBNvL20, GW07]. fewer [PR08]. FFT [GKM16]. Fibonacci [BLT25]. FIFO [And09]. File [BBM19]. Fill [BCK^{+}20, KMM11]. Fill-in [BCK^{+}20]. fillup [JS07]. filter [CDHW09, DH12]. Filters [NE19]. Filtrations [BS18]. Find [ATG^{+}14, AAY10, RW09]. Finding [ALLT11, BK24, CDEM10, Cab10, CAY10, CRSS22, CSS21, FPS22, GN08, Heg06, HMS07, KKK18, MOR13, VWY10, And10, RSS06, Joh07]. Fine [GHK^{+}24, dBBJW21]. Fine-Grained [GHK^{+}24, dBBJW21]. finite [BCN12, CFI^{+}08]. Firefighting [ABZ19]. First [GIKW19, HKN17, EP05]. First-order [GIKW19]. Fit [CCDL16]. Fixed [BBM19, BBT12, CM15, CCHM15, DFHT05, DHM14b, FPS22, FGK^{+}24b, GST23, MV15, DJP^{+}12, EKS05, KKW12, RSS06, CGK^{+}11]. Fixed-Parameter [BBT12, CM15, CCHM15, DHM14b, FGK^{+}24b, MV15, DFHT05, CGK^{+}11]. Flajolet [SSS^{+}11]. flash [BAT11]. Flat [AAE^{+}25]. flexible [AR05]. Flip [BBT12, BCC21]. Flips [IIK^{+}23]. Flow [AGLW18, KKPW24, KT18, LL25, AF07, AKR12, AGG10, BD07, BFKS14, BKMSS11, CMS07, EF12, GMT11, HL06b, IM12, Kar08, TM08]. Flow-augmentation [KKPW24]. Flows [CCG^{+}24, Wan22, BEH^{+}10, EKS05]. Fooling [Har18]. For-All [GLPS17]. Forbidden [ACGP16, GJLS17]. Forbidden-Set [ACGP16]. Forest [BKM15, CDKL20, CJKV24, DKN17, FL25, GHNR10a, GHNR10b]. Foreword [CRR13, Epp07, Gab09, HT10, LOM06, Mat10, Gab05]. Form [BLS23]. Formulas [KT19]. formulation [CKL^{+}09]. Formulations [BGK^{+}25]. Foundations [FJS14]. Fourier [MRR06]. FPT [BD11, CDP19]. FPTAS [KKW12]. Fraction [LZ25]. Fractional [CZ24, GM14, CKK10b, Kar08, Mar10]. Framework [Coh18, Har16, AKR12]. Frank [Cla10]. Fréchet [AFK^{+}15, BKN21, BFL^{+}23, CD24, CW10, FF23, GSW24, HPR14]. Free [BFF^{+}22, CHJ^{+}18, GHLL16, KRX16, ND25, SHHA16, ALL^{+}25, BNCS08, BDH^{+}20, CKS09, DKT11, FGZ21, GKPP22, LCS^{+}19, LPV18, Saw06]. Frequency [Coh18]. Frugal [AT07]. full [FMMN07, KKK^{+}10, MN08]. full-text [FMMN07, MN08]. Fully [BKS12, BGK^{+}22, CDP19, DFL^{+}25, FLS^{+}18, Iba08, Jez15, NS14, NS16, OSSW20, AHTL05, Elk11, Rod08, RNO11]. Function [BCP13, DHK16, DP06, SS09]. Functional [NZ24, NS14, CMY11]. Functions [BJLY17, CD17, Fel17, MPQS20, MA16, WZ16, AG10, GMT11, Vig14]. Fuzzy [BBB20].

galled [MSS11]. Gallery [BM20]. Galois [AK12b]. Game [EFF^{+}15, BCKV06]. Games [ADGH21, DFM23, FV19, CKK10a, DHMZ12, FFM12, FKS08, GMT11, Swa12]. Gap [FP13, HKP^{+}18, LL25]. Gaps [ABZ19]. gas [KMM11]. Gasoline [NRS18]. Gathering [DPP14, BKMSS11, RS11b, SZ10]. General [DPS18, ES16, EP16, ERV16, GMS19, MWsZ23, AAA^{+}06, JZ06, MR09, ABD^{+}08]. General-valued [MWsZ23]. Generalization [BR14, HHM^{+}18, JCCW22, KMSY25, CF05]. generalizations [VB08]. Generalized [ACER19, AFH^{+}16, BEKN23, CN19, CV20, FGG^{+}25, FMM^{+}25, HU20, CGR08, HL06a, Lev09]. generate [BS10]. Generating [BBHT17, BHPR19, Saw06]. Generation [CMM25]. Generators [ELMR21]. Generic [MRR06, OOB24, RT22]. Genome [CRTZ24]. Genus [PPSV18, Dji10]. Geodesic [Cab22, CW10, KR19, OGGW10]. Geometric [ACS^{+}22, CGK^{+}11, Cha20, CJKV24, DI25, GGI^{+}21, GS18, GRSW16, Vig14, BCEG07, BCH^{+}12, CCHP12, Epp09c, GLNS08]. Geometry [BFF^{+}22]. Getting [PUW08]. Gibbs [HK25]. Girth [DKR16, RT13, Dji10]. Gives [DH18, CFR10]. Good [ADK16, LZ25, CFR10, Kol08]. Grained [GHK^{+}24, dBBJW21]. Graph [AGW23, BRW16, BF24, CMA^{+}19, CDP21, FGL^{+}19, Gab16, GHK^{+}24, GP19, GW22, KMZ18, KN16, KRS19, SST22, SW20b, SYZ20, BKS12, BLPS13, CFI^{+}08, DFR09, DKK06, EFKN09, GKLT09, RW09, SS09, Wil10]. Graphs [ACGP16, ALL^{+}25, AFT19, ASS19, ADF^{+}15, ADD^{+}18, AKLR20, ABCH25, BLL^{+}24, BGGS24, BDH^{+}20, BSWN15, BK16, BDLP23, Cab19, Cab22, CKM^{+}24, CMT22, CEV21, CR18, CDP19, CGH17, DHK14, DPS18, ES16, EP16, EMTG23, EHL^{+}18, ELMR21, FL25, Fil25, FGZ21, FMR24, FMM^{+}25, FLS^{+}18, FLST18, FM25, GILP16, GLSY23, GST23, GJNW23, GNSW20, GKPP22, GSW24, JS23, KMTS23, KK13, KKPW24, LR15, LPV18, LMS18, LM19, Neu24, OSSW20, PSZ24, PPSV18, Wan22, WY16, AS07, ALM^{+}12, ASS08, BFKS14, BS06a, BHKK12, BKM09a, BKM09b, BGPV08, BHLR10, Cab10, DC05, Cha12, CEGS11, CSTW12, DS11, DFHT05, Dji10, DKT11, EPR10, Epp09a, Epp09c, GKK10, GT08, HKRL07, HSS07, Iba08, KMW10, KP08, KK06, Lau06, MR09, MZ12, PS10, PRV11, PRS12, RTZ08]. graphs [RZ12, RST14, VWY10, VH05, YB12]. Gray [KL06, MN18]. Greedy [FN20, LT24, CKS05, Cla10, CNP^{+}11, FP10, GR10, RS11b]. Gromov [AFN^{+}18]. Group [DHK14, KW16a]. Groups [CZ18, Neu22, AMR09, AK12b]. Guarantees [TY18, BST08]. Guarding [AFK^{+}18]. Guessing [AGKS07]. Guest [Buc08]. guided [CFI^{+}08]. Guillotine [KLM^{+}25].

Hadamard [CI17]. Half [CCLS25, HJT17]. Half-Integral [CCLS25]. halfplanes [CKS09]. Hamiltonian [FGL^{+}19]. Handed [AAGH25]. Hanoi [ASS08, BS06b, DS08]. Hard [CDL^{+}16, GKM16, BH12, Bon22, Cha20, Vis08]. hard-capacitated [BH12]. Hardness [BGK^{+}25, BM20, BKN21, CMVZ16, DKR16, CPR^{+}11, KNS^{+}07, KM12]. hash [AG10]. Hashing [AFCK25, Pag18, Wal23, Wal25, DK12, Jan05, Vio05]. Hashing-Based [Wal25]. Haste [SHHA16]. Hausdorff [AHPSW10, AFN^{+}18]. haystacks [Joh07]. Heap [CR18]. Heaps [BLT25, Elm17, HKTZ17, ST25, KT08]. heaviest [VWY10]. Heavy [BDW19, BNS19]. Helps [CCG^{+}24]. Hermite [BLS23]. hidden [FKW11]. Hierarchical [CGMZ16, LK08, ST10]. High [AFS18, AEP18, BCSV25, Bła20, GHM^{+}24, FR10, GLS10]. High-Dimensional [AEP18, BCSV25, GLS10]. High-Quality [AFS18, GHM^{+}24]. Higher [ACHKP21, TY18]. Highway [JS23]. Hinders [CCG^{+}24]. Hitters [BDW19, BNS19]. Hitting [ACS^{+}22, FLL^{+}19, GJLS17, GST23]. Hoc [KMPS16]. Holant [BG20, GLLZ21]. Hole [CSS21]. Hollow [HKTZ17]. Homomorphic [GLSY23]. Homomorphism [GHK^{+}24]. Homomorphisms [FMR24]. hop [FCFM09]. hop-optimal [FCFM09]. Hopcroft [CZ24]. hospitals [KMMP07]. hospitals-residents [KMMP07]. hotlink [Jac11, LM11]. Hub [BGGN16]. Hulls [DEK21, BB09]. Hunt [BDLP23]. Hunts [TSZ14]. Hyperbolic [BFF^{+}22, Epp09b]. Hypergraph [GHM^{+}24, Har19, Neu22, AFS12, LDX09]. Hypergraphs [FPZ23, ND25, RKH20]. Hyperplane [AFHS20]. hypertree [Mar10].

I/O [AAY10, CFRY22, MZ12]. I/O-Efficient [CFRY22, AAY10, MZ12]. Ideal [Mas21]. identical [GP08]. identification [CPR^{+}11]. Identity [BCSV25]. idle [AA14]. IDs [DLS14]. II [AAHP^{+}16, KKPW24, SST22]. III [FGL^{+}19]. Im [BCKV06]. Image [BMR22]. Images [RT14]. impatient [BCC^{+}10]. Implementation [BBHT17, HMS07]. Implicit [FLL^{+}19]. Imprecise [AFHN16]. Improved [AER15, BKK^{+}19, BFG09, BBT12, BKN21, BPR^{+}17, CHA18, CJ18, CMO^{+}08, CKP12, CLL^{+}12, CJST07, DKN17, EFM^{+}16, FGK^{+}24c, GS09, GHKS06, GHKS08, GNSW20, HIMY07, HSS07, KK16, LM11, Lac13, LS24, SW20b, Swa16, YLW08, AR06, CCKR11, CGR08, EK07, BKLP07, GHKS13]. Improving [BCC21, CKS19, GW22, JCCW22]. In- [JMR19].

Inapproximability [ABCH25, BG21, GRSW16, KS16]. inclusion [BG11]. incomplete [GLS10]. Incremental [BFGT16, GHT18, HKM^{+}12, AFM08]. Independence [BGGS24, FGYZ22, LPV18, PT16]. Independent [BHPR^{+}20, BN14, GHH22, GKPP22, Har19, LM19, LPS^{+}20, AGM^{+}08, Epp09a]. index [FV10]. indexable [RRS07]. Indexes [CEK^{+}21, BHMS11, BCH^{+}12, CHLS07, FMMN07, MN08]. Indexing [BN14, BCKM20, BFG^{+}16, BGS23, Pre21, FGGV06]. Individual [Jan05, Vio05]. Induce [FP13]. inequality [BGLZ09, Blä08]. Infants [GKM16]. infeasible [BST08]. Infinite [CKL21a, VZ21, ASS08]. Information [GMP17, AHTL05, BCKV06, CFK^{+}07, EF12, GMV09, KKK^{+}10, RS11b, MNS12]. Injecting [BLW09]. Input [GSW24]. Insert [BPL18]. Insertion [BDW19, AHRT05, Jan05]. Insertions [ACHM22]. Inspired [ALM^{+}19]. Instability [And09]. Instances [DKR16, Vis08]. Integer [BLS23, EW20, CMS07, CLLJ08]. Integral [CCLS25]. Integrality [ABZ19, HKP^{+}18, LL25]. interchanges [FZ07]. Intersection [AAE^{+}25, HIMŽ19, BK08, Epp09c]. Interval [ALM^{+}23, BFPP18, CM15, EHR16, ND25, ALM^{+}12, BHLR10, PRV11]. intervals [BNCS08]. Introduction [BKN^{+}25, DMT24, ELMS25, GBFC07, HK22, KADK22, LPW20, MWY19]. Invariant [Wim16]. Inverse [BLW09, Cab22, NZC11]. Investigation [KR18]. IP [VZ21]. IPs [KR24]. Irreducibles [BHK^{+}16]. Ising [SS09]. isolated [II09]. Isomorphism [ABH^{+}18, GNSW20, Neu22, Neu24]. Isomorphisms [iKMNZ25]. Issue [BKN^{+}25, BGN^{+}18, CRR13, DMT24, ELMS25, HK22, KADK22, LPW20, MWY19, RRSW16, Epp07, GBFC07, Gab09, HT10, Mat10]. Items [CZ18]. Iterations [MPQS20]. Iterative [CLL^{+}12]. IV [CMM25].

Jitter [HS09]. jobs [DJP^{+}12, GHKS08, GP08]. Johnson [AL13, JW13]. joint [ELR^{+}08]. Journey [HPJ21]. Jump [CCW18]. Junta [LCS^{+}19]. Juntas [Har18].

Kernel [ALM^{+}19, ALM^{+}23, BRFF^{+}12, CGH17, KW14, KPR16, Tho10a, Tho10b]. Kernelization [DLS14, GJLS17, Kra14]. Kernels [FLST18, FLL^{+}19, KLP^{+}16a, KRS19, Wah22, PRS12]. keys [BB08]. Kinetic [AKS08]. Knapsack [DHK16, GGI^{+}21, GNR16, HW19, MW24, Swa16]. Knapsacks [CZ18]. Know [SZ20]. Known [LMS18, CRV11]. Knuth [BGLZ09]. Kt [BDH^{+}20]. Kt-free [BDH^{+}20].

Label [BGGN16, CFI^{+}08, DKR16, MMS14]. Label-guided [CFI^{+}08]. Labeling [Kor10]. Labels [ACGP16]. Languages [CMM25]. Laplacian [JS25]. Large [DKR16, AGG10, RW10b]. Latency [BMSV^{+}09, INV16, GMT11]. Latency-constrained [BMSV^{+}09]. lattice [NS09]. Lattices [BHK^{+}16]. law [CSTW12]. layouts [BGPV08]. LC [JS07].

Leader [ARS^{+}14, CKP^{+}19, CDJ23, GMP17, MPY22, KKK^{+}10]. Leaf [Laf23, BLS08, DV10, BD11]. Lean [CKL^{+}21b]. Learned [BFV22]. Learning [BCSV20, Wim16]. least [Kol08]. leaves [BRFF^{+}12]. left [DMM^{+}12]. left-to-right [DMM^{+}12]. Lemma [CHL^{+}20, EW20, HH17, Har18, Har21, Har16]. Lempel [BFG09]. Length [BEH^{+}10, Cab22, LZ25, BB08, GP08, KK06]. Length-bounded [BEH^{+}10]. lengths [KMW10, MZ12]. Leontief [GN14]. Level [KR19, GRR06]. level-ancestor [GRR06]. Levels [MN18]. Lewis [WY25]. Liars [Sha16]. Light [CWN18]. Lightweight [ES16]. Like [CCLS25]. Limitations [IMM08]. limited [KKM11]. Limits [KW16a, Joh06]. Lindenstrauss [AL13, JW13]. Line [BHD^{+}21, KPR16, PS23, BFKS14, GS09, RS11b]. Linear [AW19, BCKM20, BPGN09, BSWN15, BCAD23, CHA18, CRTZ24, CCKN19, CLL14, DJO^{+}25, EP16, HS18, iKMNZ25, KLP^{+}16a, KKR^{+}20, LNR^{+}14, LMPS18, LRS18, MA16, PT16, RS17, TY18, WY16, BB09, BG11, BLS08, CKK10a, CKL^{+}09, DKT11, EPR10, FP10, GHPT05, Jan05, KMW10, MOR13, NZC11, RW09, VH05, Vio05]. Linear-Size [EP16]. linear-space [KMW10]. Linear-Time [BCAD23, CRTZ24, KKR^{+}20, RS17, BCKM20, BLS08, RW09, VH05]. Lines [DT16, BB09, GLS10]. Link [CBFWW15]. Linkages [CCLS25]. Linking [Gab17]. links [GW07]. List [FMR24, AGKS07, AGvS13]. Load [ABF^{+}18, BHS14, BKK^{+}19, GPSS15, Wal23, EDKM07, GMP05]. Load-Balancing [BKK^{+}19]. loads [And09]. Local [BNS19, CHL^{+}20, DP14, ER17, FGG^{+}25, HH17, Har16, Har18, Har21, JMR22, PS16, And10, BCM11, KP08, LK08]. local-ratio [BCM11]. Locality [Pag18]. Locality-Sensitive [Pag18]. Locating [KT19]. Location [ABF^{+}18, ANFS17, AK18, Cha13, CL22, KS16, MP22, AGG^{+}09, AMM07, BH12, BCH^{+}12, CDI^{+}12, Fot11, FS11, MV08, ST10, Svi10, SS08b]. Locations [GLLZ21]. Log* [CZ24]. Logarithmic [Cha20, EP16, AGP^{+}11, BFKS14]. Logarithmic-factor [Cha20]. Logit [FV19]. long [GN08]. Longest [AEMV24, BCAD23, JN24, EP05]. loopless [KL06]. Lopsided [Har21, Har16]. Lopsidependency [Har16]. losing [KM12]. lot [ELR^{+}08]. Lottery [HPST19]. Lovász [CHL^{+}20, HH17, Har16, Har18, Har21]. Low [CHA18, CKM^{+}24, CDP21, FLS^{+}18, FGL^{+}20, KRX16, NS09, KP08, Pet09]. Low-dimensional [NS09]. Low-rank [FGL^{+}20]. Low-treewidth [CKM^{+}24]. Lower [ACHKP21, ABHS22, BN15, BCK^{+}20, CRSS22, CDKL20, DI25, DLS14, Kra14, KT18, SCRS17, Soc16, Svi10, AR08a, AGvS13, CGK^{+}11, Cha10a, MSS11]. Lower-bounded [Svi10]. LP [CCW18, HKS11, Li17, LL25, MMS14, VZ21]. LP-based [MMS14]. LZW [Gaw13].

Machine [GSV20, HJ15, JMR19]. Machines [EMS23, SZ20, CKK10b, GP08, GLP06, TM08]. maiden [EP05]. Maintaining [ACHM22, AHTL05, Elk11]. Maintenance [HKN17, HKM^{+}12]. majorization [GMP05]. Makes [SHHA16]. Makespan [GNR15]. Making [Ruž09]. malleable [JZ06]. Management [AER15, AKM08, CJST07]. Many [BKMV20, BK24, BRFF^{+}12, Joh06]. Many-to-One [BK24]. Many-visits [BKMV20]. Map [GSW24, DFHT05]. Mapping [CDD^{+}15]. Maps [iKMNZ25]. Market [FGK^{+}16]. Markets [BGH19]. Markov [FBV09, MTZ10]. marriage [HIMY07, IMY10]. mask [BF09]. Massey [KY13]. Massively [CDP21]. MAST [BPGN09]. Matching [AKLR20, BFH21, BG14, CCW18, CLL^{+}12, CHLT14, DPS18, EMTG23, EHL^{+}18, FLN14, FGG^{+}25, Gab25, Gaw13, GSW24, HTWZ19, Jez15, JN24, LS20, NS16, PS19, PS23, ALLS07, BFG09, DC05, CS11, CM07, GGN06, HLS07]. matching-covered [DC05]. Matchings [BK24, Kav24, AFS12, GKK10, GKP08, IKM^{+}06, KMMP07, KR16, Mes14, RW10b, VH05, GKK23]. Matrices [DHS16, FLS^{+}18, GMW20, KMNS17, KLP16b]. Matrix [BLS23, FGL^{+}20, SWYZ21, WY13, KY13, YZ05]. Matroid [CLL14, GNR16, KMSY25, Swa16, Ter25]. Matroids [DFL^{+}25, FGLS19, KKR19, KW14, LMPS18]. Max [AKS17, ABG16, BR14, BCC21, BJLY17, KT18, PRV11, PS16, Wan22, BD11, CRZ20, GS17]. Max-coloring [PRV11]. Max-CSPs [CRZ20]. Max-Flow [KT18]. Max-Leaf [BD11]. Max-Min [PS16]. Max-Sum [BJLY17]. maxima [DMM^{+}12]. Maximal [BFH21, Har19, NS16, PS10, Epp09a, GKP08, IKM^{+}06]. Maximin [AMNS17]. Maximization [BF18, BFS19, CHJ^{+}18, DFL^{+}25, EMS23, BNGK^{+}09]. Maximizing [Fel17, GPSS15, MS17, WZ16]. Maximum [ER17, FGK^{+}24b, Gab25, GMW20, GKPP22, IIK^{+}23, KMNS17, LS20, MMS14, PS19, CMM09, DV10]. Maximum-Cut [ER17]. MaxMin [GNR16]. may [EK06]. Mazing [AGLW18]. mean [HL06b]. meanders [BS10]. meandric [BS10]. Means [BBB20]. Measure [DI25, GS17, AZ08, FGPS08]. measures [ABS10]. Mechanism [CKK10b]. mechanisms [AT07]. Median [BDR23, BPR^{+}17, HHL^{+}16, Li17, Swa16, Cha20, YLW08]. Medium [ARS^{+}14]. Meet [BKK17, CKL^{+}21b, CPL12]. Meet-in-the-Middle [BKK17]. Meeting [KMTS23]. meets [AFS12]. Melding [MTTZ06]. Membership [Mas21]. Memory [DEK21, FP13, SSTV23, AKM08, AGP^{+}11, AC10, BAT11, CFLM07, GLPP08, KKM11, LK08]. Memory-Adjustable [DEK21]. Memoryless [CV20, Fot11]. MEMs [Nav25]. Menger [CCLS25]. Menger-Like [CCLS25]. mergeable [GKS^{+}11]. mesh [FSP08]. Message [PRS20]. Message-Optimal [PRS20]. Metastability [FV19]. methods [BCN12]. Metric [AFN^{+}18, ACE^{+}23, GW22, HPMO23, ABS10]. Metrics [BEKN23, CGMZ16, CJ18, CJJ20, CGR08, GS09]. Middle [BKK17, MN18]. Mighty [DS19]. Migration [BBM19, GSV20, GHKS06, GHKS13]. mimicking [Wah22]. Min [AKS17, BR14, BSWN15, DHK16, GHT18, PS16]. Min-Cut [GHT18]. Min-Knapsack [DHK16]. Min-Max [BR14]. minima [BLW09]. Minimal [AFCK25, BDH^{+}20, FGPS08]. Minimization [HL13, AF07]. Minimize [RKH20, AAG09, EK06, TM08]. Minimizing [BD07, BKMSS11, DHM^{+}09, DHM14b, FS11, GLP06, HL06b]. Minimum [ABF^{+}18, ABD^{+}25, AKLR20, BBP23, BCK^{+}20, BBT12, FPZ23, GPSS15, GNR15, HS18, HVV19, INV16, KLPP23, MV15, MM09, PRS20, TY18, AGM^{+}08, ABD^{+}08, BCD12, CKL^{+}09, CLLJ08, CKS05, CS08, Fuj12, GKLT09, HL06a, HLS09, JR05, KKW12, Lev09, Nut12, PR08]. Minimum-Cost [MV15, AKLR20, Nut12]. Minimum-Flip [BBT12]. Minimum-Load [ABF^{+}18]. minmax [YLW08]. minmax-regret [YLW08]. Minor [CEV21, GST23, LR15, SST22, RW09]. Minor-closed [CEV21, SST22]. Minors [FLST18, GJLS17]. Minset [Gab16]. Minset-Poset [Gab16]. minsum [GHKS08]. Minwise [PT16]. MIS [FGG^{+}25, FN20, OSSW20]. Mischief [DPP14]. mismatch [JN24]. Mixing [FGYZ22]. mobile [FS11]. Model [AN16, CMR18, CDL^{+}25, FTK20, HPST19, LS20, PS19, And09, FCFM09]. Models [GST23, JW23, AR05, FBV09]. Modern [BBHT17]. Modification [LPS24]. Modification-disjoint [LPS24]. Moment [JW23]. moments [LK08]. Monge [GMW20, KMNS17]. monitoring [CMY11]. Monotone [AFCK25, BJLY17, FGK^{+}24a, IIK^{+}23]. monotonic [IMM08]. Monotonically [CdM22]. Monotonicity [KR19, BGLZ09, FR10]. Morphing [BLPS13]. Moser [Har16, HS17]. motif [FKW11]. Movement [DHM14b, HKKK16, DHM^{+}09, FS11]. moves [CM07]. moving [Kol08]. MST [Fuj12]. Multi [ADK16, GNR15, BS06b, CRR09]. multi-peg [BS06b]. Multi-Pivot [ADK16]. multi-scheduling [CRR09]. Multi-Vehicle [GNR15]. multicast [CRR09]. multicasting [AR05]. Multicommodity [CMS07, AKR12, BKLP07, Kar08]. multicover [CCHP12]. multicriteria [Man12]. Multicut [Wah22]. Multicut-mimicking [Wah22]. Multicuts [RS17]. Multidimensional [YZ12]. Multiflow [Hir19]. multigraph [SS08a]. multigraphs [HKS11]. multihop [AZ07]. multilabeled [BHMS11]. Multilevel [KS16]. Multipartite [EJK08]. multiparty [FIM^{+}06]. Multiple [BJKK18, CZ18, CKR12, BHLR10, FKW11, HS09, VP07]. multiple-interval [BHLR10]. Multiplication [WY13, IP11, VP07, YZ05]. Multiplierless [VP07]. multisets [RRS07]. Multistage [KR24]. multivariate [Epp06]. Multiway [Wah22].

name [AGM^{+}08]. name-independent [AGM^{+}08]. Nash [Das13, DFM23, EDKM07, GKK23, IIK^{+}23]. Natural [Li17]. Navigating [BFKR21]. Navigation [DEK21]. Near [ACHM22, ADKS25, BSWN15, Bre23, CDJ23, CMM09, CWN18, GLPS17, Wei22, WY16, CCM10, EPR10]. Near-Linear [BSWN15]. near-linear-time [EPR10]. Near-Optimal [ACHM22, CDJ23, CWN18, GLPS17, ADKS25, Bre23, CMM09, CCM10]. Nearest [AAHP^{+}16, AEP18, Gab17, IN07, AKS08]. Nearest-Neighbor [AAHP^{+}16]. Nearest-neighbor-preserving [IN07]. Nearly [CI17, DKN17, KBNvL20]. needles [Joh07]. Negations [FGRZ25]. negative [CGR08, KMW10]. negative-type [CGR08]. Negatives [Pag18].

Neighbor [AAHP^{+}16, AEP18, IN07]. Neighborhoods [AFHS20, CJ18]. Neighbors [Wei22, AKS08]. nested [Kau07]. Network [BK16, CLS^{+}22, CEV21, CFM21, DP14, EFF^{+}15, Fuk17, GHLL16, PPSV18, AR05, AAA^{+}06, And09, AZ08, CEGS11, CGNS08, DHMZ12, EKS05, FCFM09, HKKN12, JR05, Swa12]. Networks [BHS14, BFF^{+}22, CN14, CLNV14, DGG^{+}20, HKN17, KRX16, KMPS16, MPY22, Wah22, AZ07, AGG10, BH12, BMSV^{+}09, MSS11, NHK08, PR12]. nilpotence [AK12b]. No [EMS23, FG08, BRFF^{+}12, PS23]. Node [CCW18, CEV21, DHK14, Hir19, HKRL07, RS09]. Node-capacitated [Hir19, HKRL07]. node-cost [RS09]. Node-Weighted [CCW18, DHK14, CEV21]. Noisy [BFKR21, BFKR25]. Non [CGK20, OOB24]. Non-recursive [OOB24]. Non-Uniform [CGK20]. nonabelian [AMR09]. nonavailability [DJP^{+}12]. Nondecreasing [Wil10]. Nondeterminism [Kra14]. Nonlinear [HJ15]. nonmetric [ABS10]. Nonmonotone [ADV^{+}16]. nonpreemptive [GP08]. Nonsingular [BLS23]. nontrivial [KS08]. Normal [BLS23]. note [Alb10a, AHRT05, Alb10b]. Nothing [AFH^{+}16]. Novel [NW07]. Nowhere [KRS19]. NP [DS19, GKM16, Joh05, Joh06, Joh07]. NP-completeness [Joh05, Joh06, Joh07]. NP-Hard [GKM16]. NP-Mighty [DS19]. Number [SZ20, CFR10, KKW12, KS08]. Numerical [KLP16b].

O-Efficient [CFRY22, AAY10, MZ12]. Objects [AAE^{+}25]. Oblivious [BFCF^{+}17, CCW18, CR18, FV16, FLPR12, HKRL07, Har21, WW22]. Obstacles [AFS18, CW15]. occurrences [BCN12]. Odd [ALL^{+}25, CSS21, FGL^{+}19, KW14]. off [DFR09]. offline [BNCS08]. One [BK24, Kav24, CDEM10, Fot11]. One-Sided [Kav24]. Online [ACE^{+}23, ADKS25, AGGP21, BODD^{+}20, BT25, BJKK18, BHD^{+}21, BCI^{+}23, BFS19, CHJ^{+}18, CKS09, DH18, EMS23, EV06a, GSV20, GP08, HTWZ19, JLSS12, JPS22, KB06, LS20, LMMW16, MNS12, WY25, AAA^{+}06, AAG09, AKL10, BNCS08, BF07, CCL^{+}09, CJST07, CNP^{+}11, EV10, GS09, GMP05, IM12, PU07, PRV11, YZ12, HCT^{+}11, LK08]. open [BS10, GHKS06, GHKS13]. Opposite [SSA25]. Opt [ERV16]. optical [AZ08]. Optimal [ACHM22, AL13, AGvS13, AAdFM22, AD16, ABCH25, BN15, BDW19, Bła20, BRW16, BDLP23, CMA^{+}19, CDJS17, CDJ23, CWN18, CI17, CEK^{+}21, CGMY22, DH18, DKK06, Elm17, Gaw13, GLPS17, GMS19, GIN^{+}17, GT08, GRSW16, HT21, Han07, JW23, JW13, LMS18, LMMW16, MP22, PRS20, Pre21, Wei22, ADHY08, ADKS25, BKMV20, Bre23, CCM10, CMO^{+}08, CMM09, DMRW09, FCFM09, FSP08, GK09]. Optimal-Time [CEK^{+}21]. Optimality [DS08, CNP^{+}11]. Optimally [EK06, KS08, TM08, Wil10]. Optimization [BGGN16, BHLR10, BPR^{+}17, GM12, GNR16, AAA^{+}06, GJL12, Vig14, MNS12]. Optimizing [CCL^{+}09]. Oracle [BBP23, BSWN15, EP16, LR15, CSTW12]. Oracles [BHPR^{+}20, CMT22, GW20, Har21, KK06, WY13, BS06a, GLNS08]. Order [HL13, BF07, BM08, GIKW19]. Order-Preserving [HL13]. Ordered [KR19]. Ordering [BR16, CFR10, AFM08, CDHW09, DH12, KB06, HKM^{+}12]. Ordinal [ABD^{+}08, GRR06, HMS12]. Orientability [Wal25]. Orientation [IIK^{+}23]. orientations [FLM^{+}12, FSP08]. orienteering [CKP12]. Orthogonal [BRW16, Cha13, CW16, CW21, BLPS13]. Other [BCK^{+}20]. out-trees [BRFF^{+}12]. Outlier [KKK18]. Outliers [CN19, FKRS19, HPST19]. output [ST08]. output-density [ST08]. Overhead [ACHM22]. Overlapping [Wal23].

P2P [NW07]. Packet [KMPS16]. packets [JLSS12]. Packing [AARA23, ADKS25, BSSX20, CZ18, CS07, EPR13, FLL^{+}19, KLM^{+}25, BNLT07, CMS07, KNS^{+}07, HCT^{+}11]. Packings [BKK17, GGI^{+}21, LPS24]. Paging [BBC^{+}25, JPS22]. pair [AKS08]. Pairs [KT18, Cha12, DI06, MTZ10, RS11a]. Pairwise [Cab19, GW07]. pants [Epp09b]. Parallel [CDP21, DH12, FN20, HH17, Har18, Har21, CKS09, GKK^{+}09, Han07]. Parallelizable [Kol25]. Parameter [BBT12, CM15, CCHM15, DHM14b, FGK^{+}24b, GST23, HK25, MV15, DFHT05, RSS06, CGK^{+}11]. Parameterized [AMW20, BFPP18, BCK^{+}20, BM20, CFM21, CHLT14, CNP^{+}22, FL25, FLS^{+}18, GHK^{+}24, GWZ21, KW16a, LPS24, LNR^{+}14, LRS18, LPS^{+}20, MP22, PSZ24, RS17, GJL12, HLS07, SST22]. Parametric [Epp18, FBV09]. parentheses [LY08]. Parity [CLL14]. Part [FMM^{+}25]. Partial [GMW20, JS07, KMNS17, MSS11]. Partially [ADF^{+}15, HKN17]. Partition [LR15, Ter25, CLLJ08, CKS05, SS09, ZO08]. Partitioning [AAGH25, AD16, GHM^{+}24, SYZ20]. Partitions [Fil24, CAY10]. Party [KR16]. pass [Fot11]. passes [DFR09]. Path [ABHS22, AGLW18, AZ08, BCHR20, EK20, EP16, FGK^{+}24b, GWZ21, HMZ16, SP24, AT07, DI06, HSB07, MZ12]. Path-Reporting [EP16]. Paths [AFS18, BFF^{+}22, CW15, CR18, GW20, IIK^{+}25, KK13, KK16, LBSZ21, WY13, Cha12, CK07, DS11, Elk05, EPR10, GW07, HS06, HMS07, KMW10, KK06, MTZ10, RS11a, RZ12, Wil10]. Pathwidth [GJNW23]. Pattern [Gaw13, GGN06, Jez15, ALLS07, CS11]. Patterns [BGS23]. PCSPs [NZ24]. Peeling [Wal25, ALLT11]. peg [BS06b]. Perfect [AKLR20, AFCK25, BCP24, BG21, GKK10, HHM^{+}18, ND25, AG10, BCKV06]. perform [BCM11]. Performance [HJ15]. periodic [BCMSM12]. Periodicity [EMS10]. Periods [MPQS20]. Permanent [DHM^{+}14a, And09]. Permanently [BFKR21, BFKR25]. Permutation [CMM25, Wim16]. Permutation-Invariant [Wim16]. Permutations [BBHT17, EPR13, RW10a]. permuterm [FV10]. Persistent [Cha13]. Perspective [Fil25]. Perturbation [BHW20]. Phase [BBM19, CFH07, FP10]. Phase-based [BBM19]. Philippe [SSS^{+}11]. Phylogenies [HHM^{+}18]. piece [EP11]. Piecewise [MA16]. Piecewise-Linear [MA16]. Piles [DEK21]. pipelined [CDHW09, DH12]. Pivot [ADK16, AD16, WNN15]. Placement [EFM^{+}16, GKK^{+}09]. Planar [AAGH25, ASS19, AKLR20, BSWN15, BK16, Cab19, Cab22, CMT22, CEV21, DHK14, EHL^{+}18, Fil25, FM25, GST23, IIK^{+}25, KK13, KKR19, KBNvL20, MP22, PPSV18, Wan22, WY16, AMM07, BLPS13, BKM09a, BKM09b, CDI^{+}12, DS11, DFHT05, Dji10, DKT11, EPR10, GK09, GT08, KMW10, KK06, OGGW10, PS10]. Planarity [ADF^{+}15, ADD^{+}18, BFR23, KR19]. Plane [ABD^{+}25, CW15, GMS19, CS08, Saw06]. Planes [DT16]. Planted [HKP^{+}18]. player [GMT11]. player-specific [GMT11]. Plurality [ADGH21, DGM18, AR08a, AR08b, AR09]. Point [AK18, BHPR19, CRZ20, Cha13, CL22, HPJ21, KPR16, AMM07, BCH^{+}12, CFH07, CDI^{+}12]. Point-Width [CRZ20]. Pointers [BCFC^{+}25]. Points [AN16, ADGH21, DGM18, FPS22, ASW08, AHPSW10, CAY10]. policies [Jan05]. polygon [CAY10, CW10]. Polygonal [AFS18, BDR23]. Polygons [AW19, CDD^{+}15, MW24]. Polylogarithmic [ALM^{+}20, GHT18, Har18]. Polymer [BCP24]. Polynomial [AARA23, ALM^{+}23, ALL^{+}25, BCD12, BGK^{+}25, BKM15, CDL^{+}25, CDP19, DFM23, DHM^{+}14a, FLS^{+}18, GS17, GKPP22, KW14, KRS19, Laf23, LR15, LL25, LBSZ21, PRS12, PS16, SS18, AK12b, LDX09, SS09]. Polynomial-Space [GS17]. Polynomial-Time [BKM15, DFM23, FLS^{+}18, BCD12, CDL^{+}25, GKPP22, LDX09]. polynomially [Kau07]. Polynomials [Har19, MS17]. Polytopes [AAdFM22]. Popular [Kav24, KMSY25, Mes14]. Poset [Gab16]. Positive [BPR^{+}17, DHS16]. Potential [FV19]. Power [BCP13, CDD^{+}15, CSTW12, Hal12, ISG07]. power-law [CSTW12]. Powers [Laf23, BLS08]. PQ [BR16]. PQ-Ordering [BR16]. Practical [SCRS17]. practically [ADHY08]. Practice [CRTZ24]. precedence [DH12, JZ06]. precise [DK12]. Predecessor [Cha13, GMW20]. Predictions [ACE^{+}23, JPS22]. Preemption [BFS19]. Prefer [SSA25]. Prefer-Opposite [SSA25]. Prefer-Same [SSA25]. Preferential [ELMR21]. prefix [RRS07]. Preimage [MPQS20]. Preprocessing [BSWN15]. Prescribed [BH19]. Preservers [BW21, CC25]. Preserving [HL13, IN07]. Price [BGH19, GN14, DHMZ12]. Price-based [GN14]. priced [CFK^{+}07]. Pricing [AR05, BCC^{+}10]. Primal [DH18, VB08, BCM11]. Primal-dual [VB08, BCM11]. Principle [KLPP23]. Principles [ARS^{+}14]. priority [AHRT05, ADHY08, EJK08, MTTZ06]. Privacy [BNS19, FJS14, JMR22]. Private [SCRS17]. Prize [CJJ20, Fuk17, HKKN12]. Prize-Collecting [Fuk17, HKKN12]. Probability [Bła20]. Probably [LMS18]. probe [GGM10]. Probing [PT16, Jan05, Vio05]. Problem [AKS21, BEJK19, BBC^{+}25, BK16, CGK20, CCW18, CJJ20, CZ24, CMVZ16, CKL21a, DT16, EK20, Epp18, ER17, FGK^{+}24c, Fuk17, HW19, HVV19, KK13, KK16, KR18, Kra14, KS16, MMS14, Mas21, MW24, NRS18, SW20a, Soc16, TY18, AMR09, AR08a, BK08, BH12, BHZ13, BPGN09, BG11, BHKK12, BKMSS11, CF05, Cha10b, CEGS11, CCHP12, CLLJ08, CKS05, CM07, DS08, DV10, EV10, EKS05, FHR07, FN10, GS09, GKLT09, HIMY07, HL06a, IMY10, Jac11, JR05, Kar09, KKW12, KMMP07, KMM11, LM11, MV08, NS10, RS09, SZ10].

Problems [AZBG^{+}22, AGW23, AAE^{+}25, ALM^{+}20, AMW20, BG20, BFGT16, BR14, BDW19, BCC21, BR16, BFR23, BCK^{+}20, BM20, BSSX20, BF18, CFRY22, CN19, Cha20, CLL14, CV20, CFM21, CDL^{+}16, CGH17, CMWW19, CNP^{+}22, DI25, FLN14, FLL^{+}19, FGL^{+}20, FLK^{+}20, GMP23b, GKM16, GM12, GLLZ21, GWZ21, HPST19, Hir19, HL13, JW13, Khu05, Khu06, Khu07, KW16a, MP22, Mas21, PPSV18, Swa16, Wah22, AAA^{+}06, AKR12, BST08, BRFF^{+}12, BHLR10, CMM09, CEGS11, CKP12, CGNS08, CDHW09, CNP^{+}11, DFR09, EKS05, FS11, GJL12, HKKN12, HLS09, HSB07, Kar08, KNS^{+}07, Nut09, Nut12, SS09, ST08, YLW08]. processes [EP12, GLPP08, MTZ10]. Processing [GPSS15, JMR19]. processors [KS08]. Product [CDJS17, FLPS17]. Products [SWYZ21]. Programming [CHA18, CKS19, EW20, LNR^{+}14, MTK^{+}19, RST14]. programs [CMS07, FP10]. projective [AGvS13]. Promise [VZ21]. Proofs [BGK^{+}25]. propagation [IP11]. Properties [BMR22, CDJS17, GIKW19, KRS19, RT14, ABD^{+}08, MR09, PR12]. Property [CDJS17]. protocols [GN14, GR10]. Protrusion [KLP^{+}16a]. Provably [Kol08]. provider [BCKV06]. Proximity [EW20]. pseudo [II09]. pseudo-cliques [II09]. Pseudopolynomial [KX19]. Pseudoprimes [Sha16]. PTAS [AFHS20, CJ18, CJJ20, MZ23, MWsZ23]. Puzzle [NRS18].

QoS [CJST07]. quadrangle [BGLZ09]. quadrangle-inequality [BGLZ09]. Quadratic [GS18, MMS14, HLS09]. quadtrees [CFH07]. Quality [AFS18, GHM^{+}24]. Quantification [FJS14]. Quantum [AMR09, JN24, LZ25, MRR06]. Quasi [LR15, PS16]. Quasi-Polynomial [LR15, PS16]. Quasiconvex [Epp06]. Quasipolynomial [Wah22]. Quasirandom [DFS14]. Queries [AAE^{+}25, FF23, FGRZ25, GMW20, GIN^{+}17, GSW24, HMZ16, KMNS17, Pre21, BCH^{+}12, GRR06]. Query [BFKR25, FLM^{+}12]. Querying [CFK^{+}07, SWYZ21]. Quest [BG21]. Queue [AEMV24]. Queueing [BJKK18]. queues [AHRT05, EJK08, MTTZ06]. Queuing [CKR12]. Quickly [CW21, Oum08, Ruž09]. quickselects [MPV10]. Quicksort [ADK16, AD16, WNN15].

R [ADHY08]. R-tree [ADHY08]. Race [AA14]. Radio [MPY22, EK07]. RAM [Cha13]. Ramsey [ACE^{+}20, Kra14]. Ramsey-Type [Kra14]. Random [BBHT17, CP25, DGG^{+}20, ELMR21, GPSS15, HSS22, HTWZ19, BB09, BLW09, CFH07, FSP08, MZ12, PS10, PR08]. Randomized [AEP18, CV20, CKL^{+}21b, FN20, KSS09, KKM11, KW14, PR08, BKS12]. Range [ACY12, CW16, GIN^{+}17, WY18, BCEG07]. Rank [BFV22, HST15, IKM^{+}06, BF09, FGL^{+}20, Oum08]. Rank-Balanced [HST15]. rank-convolutions [BF09]. Rank-maximal [IKM^{+}06]. rank-width [Oum08]. Rank/Select [BFV22]. Ranked [SSTV23]. ranking [CFR10]. Rapid [FGYZ22]. Rate [RS11b]. Ratio [LMMW16, BCM11, BF07, BM08, Kar09]. Ray [Tao14, GK09]. Razborov [CW21]. Reachability [CC25, Lac13]. read [Wil10]. Real [BG14, BNGK^{+}09, BCMSM12, VWY10]. Real-Time [BG14, BNGK^{+}09, BCMSM12]. Realistic [GSW24]. Reallocation [BFCF^{+}17]. really [EP11]. Reasoning [AKL10]. Rebalancing [STK16]. receiver [EK06]. recognition [BLS08]. Recognizing [AFT19, Laf23]. recolorings [MSS11]. Recompression [Jez15]. Reconfiguration [LM19]. Reconstruction [CDL^{+}25, CJL17, KMZ18]. Recovery [GLPS17]. rectangle [ELR^{+}08]. rectangles [CKS09]. Rectangular [BGPV08]. Rectilinear [FLK^{+}20, GK09]. Recurrence [HJT17, Epp06]. recurrent [Kau07]. recursive [OOB24]. Reducing [CJ18]. Reduction [BGK^{+}25, MMS14, MOR13, NS09]. Reduction-Based [BGK^{+}25]. redundancy [BK08]. Reflex [CDD^{+}15]. register [DP06]. regret [YLW08]. Regular [FGLS19, BFG09, GKK10]. regulation [HS09]. rejections [AAG09]. Related [AAE^{+}25, BFGT16, BDW19, CFRY22, GWZ21, CKP12]. relations [BHMS11]. Relationship [KR19]. relative [BF07, BM08]. Relatives [Nav25]. Relaxation [Li17, VZ21, ABD^{+}08, GGG10]. relaxed [DM09]. Relay [EFM^{+}16]. Reliable [HPMO23]. Rendezvous [FP13, TSZ14, KKM11]. rent [BKLP07]. rent-or-buy [BKLP07]. Reordering [AER15]. Repairmen [HKKK16, FHR07]. Repeats [CCKN19]. repetitions [AAG09]. Repetitive [Nav25]. Replacement [GW20, RZ12, SP24, WY13, EPR10]. Reporting [EP16]. Reports [DP14]. Representation [BS18]. Representations [Gab16, FMMN07]. Representative [FLPS17]. Representing [BN15]. Required [PT16]. Rerouting [IIK^{+}25]. Resampling [Har21]. reservation [BM08]. residents [KMMP07]. Resilience [BHW20]. Resilient [FGI09]. Resistance [CLS^{+}22]. Resolution [GM12, ND25]. Resource [JMR19, CCKR11, GN14]. respect [CGK^{+}11]. response [FFM12, PUW08]. Restore [CMR18, FTK20]. Restricted [AKS17, MPQS20, Neu22, PS16, BNLT07]. restriction [GGG10]. restrictions [AMS06]. Results [BG21, EW20, FMM^{+}25, GRSW16, CPR^{+}11, GHKS06, GHKS13, HIMY07, KNS^{+}07]. Retractions [FGZ21]. Retrieval [SSTV23]. Retroactive [DIL07]. Reversal [CBFWW15]. Revisited [ABH^{+}18, CK07, HPR14, NS09]. Ride [GNR15, GHNR10a, GHNR10b]. right [DMM^{+}12]. Robust [ARS^{+}14, GMP23a, GNR16, HS06]. roommates [CF05]. Rooted [CLNV14, KL06, Saw06]. roots [Lau06]. Rounding [AZBG^{+}22, FGG^{+}25]. Roundtrip [RTZ08]. Route [CMVZ16]. Routing [AZ07, ADPP07, CGMZ16, CBFWW15, GMT11, JS23, KRX16, MZ23, AGM^{+}08, CSTW12, HKRL07, KP08, RTZ08]. Rule [HJ15]. Rumor [DFS14].

s [BRFF^{+}12, CRV11]. S-T [CRV11]. Safe [CRTZ24]. Salesman [AKS21, DT16, BHKK12]. Same [SSA25]. samples [EMS10]. Sampling [BT24, BGGS24, BCP24, BCSV25, Bre23, CHGG^{+}17, Coh18, WY25, BCEG07, FSP08, GKK10, MPV10, PT11]. Santa [AFS12]. SAT [CDL^{+}16]. Satisfaction [AZBG^{+}22, GRS17, CMM09, SS09]. savings [ISG07]. Scalable [BKN14, GHM^{+}24, IM12]. Scalably [EP12]. Scale [BFF^{+}22, KRX16].

Scale-Free [BFF^{+}22, KRX16]. Scaling [BCP13, DPS18, AA14]. Scattering [Fil24]. schedule [Wil10]. Scheduling [AFH^{+}16, BKN14, HJ15, JMR19, KMPS16, LMMW16, SZ20, AZ07, BCD12, BNLT07, BNGK^{+}09, BCMSM12, CCL^{+}09, CEGK11, CKK10b, CRR09, DJP^{+}12, EP12, EV06a, EK06, GHKS06, GHKS08, GP08, Hal12, IM12, JZ06, JLSS12, GHKS13]. Scheme [BKM15, BKM09a, BKM09b, CSTW12, LDX09, SS08a]. Schemes [BKK^{+}19, FGL^{+}20, FKRS19, HW19, JMR19, JS23, KRX16, IMM08, Kar08, KP08, Kor10, NHK08]. SDD [KLP16b]. Search [BCI^{+}23, Cha13, CGMY22, ER17, GMW20, PS16, STK16, JS07]. Searching [AAHP^{+}16, BFKR25, ACY12, FG08]. seat [BM08]. secretary [BHZ13]. secrets [AGKS07]. Secure [FIM^{+}06, PR12]. Seeds [KKR^{+}20]. Segmentation [ADV^{+}16]. Select [BFV22]. Selection [AFHN16, AFK^{+}15, CMR18, EHR16, GIN^{+}17, Cha10a, Han07]. Self [Elm17, JCCW22, ST25]. Self-Adjusting [Elm17, ST25]. Self-Improving [JCCW22]. Selfish [BHS14, ADPP07]. Semi [AAE^{+}25, ER16, FLN14, KR16, PS19]. Semi-Algebraic [AAE^{+}25]. Semi-Matching [FLN14]. Semi-matchings [KR16]. Semi-Streaming [ER16, PS19]. Semicontinuous [AFK^{+}15]. Semidefinite [DHS16]. sender [EK06]. Sensitive [CL22, Pag18]. Sensitivity [BBP23, GW20, WY13]. sensor [BMSV^{+}09, FCFM09]. separability [AAK06]. Separate [GS17]. separating [Cab10]. Separations [CKP^{+}19, JMR22]. separator [RW09]. Separators [LT24, LPS^{+}20, MOR13]. sequence [CFLM07]. Sequences [BN15, SSA25, TSZ14, FMMN07, FKW11, GGN06, Kau07, MN08]. Sequential [KR18]. Server [BEJK19, BEKN23, CKL21a, CV20, FN10]. Service [AGGP21, RKH20, BCKV06, MV08]. sessions [And09]. Set [ACGP16, ACS^{+}22, ALM^{+}19, ASS19, BHPR^{+}20, CEGS11, CCHM15, DHK16, ER16, FGG^{+}25, FLST18, FLL^{+}19, GKPP22, HPJ21, Har19, LRS18, LM19, LMM^{+}21, AAG09, BCN12, Cha10b, CCHP12, ELR^{+}08, GKLT09, PRS12, RW10a, Tho10a, Tho10b]. set-up [ELR^{+}08]. SETH [ABHS22, Bon22]. SETH-based [ABHS22]. SETH-hard [Bon22]. Sets [BHPR19, BDH^{+}20, FMM^{+}25, GS17, GNR16, JN24, LN22, LPS^{+}20, AMS06, BLW09, Epp09a, FGPS08, RSS06]. settings [CCHP12]. sex [IMY10]. sex-equal [IMY10]. Share [AMNS17]. shared [AKM08, AC10]. shared-memory [AKM08, AC10]. Shares [CF20]. Sharing [BKLP07, IMM08]. Shaving [CZ24]. Shivers [Jug24]. shooting [GK09]. shop [GHKS06, GHKS13]. Short [CP25, KSS09]. Shortcut [CD24]. Shortcuts [AFK^{+}15]. Shortest [BCHR20, BFF^{+}22, CW15, CR18, CSS21, DS11, FGK^{+}24a, KMW10, Cab10, Cha12, DI06, Elk05, HSB07, HMS07, KK06, MTZ10, MZ12, RS11a, RZ12]. shredders [Heg06]. SIAM [BKN^{+}25, ELMS25, BGN^{+}18, DMT24, HK22, KADK22]. side [EV06b]. Sided [Kav24]. Sigma [SW20a]. signatures [Ruž09]. Silent [DP14]. Simon [AMR09]. Simple [CDD^{+}15, CJL17, CGMY22, FGK^{+}16, GWZ21, NS16, AMM07, CW10, HMS07, RZ12]. Simpler [CD17, Kol25, LBSZ21, MM09, Rod08]. Simplex [DS19]. Simplicial [BS18]. Simplified [KN16]. Simultaneous [AGG^{+}09, BR16]. Single [BCHR20, CNP^{+}22, HJ15, KLP^{+}16a]. Single-Exponential [KLP^{+}16a]. Single-Machine [HJ15]. Single-Source [BCHR20]. SINR [AK18, HT21, KLPP23]. Size [EP16, EHL^{+}18, HS18, LN22, NRS18, RS11b]. Sizes [MPQS20]. sizing [ELR^{+}08]. Skeletons [CMV16]. Sketches [CD17]. Skew [RS17]. Skew-Symmetric [RS17]. Skip [AS07]. Slack [AEP18]. sleep [AA14]. Small [BCKM20, BFG^{+}16, Bre23, CGH17, DGG^{+}20, GJNW23, LPS^{+}20, MTK^{+}19, Neu24, And09, MOR13, PT11]. Smaller [ACHKP21]. Smith [HJ15]. Smolensky [CW21]. Smoothed [BCC21, CDL^{+}25, DMM^{+}12, ERV16, ER17, AK12a]. Social [GKK23]. SODA [BKN^{+}25, BGN^{+}18, DMT24, ELMS25, HK22, KADK22, Epp07, GBFC07, Gab09, HT10, MWY19, Mat10]. SODA’11 [CRR13]. SODA’12 [RRSW16]. SODA’18 [LPW20]. Solutions [HJT17, BST08]. solvable [ASS08]. Solvers [JS25]. Solving [CNP^{+}22, GM14, SW20a, DS08]. Some [Cha20, CDP19, GMS19, GRSW16, AR09, CGNS08, FP10, GKLT09, HSB07]. Sort [CFRY22, Jug24, NZC11]. Sorting [AFHN16, CMR18, DEK21, FTK20, Jug24, FZ07, FG08, GS09]. Source [BCHR20, AGG^{+}09]. Space [BCKM20, BKMV20, BFG^{+}16, CDP21, EHR16, GS17, LBSZ21, WY18, BG11, Cha10a, DFR09, EMS10, GLS10, KMW10, PT11]. Space-Constrained [EHR16]. Space-optimal [BKMV20]. Spaces [HPMO23, LT24]. Spanner [BFH21, DKR16]. Spanners [BT25, BW21, CWN18, CDKL20, ES16, EN19, HPMO23, LT24, Sol13, BKMP10, BKS12, Elk11, GLNS08, Pet09, RTZ08]. Spanning [ACE^{+}20, ABD^{+}25, BD11, FGLS19, GT16a, GT16b, HKN17, PRS20, CS08, DV10, JR05, PR08, ZO08]. Sparse [BFG^{+}16, BHPR19, Bon22, CRSS22, CI17, DHS16, EN19, Fil24, FTK20, GIKW19, GLPS17, HT21, MTK^{+}19, MWsZ23, OSSW20, PP16, Pre21, RT14, Sol13, BSSX20, Cla10, Elk11, Epp09a, MR09, YZ05]. Sparsest [CKM^{+}24, Kol25, KW16b, CGR08]. Sparsification [CDP21, KLP16b, PZ24, PPSV18]. Spatial [ADGH21, Wal25]. Special [BKN^{+}25, BGN^{+}18, CHL^{+}20, DMT24, ELMS25, HK22, KADK22, LPW20, MWY19, RRSW16, Epp07, GBFC07, Gab09, HT10, Mat10, CRR13]. specific [GMT11]. Spectral [BGGS24, FGYZ22, KLP16b]. Speed [BCP13, JN24, AA14, FFM12]. Speed-Ups [JN24]. Speeding [GKM16]. speedup [BGLZ09, EP12]. Speedups [Cha20]. Spider [Fuk17]. split [Iba08]. splittable [EV06a, GMT11]. Splitting [CCG^{+}24]. Spreading [DFS14]. Square [FGZ21]. Square-free [FGZ21]. Squarepants [Epp09b]. Squares [AARA23, HKP^{+}18, Kol08]. SRPT [TM08]. Stabbing [Tao14, ELR^{+}08]. Stability [CCDL16]. stable [CF05, HIMY07, IMY10, KMMP07]. Stackelberg [Swa12]. Stars [HMVW19]. state [AA14]. Static [FF23, NS14, ALLS07]. station [KMM11]. stationary [CRV11]. Statistics [Coh18]. steepest [AKR12]. Steiner [BKM09a, BKM09b, BKM15, CJJ20, CEGS11, CS07, CLNV14, CFM21, CDKL20, CJKV24, DHK14, DKN17, EKS05, FL25, FLK^{+}20, FM25, GMP23a, HKKN12, KBNvL20, LL25, PPSV18]. Steinitz [EW20]. Sticky [AZBG^{+}22]. Stochastic [ADKS25, BJKK18, DHK16, KR24]. Stock [NRS18]. Storage [BFCF^{+}17]. Straight [CMV16]. strategies [MPV10, Swa12]. Stream [Coh18, CCM10]. Streaming [AN16, Bła20, BG14, CJKV24, Elk11, ER16, EHL^{+}18, FF23, JW23, JW13, KR16, DFR09, FMS^{+}10, PS19]. Streams [BDW19, CCMT14, BCEG07, HS09]. Stretch [EP16, AGM^{+}08]. Strict [BLT25]. strike [LY08]. String [BCKM20, BGS23, BG14, EMTG23, FV16, JN24, BFG09, CKS05, CFLM07, CM07]. String-Matching [BG14]. Strings [Gaw13, BHMS11]. Strip [KLM^{+}25]. Strong [BG21, CC25, GHH22, ND25, AC10, HKM^{+}12]. Strong-Connectivity [CC25]. Strongly [BGMW20, KMMP07, Lac13, TSZ14]. Structure [BCSV20, BLS08, BNS19, Gab17, MRYY24, CFLM07]. Structures [BFV22, FF23, GIKW19, GIN^{+}17, HMZ16, PP16, PP18, RT22, Wal25, AKS08, DIL07, GKS^{+}11]. Study [BBB20]. Subconstant [JW13]. Subcubic [AGW23, BGMW20]. subdivisions [CDI^{+}12, GK09]. Subexponential [BFPP18, FLK^{+}20]. Subgraph [HVV19, ST08]. Subgraphs [BKK17, BCP24, Bre23, Neu24, And10, HS06, PS10, VWY10]. Subject [MS17]. Sublinear [ELMR21, GMV09, HKN17, LT24, EMS10, RS11a]. Sublinear-Time [HKN17]. Submatrices [HL13]. Submatrix [GMW20, KMNS17]. Submodular [BHZ13, BJLY17, BF18, BFS19, CHJ^{+}18, DHK16, DFL^{+}25, Fel17, FGK^{+}24a, GKK23, INV16, WZ16]. Subquadratic [Cab19, FLL^{+}19, KKK18]. Subsequence [BCAD23]. Subset [ABHS22, CCHM15, KX19, LRS18]. Subspace [WW22]. Substring [JN24, Pre21]. Subtree [ABH^{+}18, Epp09b]. Succinct [BHMS11, BCH^{+}12, GRR06, HMS12, NS14, RRS07]. Suffix [FTK20, OOB24, FGGV06, RNO11]. Sum [ABHS22, BT24, BJLY17, Cab19, HKS11, HKP^{+}18, KX19, Epp09b]. Sums [DHS16, RRS07, Vig14]. supporting [BCH^{+}12]. Surface [CJL17]. Surfaces [CdM22, RST14]. Survivable [BK16]. Survivable-Network [BK16]. Swap [BPL18]. Swap-Insert [BPL18]. switches [AKM08, AR06, CJST07]. Symmetric [Fel17, NZ24, RS17, FMS^{+}10]. Symmetry [GSV20]. Symposium [BKN^{+}25, BGN^{+}18, DMT24, ELMS25, HK22, KADK22]. Synchronized [BFR23]. Synchronizing [CP25, JN24]. System [JS25]. Systems [BCP24, DJO^{+}25, BS10].

T [CRV11]. Tabulating [Sha16]. Tandem [CCKN19]. Tangents [AW19]. tardiness [KKW12]. Tardos [Har16, HS17]. Tarski [FPS22]. tasks [EV06a, HL06b, JZ06]. Tau [SW20a]. Taxes [CKK10a]. Techniques [RT22, GGG10]. Telling [CDD^{+}15]. Terminals [KBNvL20]. Terrain [AFK^{+}18, AAY10]. Terrains [DHPR16]. Test [GNSW20]. Testers [BMR22, CDJS17]. Testing [ADF^{+}15, AK12b, BCSV25, CDJS17, CS08, ELR20, Epp09c, LCS^{+}19, Neu24, RT14, EMS10, FLM^{+}12]. Text [BN14, BFG^{+}16, Nav25, Pre21, ALLS07, CHLS07, FMMN07, MN08]. texts [BFG09]. Their [ACE^{+}20, Fil24, KMNS17, KMSY25, AG10]. Theorem [HHM^{+}18, LS24]. Theoretical [BBB20]. Theory [CRTZ24, HJT17]. thick [KT08]. Thin [BKK17, KT08]. Three [AAE^{+}25, DKT11]. Three-coloring [DKT11]. Threshold [Wal25]. Thresholds [Wal23]. Throughput [BNGK^{+}09, EMS23, CCL^{+}09, GR10]. Tight [AFCK25, BLL^{+}24, BODD^{+}20, BHD^{+}21, BD11, CHGG^{+}17, CGH17, CV07, DJP^{+}12, FN20, FMR24, FMM^{+}25, KLM^{+}25, KPR16, Soc16, WW22, WY18, CDEM10, KBNvL20, MSS11]. Tightening [CdM22]. Tiling [AARA23]. Time [AARA23, AW19, ALL^{+}25, ATG^{+}14, BKMV20, BGK^{+}22, BKK17, Bon22, BSWN15, BKM15, BG14, BGMW20, BCAD23, CRTZ24, CDJ23, CBFWW15, CCKN19, CEK^{+}21, CRR09, CNP^{+}22, DFM23, DHM^{+}14a, FLS^{+}18, Gab25, GLPS17, GMP17, GS18, GHT18, HSS22, HKN17, KMTS23, iKMNZ25, KKR^{+}20, KX19, Laf23, LR15, LN22, LRS18, PRS20, RS17, RKH20, SS18, SHHA16, WY16, AF07, AHRT05, AZ07, AK12b, BD07, BCD12, BNGK^{+}09, BS06a, BCKM20, BPGN09, BKMSS11, BCMSM12, BLS08, Cha10a, Cha12, CDL^{+}25, DKT11, EPR10, EDKM07, GKLT09, GN14, GLP06, GKPP22, GT08, HP22, HL06b, IM12, JS07, KMMP07, KMW10, LDX09, MOR13, NZC11, RW09, TM08, VH05]. Time- [BKMV20, PRS20]. Time-dependent [CRR09]. Time-Energy [CDJ23]. time-space [Cha10a]. time-varying [AZ07]. Times [BH19, GPSS15, JMR19]. Tiny [BCFC^{+}25, Sol13]. Tolerant [BCHR20, BMR22, HHL^{+}16, PP16, PP18, SS08b]. tolls [Swa12]. Tool [KLPP23]. Tools [AK18]. Top [GIN^{+}17, RT22, AHTL05]. Top- [GIN^{+}17, RT22]. Topological [CFRY22, FLST18, GST23, Neu24, AFM08, KB06, HKM^{+}12]. Tossing [BBHT17]. total [BGLZ09, GLP06, KKW12]. Tournaments [LMM^{+}21, CFR10]. Tours [CKS19, GGG10]. Tower [DS08, KR24, BS06b]. Tower-On [KR24]. Trace [CDL^{+}25]. Tracking [Bła20, YZ12]. Tractability [DHM14b, FGK^{+}24b, GRS17, CGK^{+}11]. Tractable [CM15, CCHM15, GST23, HIMŽ19, RSS06]. Tradeoff [CCDL16, RS09]. Tradeoffs [CDJ23, GMP17]. Trading [DFR09, BCFN07]. Traffic [DP14, ADPP07]. train [Wil10]. Trajectories [ADV^{+}16]. Transform [CI17, AL13, NZC11]. Transforms [BHK^{+}16, MRR06, JW13]. transitive [Rod08]. Translation [BBC^{+}25, BKN21, AHPSW10]. transpositions [FZ07]. Transversal [ALL^{+}25, KW14]. Transversals [GHH22]. Traveling [AKS21, DT16, BHKK12, FHR07]. Treasure [BDLP23, TSZ14]. Tree [Adj19, AGP^{+}11, BGMW20, CJJ20, CKS19, DHK14, FV16, FLK^{+}20, FM25, GMP23a, GT16a, GT16b, KBNvL20, LL25, ADHY08, BG11, BKM09a, BKM09b, CMS07, DMRW09, Epp09b, Fuj12, GNSW20, HMS12, JR05, PR08, BD11]. Trees [ACE^{+}20, ABZ19, AFN^{+}18, ABD^{+}25, BEJK19, BBT12, BFKR21, BFKR25, BCI^{+}23, CMM25, CZ24, CGMY22, FP13, GT16a, GT16b, GMP17, GMS19, HST15, HKN17, MZ23, NS14, PRS20, STK16, TLT21, ABD^{+}08, AHTL05, BHMS11, BRFF^{+}12, CPR^{+}11, CS07, CS08, DP06, DM09, FSP08, GRR06, GKS^{+}11, HS06, HMS12, KL06, MSS11, RRS07, RNO11, Saw06, ZO08]. Treewidth [Cab22, CNP^{+}22, FMR24, FMM^{+}25, FLS^{+}18, GJNW23, JS23, LMS18, BFK^{+}12, CKM^{+}24, MOR13]. triangle [Blä08, DKT11]. triangle-free [DKT11]. triangulation [CDI^{+}12]. triangulations [OGGW10]. tries [JS07]. trim [Fuj12]. True [BLL^{+}24]. Truncation [LMPS18]. Truthful [AGG10]. TSP [AFHS20, BKMV20, BHD^{+}21, Blä08, CJ18, CJJ20, CKS19, ERV16, GMP23a, Man12, dBBJW21]. Tutte [DHM^{+}14a]. Tverberg [ASW08]. Two [AW19, Adj19, AAGH25, BK16, CZ24, CGMY22, CY25, CHLT14, FP10, HS18, HW19, KK16, KR16, MW24, WY18, dBBJW21, DS11, GP08]. Two-Dimensional [CHLT14, HW19, MW24, WY18, CZ24]. Two-Edge [BK16]. two-face [DS11]. Two-Handed [AAGH25]. Two-Party [KR16]. Two-phase [FP10]. Two-Way [CY25, CGMY22]. Type [HPST19, Kra14, KW16b, CGR08].

UET [HL06b]. UGC [GRSW16]. ultrametrics [ABD^{+}08]. Ultrasparse [JS25]. Ultrasparsifiers [JS25]. un- [GMT11]. Uncapacitated [KS16]. Uncertain [BFL^{+}23, ACY12, MNS12]. Uncertainty [AAHP^{+}16, GM12, GNR16]. uncrossable [Nut12]. Underlying [BFF^{+}22]. Undirected [Bon22, KKPW24, Cha12, CEGS11, MZ12]. Unequal [CF20]. ungapped [FBV09]. Unified [CJJ20]. Uniform [BEKN23, CGK20, DKN17, GJLS17, Li17, Ruž08, GKK10, GLP06]. Uniformly [OSSW20]. Union [ACHM22, ATG^{+}14, AAY10]. Union-Find [ATG^{+}14]. unique [BLW09]. Unit [ACHM22, BLL^{+}24, PSZ24, EV10]. Unit-Disk [BLL^{+}24]. Universal [ABCH25, GMP23b, TSZ14, RW10a]. Unknown [BGH19]. Unless [BGMW20]. Unrelated [EMS23, CKK10b]. Unrestricted [AL13]. Unsplittable [AGLW18, ADPP07, AGG10, BFKS14, BH12]. Untrusted [ACE^{+}23]. Unweighted [Bon22, BDLP23, BS06a, Cha12, RZ12]. Update [BGK^{+}22, GHT18, AGvS13]. Updates [BJLY17]. Updating [DM09]. Upper [BN15, CDKL20, HCT^{+}11, MSS11]. Ups [JN24]. Useful [KLPP23]. Using [CKS19, EW20, GKM16, LNR^{+}14, MP22, KS08, PR08]. utilities [GN14]. utilizes [TM08].

Validity [BF24]. Valuation [CD17]. Valuations [GKK23]. value [GGM10]. valued [MWsZ23]. variable [BB08]. variable-length [BB08]. Variant [HL13]. Variants [BDH^{+}20, dBBJW21, JR05]. Variations [KR19]. varying [AZ07]. VCSPs [HIMŽ19, VZ21]. Vector [SWYZ21]. Vectors [CW21]. Vegas [Wei22]. Vehicle [GNR15, JS23, MZ23]. Verification [KMZ18]. version [BNLT07]. Vertex [ALM^{+}19, ALM^{+}23, CHGG^{+}17, CCHM15, HTWZ19, LN22, LRS18, LMM^{+}21, Wan22, DS11, GKP08, HL06a, Kar09, Kor10, RSS06, Tho10a, Tho10b, VB08, Vis08]. vertex-disjoint [DS11].

Vertex-Weighted [HTWZ19]. vertical [GK09]. Vertices [CMT22]. Very [EN19]. Via [KLP^{+}16a, APF^{+}10, AGKS07, ANFS17, AK18, AFK^{+}15, AKR12, BT24, BGGS24, BKK17, BHPR19, CMM25, CKM^{+}24, CCW18, CY25, EPR13, FGPS08, FPZ23, GGI^{+}21, GKK10, GGG10, GKP08, GM14, HH17, HKS11, HIMŽ19, HHM^{+}18, KW14, MOR13, Nut12, PT11, RS17, WY13]. video [EK06]. video-on-demand [EK06]. visits [BKMV20]. Voronoi [DHPR16, MP22]. Voting [ADGH21]. vs [Bon22, GMP17, NZ24, RS11b].

wait [EP05]. Walks [CMA^{+}19, DGG^{+}20, HSS22, SS18]. Walsh [CI17]. Warping [GS18]. Waste [SHHA16]. Way [CY25, CGMY22]. Weak [AFT19, FCFM09]. weakly [HSS07]. Weight [GKPP22, PS19, WY25, KM12]. Weighted [Adj19, ALM^{+}20, BT24, CCW18, CV20, DHK14, DPS18, GHH22, HTWZ19, JPS22, Mes14, AKL10, BD07, CEV21, CFR10, GR10, GKP08, KKW12, VWY10, VH05, Wil10]. Weighted- [ALM^{+}20]. Weights [DKN17]. Welfare [GKK23]. Well [FV19, BCM11]. Well-Behaved [FV19]. Wideness [KRS19]. Width [AN16, CCG^{+}24, CRZ20, FL25, GHK^{+}24, MTK^{+}19, CDP19, FGL^{+}19, GNSW20, Mar10, Oum08]. Williams [IIK^{+}23]. Windows [BNLT07]. Windrose [ADD^{+}18]. wins [CFR10]. Wireless [Hal12, KLPP23, KMPS16, AZ07, BKMSS11, EF12, NHK08, PR12]. Without [STK16, KR18, Pag18]. Witness [HH17]. witnesses [ALLT11]. Wolfe [Cla10]. Word [Cha13]. Words [CP25, CCKN19, BCN12, KSS09]. Workspace [AW19]. World [DGG^{+}20]. worst [ADHY08, BF07, BM08, CV07]. worst-case [ADHY08, CV07]. Writing [KS08]. Writing-all [KS08].

Yao [BGLZ09].

zero [Kau07]. Zeros [GLLZ21]. Zeta [BHK^{+}16]. Zip [TLT21]. Ziv [BFG09].

References

Albers:2014:RIN

[AA14] Susanne Albers and Antonios Antoniadis. Race to idle: New algorithms for speed scaling with a sleep state. ACM Transactions on Algorithms, 10(2):9:1–9:??, February 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Alon:2006:GAO

[AAA^{+}06] Noga Alon, Baruch Awerbuch, Yossi Azar, Niv Buchbinder, and Joseph (Seffi) Naor. A general approach to online network optimization problems. ACM Transactions on Algorithms, 2(4):

640–660, October 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Arya:2022:OBC

[AAdFM22] Rahul Arya, Sunil Arya, Guilherme D. da Fonseca, and David Mount. Optimal bound on the combinatorial complexity of approximating polytopes. ACM Transactions on Algorithms, 18 (4):35:1–35:??, October 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3559106.

Agarwal:2025:IQF

[AAE^{+}25] Pankaj K. Agarwal, Boris Aronov, Esther Ezra, Matthew J. Katz, and Micha Sharir. Intersection queries for flat semi-algebraic objects in three dimensions and related problems. ACM Transactions on Algorithms, 21(3):25:1–25:59, July 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Alon:2009:ACM

[AAG09] Noga Alon, Yossi Azar, and Shai Gutner. Admission control to minimize rejections and online set cover with repetitions. ACM Transactions on Algorithms, 6(1):11:1–11:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Agarwal:2025:THP

[AAGH25] Pankaj K. Agarwal, Boris Aronov, Tzvika Geft, and Dan Halperin. On two-handed planar assembly partitioning with connectivity constraints. ACM Transactions on Algorithms, 21(2): 19:1–19:??, April 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Agarwal:2016:NNS

[AAHP^{+}16] Pankaj K. Agarwal, Boris Aronov, Sariel Har-Peled, Jeff M. Phillips, Ke Yi, and Wuzhou Zhang. Nearest-neighbor searching under uncertainty II. ACM Transactions on Algorithms, 13 (1):3:1–3:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Agarwal:2006:EAB

[AAK06] Pankaj K. Agarwal, Boris Aronov, and Vladlen Koltun. Efficient algorithms for bichromatic separability. ACM Transactions on Algorithms, 2(2):209–227, April 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Aamand:2023:TSP

[AARA23] Anders Aamand, Mikkel Abrahamsen, Peter M. R. Rasmussen, and Thomas D. Ahle. Tiling with squares and packing dominos in polynomial time. ACM Transactions on Algorithms, 19 (3):30:1–30:??, July 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3597932.

Agarwal:2010:EBU

[AAY10] Pankaj K. Agarwal, Lars Arge, and Ke Yi. I/O-efficient batched union–find and its applications to terrain analysis. ACM Transactions on Algorithms, 7(1):11:1–11:??, November 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Austrin:2025:OIU

[ABCH25] Per Austrin, Jonah Brown-Cohen, and Johan Håstad. Optimal inapproximability with universal factor graphs. ACM Transactions on Algorithms, 21(3):32:1–32:39, July 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Alon:2008:OEM

[ABD^{+}08] Noga Alon, Mihai Bãdoiu, Erik D. Demaine, Martin Farach-Colton, Mohammadtaghi Hajiaghayi, and Anastasios Sidiropoulos. Ordinal embeddings of minimum relaxation: General properties, trees, and ultrametrics. ACM Transactions on Algorithms, 4(4):46:1–46:??, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Akitaya:2025:MPB

[ABD^{+}25] Hugo A. Akitaya, Ahmad Biniaz, Erik D. Demaine, Linda Kleist, Frederick Stock, and Csaba D. Tóth. Minimum plane bichromatic spanning trees. ACM Transactions on Algorithms, 21(4):48:1– 48:14, October 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ahmadian:2018:AAM

[ABF^{+}18] Sara Ahmadian, Babak Behsaz, Zachary Friggstad, Amin Jorati, Mohammad R. Salavatipour, and Chaitanya Swamy. Approximation algorithms for minimum-load k-facility location. ACM Transactions on Algorithms, 14(2):16:1–16:??, June 2018. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Austrin:2016:BBB

[ABG16] Per Austrin, Siavosh Benabbas, and Konstantinos Georgiou. Better balance by being biased: a 0.8776-approximation for max bisection. ACM Transactions on Algorithms, 13(1):2:1–2:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Abboud:2018:SIR

[ABH^{+}18] Amir Abboud, Arturs Backurs, Thomas Dueholm Hansen, Virginia Vassilevska Williams, and Or Zamir. Subtree isomorphism revisited. ACM Transactions on Algorithms, 14(3):27:1–27:??, July 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3093239.

Abboud:2022:SBL

[ABHS22] Amir Abboud, Karl Bringmann, Danny Hermelin, and Dvir Shabtay. SETH-based lower bounds for subset sum and bicriteria path. ACM Transactions on Algorithms, 18(1):6:1–6:22, January 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3450524.

Ackermann:2010:CMN

[ABS10] Marcel R. Ackermann, Johannes Blömer, and Christian Sohler. Clustering for metric and nonmetric distance measures. ACM Transactions on Algorithms, 6(4):59:1–59:??, August 2010. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Adjiashvili:2019:FTB

[ABZ19] David Adjiashvili, Andrea Baggio, and Rico Zenklusen. Firefighting on trees beyond integrality gaps. ACM Transactions on Algorithms, 15(2):1–33, May 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3173046.

Aspnes:2010:ASM

[AC10] James Aspnes and Keren Censor. Approximate shared-memory counting despite a strong adversary. ACM Transactions on Algorithms, 6(2):25:1–25:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Abraham:2020:RST

[ACE^{+}20] Ittai Abraham, Shiri Chechik, Michael Elkin, Arnold Filtser, and Ofer Neiman. Ramsey spanning trees and their applications. ACM Transactions on Algorithms, 16(2):19:1–19:21, April 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3371039.

Antoniadis:2023:OMA

[ACE^{+}23] Antonios Antoniadis, Christian Coester, Marek Eliás, Adam Polak, and Bertrand Simon. Online metric algorithms with untrusted predictions. ACM Transactions on Algorithms, 19 (2):19:1–19:??, April 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3582689.

Adamaszek:2019:LKC

[ACER19] Anna Adamaszek, Artur Czumaj, Matthias Englert, and Harald Räcke. An O(log k)-competitive algorithm for generalized caching. ACM Transactions on Algorithms, 15(1):1–18, January 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3280826.

Abraham:2016:FSD

[ACGP16] Ittai Abraham, Shiri Chechik, Cyril Gavoille, and David Peleg. Forbidden-set distance labels for graphs of bounded doubling dimension. ACM Transactions on Algorithms, 12(2):22:1–22:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Abboud:2021:SCH

[ACHKP21] Amir Abboud, Keren Censor-Hillel, Seri Khoury, and Ami Paz. Smaller cuts, higher lower bounds. ACM Transactions on Algorithms, 17(4):30:1–30:40, October 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3469834.

Agarwal:2022:MUU

[ACHM22] Pankaj K. Agarwal, Ravid Cohen, Dan Halperin, and Wolfgang Mulzer. Maintaining the union of unit discs under insertions with near-optimal overhead. ACM Transactions on Algorithms, 18(3):26:1–26:??, July 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3527614.

Agarwal:2022:DGS

[ACS^{+}22] Pankaj Agarwal, Hsien-Chih Chang, Subhash Suri, Allen Xiao, and Jie Xue. Dynamic geometric set cover and hitting set. ACM Transactions on Algorithms, 18(4):40:1–40:??, October 2022. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3551639.

Agarwal:2012:RSU

[ACY12] Pankaj K. Agarwal, Siu-Wing Cheng, and Ke Yi. Range searching on uncertain data. ACM Transactions on Algorithms, 8(4):43:1– 43:??, September 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Aumuller:2016:OPD

[AD16] Martin Aumüller and Martin Dietzfelbinger. Optimal partitioning for dual-pivot Quicksort. ACM Transactions on Algorithms, 12 (2):18:1–18:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Angelini:2018:WPE

[ADD^{+}18] Patrizio Angelini, Giordano Da Lozzo, Giuseppe Di Battista, Valentino Di Donato, Philipp Kindermann, Günter Rote, and Ignaz Rutter. Windrose planarity: Embedding graphs with direction-constrained edges. ACM Transactions on Algorithms, 14(4):1–24, October 2018. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3239561.

Angelini:2015:TPP

[ADF^{+}15] Patrizio Angelini, Giuseppe Di Battista, Fabrizio Frati, Vít Jelínek, Jan Kratochvíl, Maurizio Patrignani, and Ignaz Rutter. Testing planarity of partially embedded graphs. ACM Transactions on Algorithms, 11(4):32:1–32:??, June 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Aronov:2021:PPS

[ADGH21] Boris Aronov, Mark De Berg, Joachim Gudmundsson, and Michael Horton. On β-plurality points in spatial voting games. ACM Transactions on Algorithms, 17(3):24:1–24:21, August 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3459097.

Arge:2008:PRT

[ADHY08] Lars Arge, Mark De Berg, Herman Haverkort, and Ke Yi. The priority R-tree: a practically efficient and worst-case optimal Rtree. ACM Transactions on Algorithms, 4(1):9:1–9:??, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Adjiashvili:2019:BAF

[Adj19] David Adjiashvili. Beating approximation factor two for weighted tree augmentation with bounded costs. ACM Transactions on Algorithms, 15(2):1–26, May 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3182395.

Aumuller:2016:HGM

[ADK16] Martin Aumüller, Martin Dietzfelbinger, and Pascal Klaue. How good is multi-pivot Quicksort? ACM Transactions on Algorithms, 13(1):8:1–8:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ayyadevara:2025:NOA

[ADKS25] Nikhil Ayyadevara, Rajni Dabas, Arindam Khan, and K. V. N. Sreenivas. Near-optimal algorithms for stochastic online bin packing. ACM Transactions on Algorithms, 21(2):24:1–24:??, April 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Auletta:2007:RSU

[ADPP07] Vincenzo Auletta, Roberto De Prisco, Paolo Penna, and Giuseppe Persiano. Routing selfish unsplittable traffic. ACM Transactions on Algorithms, 3(4):52:1–52:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Aronov:2016:STN

[ADV^{+}16] Boris Aronov, Anne Driemel, Marc Van Kreveld, Maarten Löffler, and Frank Staals. Segmentation of trajectories on nonmonotone criteria. ACM Transactions on Algorithms, 12(2):26:1–26:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Amir:2012:CDC

[AEL^{+}12] Amihood Amir, Estrella Eisenberg, Avivit Levy, Ely Porat, and Natalie Shapira. Cycle detection and correction. ACM Transactions on Algorithms, 9(1):13:1–13:??, December 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Antoniadis:2024:BBC

[AEMV24] Antonios Antoniadis, Matthias Englert, Nicolaos Matsakis, and Pavel Veselý. Breaking the barrier of 2 for the competitiveness of longest queue drop. ACM Transactions on Algorithms, 20 (4):38:1–38:??, October 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3676887.

Anagnostopoulos:2018:RES

[AEP18] Evangelos Anagnostopoulos, Ioannis Z. Emiris, and Ioannis Psarros. Randomized embeddings with slack and high-dimensional approximate nearest neighbor. ACM Transactions on Algorithms, 14 (2):18:1–18:??, June 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Avigdor-Elgrabli:2015:ICA

[AER15] Noa Avigdor-Elgrabli and Yuval Rabani. An improved competitive algorithm for reordering buffer management. ACM Transactions on Algorithms, 11(4):35:1–35:??, June 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Albers:2007:EEA

[AF07] Susanne Albers and Hiroshi Fujiwara. Energy-efficient algorithms for flow time minimization. ACM Transactions on Algorithms, 3 (4):49:1–49:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Assadi:2025:TBM

[AFCK25] Sepehr Assadi, Martín Farach-Colton, and William Kuszmaul. Tight bounds for monotone minimal perfect hashing. ACM Transactions on Algorithms, 21(4):36:1–36:23, October 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Adany:2016:ANG

[AFH^{+}16] Ron Adany, Moran Feldman, Elad Haramaty, Rohit Khandekar, Baruch Schieber, Roy Schwartz, Hadas Shachnai, and Tami Tamir. All-or-nothing generalized assignment with application to scheduling advertising campaigns. ACM Transactions on Algorithms, 12(3):38:1–38:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ajtai:2016:SSI

[AFHN16] Miklós Ajtai, Vitaly Feldman, Avinatan Hassidim, and Jelani Nelson. Sorting and selection with imprecise comparisons. ACM Transactions on Algorithms, 12(2):19:1–19:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Antoniadis:2020:PET

[AFHS20] Antonios Antoniadis, Krzysztof Fleszar, Ruben Hoeksma, and Kevin Schewior. A PTAS for Euclidean TSP with hyperplane neighborhoods. ACM Transactions on Algorithms, 16(3):38:1– 38:16, June 2020. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3383466.

Avraham:2015:DSF

[AFK^{+}15] Rinat Ben Avraham, Omrit Filtser, Haim Kaplan, Matthew J. Katz, and Micha Sharir. The discrete and semicontinuous Fréchet distance with shortcuts via approximate distance counting and selection. ACM Transactions on Algorithms, 11(4):29:1–29:??, June 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ashok:2018:EAT

[AFK^{+}18] Pradeesha Ashok, Fedor V. Fomin, Sudeshna Kolay, Saket Saurabh, and Meirav Zehavi. Exact algorithms for terrain guarding. ACM Transactions on Algorithms, 14(2):25:1–25:??, June 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ajwani:2008:AIT

[AFM08] Deepak Ajwani, Tobias Friedrich, and Ulrich Meyer. An O(n^{2.75}) algorithm for incremental topological ordering. ACM Transactions on Algorithms, 4(4):39:1–39:??, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Agarwal:2018:CGH

[AFN^{+}18] Pankaj K. Agarwal, Kyle Fox, Abhinandan Nath, Anastasios Sidiropoulos, and Yusu Wang. Computing the Gromov–Hausdorff distance for metric trees. ACM Transactions on Algorithms, 14 (2):24:1–24:??, June 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Asadpour:2012:SCM

[AFS12] Arash Asadpour, Uriel Feige, and Amin Saberi. Santa Claus meets hypergraph matchings. ACM Transactions on Algorithms, 8(3):24:1–24:??, July 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Agarwal:2018:EAC

[AFS18] Pankaj K. Agarwal, Kyle Fox, and Oren Salzman. An efficient algorithm for computing high-quality paths amid polygonal obstacles. ACM Transactions on Algorithms, 14(4):1–21, October 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3230650.

Akitaya:2019:RWE

[AFT19] Hugo A. Akitaya, Radoslav Fulek, and Csaba D. Tóth. Recognizing weak embeddings of graphs. ACM Transactions on Algorithms, 15(4):1–27, October 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3344549.

Alon:2010:BFP

[AG10] Noga Alon and Shai Gutner. Balanced families of perfect hash functions and their applications. ACM Transactions on Algorithms, 6(3):54:1–54:??, June 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Andreev:2009:SSL

[AGG^{+}09] Konstantin Andreev, Charles Garrod, Daniel Golovin, Bruce Maggs, and Adam Meyerson. Simultaneous source location. ACM Transactions on Algorithms, 6(1):16:1–16:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Azar:2010:TUF

[AGG10] Yossi Azar, Iftah Gamzu, and Shai Gutner. Truthful unsplittable flow for large capacity networks. ACM Transactions on Algorithms, 6(2):36:1–36:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Azar:2021:OSD

[AGGP21] Yossi Azar, Arun Ganesh, Rong Ge, and Debmalya Panigrahi. Online service with delay. ACM Transactions on Algorithms, 17 (3):23:1–23:31, August 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3459925.

Alon:2007:GSE

[AGKS07] Noga Alon, Venkatesan Guruswami, Tali Kaufman, and Madhu Sudan. Guessing secrets efficiently via list decoding. ACM Transactions on Algorithms, 3(4):42:1–42:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Anagnostopoulos:2018:MAU

[AGLW18] Aris Anagnostopoulos, Fabrizio Grandoni, Stefano Leonardi, and Andreas Wiese. A mazing 2+ϵ approximation for unsplittable flow on a path. ACM Transactions on Algorithms, 14(4):1–23, October 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3242769.

Abraham:2008:CNI

[AGM^{+}08] Ittai Abraham, Cyril Gavoille, Dahlia Malkhi, Noam Nisan, and Mikkel Thorup. Compact name-independent routing with minimum stretch. ACM Transactions on Algorithms, 4(3):37:1–37:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ambuhl:2011:TEL

[AGP^{+}11] Christoph Ambühl, Leszek Gasieniec, Andrzej Pelc, Tomasz Radzik, and Xiaohui Zhang. Tree exploration with logarithmic memory. ACM Transactions on Algorithms, 7(2):17:1–17:??, March 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ambuhl:2013:OLB

[AGvS13] Christoph Ambühl, Bernd Gärtner, and Bernhard von Stengel. Optimal lower bounds for projective list update algorithms. ACM Transactions on Algorithms, 9(4):31:1–31:??, September 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Abboud:2023:SEB

[AGW23] Amir Abboud, Fabrizio Grandoni, and Virginia Vassilevska Williams. Subcubic equivalences between graph centrality problems, APSP, and diameter. ACM Transactions on Algorithms, 19(1):3:1–3:??, January 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3563393.

Agarwal:2010:HDU

[AHPSW10] Pankaj K. Agarwal, Sariel Har-Peled, Micha Sharir, and Yusu Wang. Hausdorff distance under translation for points and balls. ACM Transactions on Algorithms, 6(4):71:1–71:??, August 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Alstrup:2005:BBC

[AHRT05] Stephen Alstrup, Thore Husfeldt, Theis Rauhe, and Mikkel Thorup. Black box for constant-time insertion in priority queues (note). ACM Transactions on Algorithms, 1(1):102–106, July 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Alstrup:2005:MIF

[AHTL05] Stephen Alstrup, Jacob Holm, Mikkel Thorup, and Kristian De Lichtenberg. Maintaining information in fully dynamic trees with top trees. ACM Transactions on Algorithms, 1(2):243–264, October 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Andoni:2012:SCE

[AK12a] Alexandr Andoni and Robert Krauthgamer. The smoothed complexity of edit distance. ACM Transactions on Algorithms, 8 (4):44:1–44:??, September 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Arvind:2012:TNG

[AK12b] V. Arvind and Piyush P. Kurur. Testing nilpotence of Galois groups in polynomial time. ACM Transactions on Algorithms, 8 (3):32:1–32:??, July 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Aronov:2018:BPL

[AK18] Boris Aronov and Matthew J. Katz. Batched point location in SINR diagrams via algebraic tools. ACM Transactions on Algorithms, 14(4):1–29, October 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3209678.

Aminof:2010:RAO

[AKL10] Benjamin Aminof, Orna Kupferman, and Robby Lampert. Reasoning about online algorithms with weighted automata. ACM Transactions on Algorithms, 6(2):28:1–28:??, March 2010. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Asathulla:2020:FAM

[AKLR20] Mudabir Kabir Asathulla, Sanjeev Khanna, Nathaniel Lahn, and Sharath Raghvendra. A faster algorithm for minimum-cost bipartite perfect matching in planar graphs. ACM Transactions on Algorithms, 16(1):2:1–2:30, January 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3365006.

Aiello:2008:CBM

[AKM08] William Aiello, Alex Kesselman, and Yishay Mansour. Competitive buffer management for shared-memory switches. ACM Transactions on Algorithms, 5(1):3:1–3:??, November 2008. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Azar:2010:DEC

[AKPS10] Yossi Azar, Shay Kutten, and Boaz Patt-Shamir. Distributed error confinement. ACM Transactions on Algorithms, 6(3):48:1– 48:??, June 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Awerbuch:2012:DAM

[AKR12] Baruch Awerbuch, Rohit Khandekar, and Satish Rao. Distributed algorithms for multicommodity flow problems via approximate steepest descent framework. ACM Transactions on Algorithms, 9(1):3:1–3:??, December 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Agarwal:2008:KDD

[AKS08] Pankaj K. Agarwal, Haim Kaplan, and Micha Sharir. Kinetic and dynamic data structures for closest pair and all nearest neighbors. ACM Transactions on Algorithms, 5(1):4:1–4:??, November 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Annamalai:2017:CAR

[AKS17] Chidambaram Annamalai, Christos Kalaitzis, and Ola Svensson. Combinatorial algorithm for restricted max–min fair allocation. ACM Transactions on Algorithms, 13(3):37:1–37:??, August 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

An:2021:AAB

[AKS21] Hyung-Chan An, Robert Kleinberg, and David B. Shmoys. Approximation algorithms for the bottleneck asymmetric traveling salesman problem. ACM Transactions on Algorithms, 17 (4):35:1–35:12, October 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3478537.

Ailon:2013:AOU

[AL13] Nir Ailon and Edo Liberty. An almost optimal unrestricted fast Johnson–Lindenstrauss Transform. ACM Transactions on Algorithms, 9(3):21:1–21:??, June 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Alber:2010:EN

[Alb10a] Susanne Alber. Editorial note. ACM Transactions on Algorithms, 6(4):57:1–57:??, August 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Albers:2010:EN

[Alb10b] Susanne Albers. Editorial Note. ACM Transactions on Algorithms, 6(2):22:1–22:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Agrawal:2025:OCT

[ALL^{+}25] Akanksha Agrawal, Paloma T. Lima, Daniel Lokshtanov, Paweł Rzażewski, Saket Saurabh, and Roohani Sharma. Odd cycle transversal on P_{5}-free graphs in polynomial time. ACM Transactions on Algorithms, 21(2):16:1–16:??, April 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Amir:2007:DTS

[ALLS07] Amihood Amir, Gad M. Landau, Moshe Lewenstein, and Dina Sokol. Dynamic text and static pattern matching. ACM Transactions on Algorithms, 3(2):19:1–19:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Aumann:2011:FWP

[ALLT11] Yonatan Aumann, Moshe Lewenstein, Noa Lewenstein, and Dekel Tsur. Finding witnesses by peeling. ACM Transactions on Algorithms, 7(2):24:1–24:??, March 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Aumann:2012:DIG

[ALM^{+}12] Yonatan Aumann, Moshe Lewenstein, Oren Melamud, Ron Pinter, and Zohar Yakhini. Dotted interval graphs. ACM Transactions on Algorithms, 8(2):9:1–9:??, April 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Agrawal:2019:FVS

[ALM^{+}19] Akanksha Agrawal, Daniel Lokshtanov, Pranabendu Misra, Saket Saurabh, and Meirav Zehavi. Feedback vertex set inspired kernel for chordal vertex deletion. ACM Transactions on Algorithms, 15 (1):1–28, January 2019. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3284356.

Agrawal:2020:PAA

[ALM^{+}20] Akanksha Agrawal, Daniel Lokshtanov, Pranabendu Misra, Saket Saurabh, and Meirav Zehavi. Polylogarithmic approximation algorithms for weighted-F -deletion problems. ACM Transactions on Algorithms, 16(4):51:1–51:38, September 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/10.1145/3389338.

Agrawal:2023:PKI

[ALM^{+}23] Akanksha Agrawal, Daniel Lokshtanov, Pranabendu Misra, Saket Saurabh, and Meirav Zehavi. Polynomial kernel for interval vertex deletion. ACM Transactions on Algorithms, 19(2):11:1–11:??, April 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3571075.

Arya:2007:SEB

[AMM07] Sunil Arya, Theocharis Malamatos, and David M. Mount. A simple entropy-based algorithm for planar point location. ACM Transactions on Algorithms, 3(2):17:1–17:17, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Amanatidis:2017:AAC

[AMNS17] Georgios Amanatidis, Evangelos Markakis, Afshin Nikzad, and Amin Saberi. Approximation algorithms for computing maximin share allocations. ACM Transactions on Algorithms, 13(4):52:1– 52:??, December 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Alagic:2009:QAS

[AMR09] Gorjan Alagic, Cristopher Moore, and Alexander Russell. Quantum algorithms for Simon’s problem over nonabelian groups. ACM Transactions on Algorithms, 6(1):19:1–19:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Alon:2006:ACS

[AMS06] Noga Alon, Dana Moshkovitz, and Shmuel Safra. Algorithmic construction of sets for k-restrictions. ACM Transactions on Algorithms, 2(2):153–177, April 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Alman:2020:DPP

[AMW20] Josh Alman, Matthias Mnich, and Virginia Vassilevska Williams. Dynamic parameterized problems and algorithms. ACM Transactions on Algorithms, 16(4):45:1–45:46, September 2020. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3395037.

Andoni:2016:WPS

[AN16] Alexandr Andoni and Huy L. Nguyên. Width of points in the streaming model. ACM Transactions on Algorithms, 12(1):5:1– 5:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Andrews:2009:IFP

[And09] Matthew Andrews. Instability of FIFO in the permanent sessions model at arbitrarily small network loads. ACM Transactions on Algorithms, 5(3):33:1–33:??, July 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Andersen:2010:LAF

[And10] Reid Andersen. A local algorithm for finding dense subgraphs. ACM Transactions on Algorithms, 6(4):60:1–60:??, August 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

An:2017:DFL

[ANFS17] Hyung-Chan An, Ashkan Norouzi-Fard, and Ola Svensson. Dynamic facility location via exponential clocks. ACM Transactions on Algorithms, 13(2):21:1–21:??, May 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Aggarwal:2010:AAC

[APF^{+}10] Gagan Aggarwal, Rina Panigrahy, Tomás Feder, Dilys Thomas, Krishnaram Kenthapadi, Samir Khuller, and An Zhu. Achieving anonymity via clustering. ACM Transactions on Algorithms, 6 (3):49:1–49:??, June 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Andoni:2017:E

[APP17] Alexandr Andoni, Debmalya Panigrahi, and Marcin Pilipczuk. Editorial. ACM Transactions on Algorithms, 13(2):18:1–18:??, May 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Adler:2005:PMM

[AR05] Micah Adler and Dan Rubenstein. Pricing multicasting in more flexible network models. ACM Transactions on Algorithms, 1(1): 48–73, July 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Azar:2006:IAC

[AR06] Yossi Azar and Yossi Richter. An improved algorithm for CIOQ switches. ACM Transactions on Algorithms, 2(2):282–295, April 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Alonso:2008:ACL

[AR08a] Laurent Alonso and Edward M. Reingold. Average-case lower bounds for the plurality problem. ACM Transactions on Algorithms, 4(3):27:1–27:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Alonso:2008:DP

[AR08b] Laurent Alonso and Edward M. Reingold. Determining plurality. ACM Transactions on Algorithms, 4(3):26:1–26:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Alonso:2009:ACA

[AR09] Laurent Alonso and Edward M. Reingold. Average-case analysis of some plurality algorithms. ACM Transactions on Algorithms, 5(2):17:1–17:??, March 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Awerbuch:2014:PRM

[ARS^{+}14] Baruch Awerbuch, Andrea Richa, Christian Scheideler, Stefan Schmid, and Jin Zhang. Principles of robust medium access and an application to leader election. ACM Transactions on Algorithms, 10(4):24:1–24:??, August 2014. ISSN 1549-6325 (print), 1549-6333 (electronic).

Aspnes:2007:SG

[AS07] James Aspnes and Gauri Shah. Skip graphs. ACM Transactions on Algorithms, 3(4):37:1–37:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Azriel:2008:IFS

[ASS08] Dany Azriel, Noam Solomon, and Shay Solomon. On an infinite family of solvable Hanoi graphs. ACM Transactions on Algorithms, 5(1):13:1–13:??, November 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Amiri:2019:DDS

[ASS19] Saeed Akhoondian Amiri, Stefan Schmid, and Sebastian Siebertz. Distributed dominating set approximations beyond planar graphs. ACM Transactions on Algorithms, 15(3):1–18, July 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3326170.

Agarwal:2008:ACT

[ASW08] Pankaj K. Agarwal, Micha Sharir, and Emo Welzl. Algorithms for center and Tverberg points. ACM Transactions on Algorithms, 5(1):5:1–5:??, November 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Archer:2007:FPM

[AT07] Aaron Archer and Éva Tardos. Frugal path mechanisms. ACM Transactions on Algorithms, 3(1):??, February 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Alstrup:2014:UFC

[ATG^{+}14] Stephen Alstrup, Mikkel Thorup, Inge Li Gørtz, Theis Rauhe, and Uri Zwick. Union-find with constant time deletions. ACM Transactions on Algorithms, 11(1):6:1–6:??, August 2014. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Abrahamsen:2019:CTT

[AW19] Mikkel Abrahamsen and Bartosz Walczak. Common tangents of two disjoint polygons in linear time and constant workspace. ACM Transactions on Algorithms, 15(1):1–21, January 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https:

//dl.acm.org/doi/abs/10.1145/3284355.

Andrews:2007:RSM

[AZ07] Matthew Andrews and Lisa Zhang. Routing and scheduling in multihop wireless networks with time-varying channels. ACM Transactions on Algorithms, 3(3):33:1–33:??, August 2007. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Anshelevich:2008:PDU

[AZ08] Elliot Anshelevich and Lisa Zhang. Path decomposition under a new cost measure with applications to optical network design. ACM Transactions on Algorithms, 4(1):15:1–15:??, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Abbasi-Zadeh:2022:SBR

[AZBG^{+}22] Sepehr Abbasi-Zadeh, Nikhil Bansal, Guru Guruganesh, Aleksandar Nikolov, Roy Schwartz, and Mohit Singh. Sticky Brownian rounding and its applications to constraint satisfaction problems. ACM Transactions on Algorithms, 18(4):33:1–33:??, October 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3459096.

Ben-Aroya:2011:CAF

[BAT11] Avraham Ben-Aroya and Sivan Toledo. Competitive analysis of flash memory algorithms. ACM Transactions on Algorithms, 7(2):

23:1–23:??, March 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Blandford:2008:CDV

[BB08] Daniel K. Blandford and Guy E. Blelloch. Compact dictionaries for variable-length keys and data with applications. ACM Transactions on Algorithms, 4(2):17:1–17:??, May 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Berend:2009:LAC

[BB09] Daniel Berend and Vladimir Braverman. A linear algorithm for computing convex hulls for random lines. ACM Transactions on Algorithms, 5(4):42:1–42:??, October 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Blomer:2020:CTS

[BBB20] Johannes Blömer, Sascha Brauer, and Kathrin Bujna. A complexity theoretical study of fuzzy K-means. ACM Transactions on Algorithms, 16(4):53:1–53:25, September 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/10.1145/3409385.

Bender:2025:PAT

[BBC^{+}25] Michael A. Bender, Abhishek Bhattacharjee, Alex Conway, Martín Farach-Colton, Rob Johnson, Sudarsun Kannan, William Kuszmaul, Nirjhar Mukherjee, Don Porter, Guido Tagliavini, Janet Vorobyeva, and Evan West. Paging and the addresstranslation problem. ACM Transactions on Algorithms, 21(4): 43:1–43:22, October 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bacher:2017:GRP

[BBHT17] Axel Bacher, Olivier Bodini, Hsien-Kuei Hwang, and Tsung-Hsi Tsai. Generating random permutations by coin tossing: Classical algorithms, new analysis, and modern implementation. ACM Transactions on Algorithms, 13(2):24:1–24:43, May 2017. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bienkowski:2019:DBF

[BBM19] Marcin Bienkowski, Jarosław Byrka, and Marcin Mucha. Dynamic beats fixed: On phase-based algorithms for file migration.

ACM Transactions on Algorithms, 15(4):1–21, October 2019. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3340296.

Baswana:2023:MCD

[BBP23] Surender Baswana, Koustav Bhanja, and Abhyuday Pandey. Minimum+1 (s, t)-cuts and dual-edge sensitivity oracle. ACM Transactions on Algorithms, 19(4):38:1–38:??, October 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https:

//dl.acm.org/doi/10.1145/3623271.

Bocker:2012:IFP

[BBT12] Sebastian Böcker, Quang Bao Anh Bui, and Anke Truss. Improved fixed-parameter algorithms for minimum-flip consensus trees. ACM Transactions on Algorithms, 8(1):7:1–7:??, January 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bringmann:2023:LTA

[BCAD23] Karl Bringmann, Vincent Cohen-Addad, and Debarati Das. A linear-time n^{0.4}-approximation for longest common subsequence. ACM Transactions on Algorithms, 19(1):9:1–9:??, January 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3568398.

Bansal:2010:DPI

[BCC^{+}10] Nikhil Bansal, Ning Chen, Neva Cherniavsky, Atri Rurda, Baruch Schieber, and Maxim Sviridenko. Dynamic pricing for impatient bidders. ACM Transactions on Algorithms, 6(2):35:1–35:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bibak:2021:ISC

[BCC21] Ali Bibak, Charles Carlson, and Karthekeyan Chandrasekaran. Improving the smoothed complexity of FLIP for max cut problems. ACM Transactions on Algorithms, 17(3):19:1–19:38, August 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3454125.

Baptiste:2012:PTA

[BCD12] Philippe Baptiste, Marek Chrobak, and Christoph Dürr. Polynomial-time algorithms for minimum energy scheduling.

ACM Transactions on Algorithms, 8(3):26:1–26:??, July 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bagchi:2007:DSR

[BCEG07] Amitabha Bagchi, Amitabh Chaudhary, David Eppstein, and Michael T. Goodrich. Deterministic sampling and range counting in geometric data streams. ACM Transactions on Algorithms, 3 (2):16:1–16:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bender:2025:TP

[BCFC^{+}25] Michael Bender, Alex Conway, Martín Farach-Colton, William Kuszmaul, and Guido Tagliavini. Tiny pointers. ACM Transactions on Algorithms, 21(4):38:1–38:43, October 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bhatia:2007:AAB

[BCFN07] Randeep Bhatia, Julia Chuzhoy, Ari Freund, and Joseph (Seffi) Naor. Algorithmic aspects of bandwidth trading. ACM Transactions on Algorithms, 3(1):??, February 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bose:2012:SGI

[BCH^{+}12] Prosenjit Bose, Eric Y. Chen, Meng He, Anil Maheshwari, and Pat Morin. Succinct geometric indexes supporting point location queries. ACM Transactions on Algorithms, 8(2):10:1–10:??, April 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Baswana:2020:ASS

[BCHR20] Surender Baswana, Keerti Choudhary, Moazzam Hussain, and Liam Roditty. Approximate single-source fault tolerant shortest path. ACM Transactions on Algorithms, 16(4):44:1–44:22, September 2020. CODEN ???? ISSN 1549-6325 (print), 1549-

6333 (electronic). URL https://dl.acm.org/doi/10.1145/
3397532.

Bose:2023:COS

[BCI^{+}23] Prosenjit Bose, Jean Cardinal, John Iacono, Grigorios Koumoutsos, and Stefan Langerman. Competitive online search trees on trees. ACM Transactions on Algorithms, 19(3):25:1–25:??, July 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3595180.

Bliznets:2020:LBP

[BCK^{+}20] Ivan Bliznets, Marek Cygan, Paweł Komosa, Michał Pilipczuk, and Lukás Mach. Lower bounds for the parameterized complexity of minimum fill-in and other completion problems. ACM Transactions on Algorithms, 16(2):25:1–25:31, April 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3381426.

Belazzougui:2020:LTS

[BCKM20] Djamal Belazzougui, Fabio Cunial, Juha Kärkkäinen, and Veli Mäkinen. Linear-time string indexing and analysis in small space. ACM Transactions on Algorithms, 16(2):17:1–17:54, April 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3381417.

Beier:2006:CES

[BCKV06] Rene Beier, Artur Czumaj, Piotr Krysta, and Berthold Vöcking. Computing equilibria for a service provider game with (im)perfect information. ACM Transactions on Algorithms, 2(4):679–706, October 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Borodin:2011:HWC

[BCM11] Allan Borodin, David Cashman, and Avner Magen. How well can primal-dual and local-ratio algorithms perform? ACM Transactions on Algorithms, 7(3):29:1–29:??, July 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bonifaci:2012:ACP

[BCMSM12] Vincenzo Bonifaci, Ho-Leung Chan, Alberto Marchetti-Spaccamela, and Nicole Megow. Algorithms and complexity for periodic realtime scheduling. ACM Transactions on Algorithms, 9(1):6:1–6:??, December 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bassino:2012:COF

[BCN12] Frédérique Bassino, Julien Clément, and Pierre Nicodème. Counting occurrences for a finite set of words: combinatorial methods. ACM Transactions on Algorithms, 8(3):31:1–31:??, July 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bansal:2013:SSA

[BCP13] Nikhil Bansal, Ho-Leung Chan, and Kirk Pruhs. Speed scaling with an arbitrary power function. ACM Transactions on Algorithms, 9(2):18:1–18:??, March 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Blanca:2024:FPS

[BCP24] Antonio Blanca, Sarah Cannon, and Will Perkins. Fast and perfect sampling of subgraphs and polymer systems. ACM Transactions on Algorithms, 20(1):5:1–5:??, January 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/10.1145/3632294.

Blanca:2020:SLH

[BCSV20] Antonio Blanca, Zongchen Chen, Daniel Stefankoviè, and Eric Vigoda. Structure learning of H-colorings. ACM Transactions on Algorithms, 16(3):36:1–36:28, June 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3382207.

Blanca:2025:CHD

[BCSV25] Antonio Blanca, Zongchen Chen, Daniel STefankovic, and Eric Vigoda. Complexity of high-dimensional identity testing with coordinate conditional sampling. ACM Transactions on Algorithms, 21(1):7:1–7:??, January 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bansal:2007:MWF

[BD07] Nikhil Bansal and Kedar Dhamdhere. Minimizing weighted flow time. ACM Transactions on Algorithms, 3(4):39:1–39:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bonsma:2011:TBF

[BD11] Paul Bonsma and Frederic Dorn. Tight bounds and a fast FPT algorithm for directed Max-Leaf Spanning Tree. ACM Transactions on Algorithms, 7(4):44:1–44:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bonamy:2020:EMD

[BDH^{+}20] Marthe Bonamy, Oscar Defrain, Marc Heinrich, Michał Pilipczuk, and Jean-Florent Raymond. Enumerating minimal dominating sets in Kt-free graphs and variants. ACM Transactions on Algorithms, 16(3):39:1–39:23, June 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3386686.

Bouchard:2023:AOD

[BDLP23] Sébastien Bouchard, Yoann Dieudonné, Arnaud Labourel, and Andrzej Pelc. Almost-optimal deterministic treasure hunt in unweighted graphs. ACM Transactions on Algorithms, 19(3):22:1– 22:??, July 2023. CODEN ???? ISSN 1549-6325 (print), 1549-

6333 (electronic). URL https://dl.acm.org/doi/10.1145/
3588437.

Buchin:2023:AMC

[BDR23] Maike Buchin, Anne Driemel, and Dennis Rohde. Approximating (k, l)-median clustering for polygonal curves. ACM Transactions on Algorithms, 19(1):4:1–4:??, January 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3559764.

Bhattacharyya:2019:OAH

[BDW19] Arnab Bhattacharyya, Palash Dey, and David P. Woodruff. An optimal algorithm for l_{1}-heavy hitters in insertion streams and related problems. ACM Transactions on Algorithms, 15(1):1– 27, January 2019. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3264427.

Baier:2010:LBC

[BEH^{+}10] Georg Baier, Thomas Erlebach, Alexander Hall, Ekkehard Köhler, Petr Kolman, Ondrej Pangrác, Heiko Schilling, and Martin Skutella. Length-bounded cuts and flows. ACM Transactions on Algorithms, 7(1):4:1–4:??, November 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bansal:2019:HKS

[BEJK19] Nikhil Bansal, Marek Eliéš, Łukasz Jeż, and Grigorios Koumoutsos. The (h, k)-server problem on bounded depth trees. ACM Transactions on Algorithms, 15(2):1–26, May 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3301314.

Bansal:2023:CAG

[BEKN23] Nikhil Bansal, Marek Eliás, Grigorios Koumoutsos, and Jesper Nederlof. Competitive algorithms for generalized k-server in uniform metrics. ACM Transactions on Algorithms, 19(1):8:1–8:??, January 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3568677.

Boyar:2007:RWO

[BF07] Joan Boyar and Lene M. Favrholdt. The relative worst order ratio for online algorithms. ACM Transactions on Algorithms, 3 (2):22:1–22:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Babai:2009:CRC

[BF09] László Babai and Pedro F. Felzenszwalb. Computing rankconvolutions with a mask. ACM Transactions on Algorithms, 6(1):20:1–20:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Buchbinder:2018:DAS

[BF18] Niv Buchbinder and Moran Feldman. Deterministic algorithms for submodular maximization problems. ACM Transactions on Algorithms, 14(3):32:1–32:??, July 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3184990.

Blasius:2024:EVA

[BF24] Thomas Bläsius and Philipp Fischbeck. On the external validity of average-case analyses of graph algorithms. ACM Transactions on Algorithms, 20(1):10:1–10:??, January 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/10.1145/3633778.

Bender:2017:COS

[BFCF^{+}17] Michael A. Bender, Martín Farach-Colton, Sándor P. Fekete, Jeremy T. Fineman, and Seth Gilbert. Cost-oblivious storage reallocation. ACM Transactions on Algorithms, 13(3):38:1–38:??, August 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Blasius:2022:ESP

[BFF^{+}22] Thomas Bläsius, Cedric Freiberger, Tobias Friedrich, Maximilian Katzmann, Felix Montenegro-Retana, and Marianne Thieffry. Efficient shortest paths in scale-free networks with underlying hyperbolic geometry. ACM Transactions on Algorithms, 18(2):19:1–19:32, April 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3516483.

Bille:2009:IAS

[BFG09] Philip Bille, Rolf Fagerberg, and Inge Li Gørtz. Improved approximate string matching and regular expression matching on Ziv– Lempel compressed texts. ACM Transactions on Algorithms, 6 (1):3:1–3:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bille:2016:STI

[BFG^{+}16] Philip Bille, Johannes Fischer, Inge Li Gørtz, Tsvi Kopelowitz, Benjamin Sach, and Hjalte Wedel Vildhøj. Sparse text indexing in small space. ACM Transactions on Algorithms, 12(3):39:1–39:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bender:2016:NAI

[BFGT16] Michael A. Bender, Jeremy T. Fineman, Seth Gilbert, and Robert E. Tarjan. A new approach to incremental cycle detection and related problems. ACM Transactions on Algorithms, 12 (2):14:1–14:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bernstein:2021:DAD

[BFH21] Aaron Bernstein, Sebastian Forster, and Monika Henzinger. A deamortization approach for dynamic spanner and dynamic maximal matching. ACM Transactions on Algorithms, 17(4):29:1– 29:51, October 2021. CODEN ???? ISSN 1549-6325 (print), 1549-

6333 (electronic). URL https://dl.acm.org/doi/10.1145/
3469833.

Bodlaender:2012:EAT

[BFK^{+}12] Hans L. Bodlaender, Fedor V. Fomin, Arie M. C. A. Koster, Dieter Kratsch, and Dimitrios M. Thilikos. On exact algorithms for treewidth. ACM Transactions on Algorithms, 9(1):12:1–12:??, December 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Boczkowski:2021:NTP

[BFKR21] Lucas Boczkowski, Uriel Feige, Amos Korman, and Yoav Rodeh. Navigating in trees with permanently noisy advice. ACM Transactions on Algorithms, 17(2):15:1–15:27, June 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/10.1145/3448305.

Boczkowski:2025:QCS

[BFKR25] Lucas Boczkowski, Uriel Feige, Amos Korman, and Yoav Rodeh. The query complexity of searching trees with permanently noisy advice. ACM Transactions on Algorithms, 21(2):18:1–18:??, April 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bansal:2014:LAU

[BFKS14] Nikhil Bansal, Zachary Friggstad, Rohit Khandekar, and Mohammad R. Salavatipour. A logarithmic approximation for unsplittable flow on line graphs. ACM Transactions on Algorithms, 10(1):1:1–1:??, January 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Buchin:2023:FDU

[BFL^{+}23] Kevin Buchin, Chenglin Fan, Maarten Löffler, Aleksandr Popov, Benjamin Raichel, and Marcel Roeloffzen. Fréchet distance for uncertain curves. ACM Transactions on Algorithms, 19(3):29:1– 29:??, July 2023. CODEN ???? ISSN 1549-6325 (print), 1549-

6333 (electronic). URL https://dl.acm.org/doi/10.1145/
3597640.

Bliznets:2018:SPA

[BFPP18] Ivan Bliznets, Fedor V. Fomin, Marcin Pilipczuk, and MichaL Pilipczuk. Subexponential parameterized algorithm for interval completion. ACM Transactions on Algorithms, 14(3):35:1–35:??, July 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3186896.

Blasius:2023:SPA

[BFR23] Thomas Bläsius, Simon D. Fink, and Ignaz Rutter. Synchronized planarity with applications to constrained planarity problems. ACM Transactions on Algorithms, 19(4):34:1–34:??, October 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3607474.

Buchbinder:2019:OSM

[BFS19] Niv Buchbinder, Moran Feldman, and Roy Schwartz. Online submodular maximization with preemption. ACM Transactions on Algorithms, 15(3):1–31, July 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3309764.

Boffa:2022:LAD

[BFV22] Antonio Boffa, Paolo Ferragina, and Giorgio Vinciguerra. A learned approach to design compressed rank/select data structures. ACM Transactions on Algorithms, 18(3):24:1–24:??, July 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3524060.

Bille:2011:TIP

[BG11] Philip Bille and Inge Li Gortz. The tree inclusion problem: In linear space and faster. ACM Transactions on Algorithms, 7(3): 38:1–38:47, July 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Breslauer:2014:RTS

[BG14] Dany Breslauer and Zvi Galil. Real-time streaming stringmatching. ACM Transactions on Algorithms, 10(4):22:1–22:??, August 2014. ISSN 1549-6325 (print), 1549-6333 (electronic).

Backens:2020:HCA

[BG20] Miriam Backens and Leslie Ann Goldberg. Holant clones and the approximability of conservative Holant problems. ACM Transactions on Algorithms, 16(2):23:1–23:55, April 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3381425.

Brakensiek:2021:QSI

[BG21] Joshua Brakensiek and Venkatesan Guruswami. The quest for strong inapproximability results with perfect completeness. ACM Transactions on Algorithms, 17(3):27:1–27:35, August 2021. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3459668.

Babenko:2016:AHL

[BGGN16] Maxim Babenko, Andrew V. Goldberg, Anupam Gupta, and Viswanath Nagarajan. Algorithms for hub label optimization. ACM Transactions on Algorithms, 13(1):16:1–16:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bezakova:2024:FSS

[BGGS24] Ivona Bezáková, Andreas Galanis, Leslie Ann Goldberg, and Daniel Stefankovic. Fast sampling via spectral independence beyond bounded-degree graphs. ACM Transactions on Algorithms, 20(1):7:1–7:??, January 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3631354.

Bei:2019:APA

[BGH19] Xiaohui Bei, Jugal Garg, and Martin Hoefer. Ascending-price algorithms for unknown markets. ACM Transactions on Algorithms, 15(3):1–33, July 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3319394.

Bhattacharya:2022:FDC

[BGK^{+}22] Sayan Bhattacharya, Fabrizio Grandoni, Janardhan Kulkarni, Quanquan C. Liu, and Shay Solomon. Fully dynamic (Δ + 1)- coloring in O(1) update time. ACM Transactions on Algorithms, 18(2):10:1–10:25, April 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3494539.

Belova:2025:PFB

[BGK^{+}25] Tatiana Belova, Alexander Golovnev, Alexander S. Kulikov, Ivan Mihajlin, and Denil Sharipov. Polynomial formulations as a barrier for reduction-based hardness proofs. ACM Transactions on Algorithms, 21(4):40:1–40:44, October 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bein:2009:KYQ

[BGLZ09] Wolfgang Bein, Mordecai J. Golin, Lawrence L. Larmore, and Yan Zhang. The Knuth–Yao quadrangle-inequality speedup is a consequence of total monotonicity. ACM Transactions on Algorithms, 6(1):17:1–17:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bringmann:2020:TED

[BGMW20] Karl Bringmann, Paweł Gawrychowski, Shay Mozes, and Oren Weimann. Tree edit distance cannot be computed in strongly subcubic time (unless APSP can). ACM Transactions on Algorithms, 16(4):48:1–48:22, September 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3381878.

Bhattacharyya:2018:EAS

[BGN^{+}18] Arnab Bhattacharyya, Fabrizio Grandoni, Aleksandar Nikolov, Barna Saha, Saket Saurabh, Aravindan Vijayaraghavan, and Qin Zhang. Editorial: ACM-SIAM Symposium on Discrete Algorithms (SODA) 2016 special issue. ACM Transactions on Algorithms, 14(3):26:1–26:??, July 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3230647.

Buchsbaum:2008:RLC

[BGPV08] Adam L. Buchsbaum, Emden R. Gansner, Cecilia M. Procopiuc, and Suresh Venkatasubramanian. Rectangular layouts and contact graphs. ACM Transactions on Algorithms, 4(1):8:1–8:??, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bille:2023:SIC

[BGS23] Philip Bille, Inge Li Gørtz, and Teresa Anna Steiner. String indexing with compressed patterns. ACM Transactions on Algorithms, 19(4):32:1–32:??, October 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3607141.

Bateni:2012:APC

[BH12] Mohammadhossein Bateni and Mohammadtaghi Hajiaghayi. Assignment problem in content distribution networks: Unsplittable hard-capacitated facility location. ACM Transactions on Algorithms, 8(3):20:1–20:??, July 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Brenner:2019:FCB

[BH19] Ulrich Brenner and Anna Hermann. Faster carry bit computation for adder circuits with prescribed arrival times. ACM Transactions on Algorithms, 15(4):1–23, October 2019. CODEN ????

ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3340321.

Bjelde:2021:TBO

[BHD^{+}21] Antje Bjelde, Jan Hackfeld, Yann Disser, Christoph Hansknecht, Maarten Lipmann, Julie Meißner, Miriam SchlÖter, Kevin Schewior, and Leen Stougie. Tight bounds for online TSP on the line. ACM Transactions on Algorithms, 17(1):3:1–3:58, January 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3422362.

Bjorklund:2016:FZT

[BHK^{+}16] Andreas Björklund, Thore Husfeldt, Petteri Kaski, Mikko Koivisto, Jesper Nederlof, and Pekka Parviainen. Fast zeta transforms for lattices with few irreducibles. ACM Transactions on Algorithms, 12(1):4:1–4:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bjorklund:2012:TSP

[BHKK12] Andreas Björklund, Thore Husfeldt, Petteri Kaski, and Mikko Koivisto. The traveling salesman problem in bounded degree graphs. ACM Transactions on Algorithms, 8(2):18:1–18:??, April 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Butman:2010:OPM

[BHLR10] Ayelet Butman, Danny Hermelin, Moshe Lewenstein, and Dror Rawitz. Optimization problems in multiple-interval graphs. ACM Transactions on Algorithms, 6(2):40:1–40:??, March 2010. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Barbay:2011:SIS

[BHMS11] Jérémy Barbay, Meng He, J. Ian Munro, and Srinivasa Rao Satti. Succinct indexes for strings, binary relations and multilabeled trees. ACM Transactions on Algorithms, 7(4):52:1–52:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Blum:2019:SAG

[BHPR19] Avrim Blum, Sariel Har-Peled, and Benjamin Raichel. Sparse approximation via generating point sets. ACM Transactions on Algorithms, 15(3):1–16, July 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3302249.

Beame:2020:EEI

[BHPR^{+}20] Paul Beame, Sariel Har-Peled, Sivaramakrishnan Natarajan Ramamoorthy, Cyrus Rashtchian, and Makrand Sinha. Edge estimation with independent set oracles. ACM Transactions on Algorithms, 16(4):52:1–52:27, September 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3404867.

Berenbrink:2014:DSL

[BHS14] Petra Berenbrink, Martin Hoefer, and Thomas Sauerwald. Distributed selfish load balancing on networks. ACM Transactions on Algorithms, 11(1):2:1–2:??, August 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Balcan:2020:CCU

[BHW20] Maria-Florina Balcan, Nika Haghtalab, and Colin White. k-center clustering under perturbation resilience. ACM Transactions on Algorithms, 16(2):22:1–22:39, April 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3381424.

Bateni:2013:SSP

[BHZ13] Mohammadhossein Bateni, Mohammadtaghi Hajiaghayi, and Morteza Zadimoghaddam. Submodular secretary problem and extensions. ACM Transactions on Algorithms, 9(4):32:1–32:??, September 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bienkowski:2018:DOS

[BJKK18] Marcin Bienkowski, Tomasz Jurdzinski, Mirosław Korzeniowski, and Dariusz R. Kowalski. Distributed online and stochastic queueing on a multiple access channel. ACM Transactions on Algorithms, 14(2):21:1–21:??, June 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Borodin:2017:MSD

[BJLY17] Allan Borodin, Aadhar Jain, Hyun Chul Lee, and Yuli Ye. Maxsum diversification, monotone submodular functions, and dynamic updates. ACM Transactions on Algorithms, 13(3):41:1– 41:??, August 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Barbay:2008:ARA

[BK08] Jérémy Barbay and Claire Kenyon. Alternation and redundancy analysis of the intersection problem. ACM Transactions on Algorithms, 4(1):4:1–4:??, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Borradaile:2016:TEC

[BK16] Glencora Borradaile and Philip Klein. The two-edge connectivity survivable-network design problem in planar graphs. ACM Transactions on Algorithms, 12(3):30:1–30:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Boehmer:2024:CFF

[BK24] Niclas Boehmer and Tomohiro Koana. The complexity of finding fair many-to-one matchings. ACM Transactions on Algorithms, 20(2):17:1–17:??, April 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3649220.

Bjorklund:2017:CTS

[BKK17] Andreas Björklund, Petteri Kaski, and Lukasz Kowalik. Counting thin subgraphs via packings faster than meet-in-the-middle time. ACM Transactions on Algorithms, 13(4):48:1–48:??, December 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Berenbrink:2019:IAD

[BKK^{+}19] Petra Berenbrink, Ralf Klasing, Adrian Kosowski, Frederik Mallmann-Trenn, and Przemysław Uznański. Improved analysis of deterministic load-balancing schemes. ACM Transactions on Algorithms, 15(1):1–22, January 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3282435.

Becchetti:2007:SCM

[BKLP07] L. Becchetti, J. Könemann, S. Leonardi, and M. Páal. Sharing the cost more efficiently: Improved approximation for multicommodity rent-or-buy. ACM Transactions on Algorithms, 3(2): 23:1–23:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Borradaile:2009:ASS

[BKM09a] Glencora Borradaile, Philip Klein, and Claire Mathieu. An O(n log n) approximation scheme for Steiner tree in planar graphs. ACM Transactions on Algorithms, 5(3):31:1–31:??, July 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Borradaile:2009:LAS

[BKM09b] Glencora Borradaile, Philip Klein, and Claire Mathieu. An O(n log n) approximation scheme for Steiner tree in planar graphs. ACM Transactions on Algorithms, 5(3):31:1–31:??, July 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Borradaile:2015:PTA

[BKM15] Glencora Borradaile, Philip N. Klein, and Claire Mathieu. A polynomial-time approximation scheme for Euclidean Steiner forest. ACM Transactions on Algorithms, 11(3):19:1–19:??, January 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Baswana:2010:ASS

[BKMP10] Surender Baswana, Telikepalli Kavitha, Kurt Mehlhorn, and Seth Pettie. Additive spanners and (α, β)-spanners. ACM Transactions on Algorithms, 7(1):5:1–5:??, November 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bonifaci:2011:MFT

[BKMSS11] Vincenzo Bonifaci, Peter Korteweg, Alberto Marchetti-Spaccamela, and Leen Stougie. Minimizing flow time in the wireless gathering problem. ACM Transactions on Algorithms, 7(3):33:1–33:??, July 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Berger:2020:TSO

[BKMV20] André Berger, László Kozma, Matthias Mnich, and Roland Vincze. Time- and space-optimal algorithm for the many-visits TSP. ACM Transactions on Algorithms, 16(3):35:1–35:22, June 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3382038.

Bansal:2014:BSA

[BKN14] Nikhil Bansal, Ravishankar Krishnaswamy, and Viswanath Nagarajan. Better scalable algorithms for broadcast scheduling. ACM Transactions on Algorithms, 11(1):3:1–3:??, August 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bringmann:2021:DFD

[BKN21] Karl Bringmann, Marvin KüNnemann, and André Nusser. Discrete Fréchet distance under translation: Conditional hardness and an improved algorithm. ACM Transactions on Algorithms, 17(3):25:1–25:42, August 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3460656.

Bansal:2025:IAS

[BKN^{+}25] Nikhil Bansal, Eun Jung Kim, Viswanath Nagarajan, Aaron Potechin, and Lars Rohwedder. Introduction: ACM–SIAM Symposium on Discrete Algorithms (SODA) 2023 special issue. ACM Transactions on Algorithms, 21(4):34:1–34:2, October 2025. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Baswana:2012:FDR

[BKS12] Surender Baswana, Sumeet Khurana, and Soumojit Sarkar. Fully dynamic randomized algorithms for graph spanners. ACM Transactions on Algorithms, 8(4):35:1–35:??, September 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Blaser:2008:NAA

[Blä08] Markus Bläser. A new approximation algorithm for the asymmetric TSP with triangle inequality. ACM Transactions on Algorithms, 4(4):47:1–47:??, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Blasiok:2020:OST

[Bła20] Jarosław Błasiok. Optimal streaming and tracking distinct elements with high probability. ACM Transactions on Algorithms, 16(1):3:1–3:28, January 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3309193.

Bandyapadhyay:2024:TCD

[BLL^{+}24] Sayan Bandyapadhyay, William Lochet, Daniel Lokshtanov, Saket Saurabh, and Jie Xue. True contraction decomposition and almost ETH-tight bipartization for unit-disk graphs. ACM Transactions on Algorithms, 20(3):20:1–20:??, July 2024. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3656042.

Biedl:2013:MOP

[BLPS13] Therese Biedl, Anna Lubiw, Mark Petrick, and Michael Spriggs. Morphing orthogonal planar graph drawings. ACM Transactions on Algorithms, 9(4):29:1–29:??, September 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Brandstadt:2008:SLT

[BLS08] Andreas Brandstädt, Van Bang Le, and R. Sritharan. Structure and linear-time recognition of 4-leaf powers. ACM Transactions on Algorithms, 5(1):11:1–11:??, November 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Birmpilis:2023:CAC

[BLS23] Stavros Birmpilis, George Labahn, and Arne Storjohann. A cubic algorithm for computing the Hermite normal form of a nonsingular integer matrix. ACM Transactions on Algorithms, 19 (4):37:1–37:??, October 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3617996.

Brodal:2025:SFH

[BLT25] Gerth Stølting Brodal, George Lagogiannis, and Robert E. Tarjan. Strict Fibonacci heaps. ACM Transactions on Algorithms, 21 (2):15:1–15:??, April 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bruss:2009:IAI

[BLW09] F. Thomas Bruss, Guy Louchard, and Mark Daniel Ward. Inverse auctions: Injecting unique minima into random sets. ACM Transactions on Algorithms, 6(1):21:1–21:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Boyar:2008:RWO

[BM08] Joan Boyar and Paul Medvedev. The relative worst order ratio applied to seat reservation. ACM Transactions on Algorithms, 4(4):48:1–48:??, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bonnet:2020:PHA

[BM20] Édouard Bonnet and Tillmann Miltzow. Parameterized hardness of art gallery problems. ACM Transactions on Algorithms, 16 (4):42:1–42:23, September 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3398684.

Berman:2022:TTI

[BMR22] Piotr Berman, Meiram Murzabulatov, and Sofya Raskhodnikova. Tolerant testers of image properties. ACM Transactions on Algorithms, 18(4):37:1–37:??, October 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3531527.

Becchetti:2009:LCA

[BMSV^{+}09] Luca Becchetti, Alberto Marchetti-Spaccamela, Andrea Vitaletti, Peter Korteweg, Martin Skutella, and Leen Stougie. Latencyconstrained aggregation in sensor networks. ACM Transactions on Algorithms, 6(1):13:1–13:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Belazzougui:2014:AIC

[BN14] Djamal Belazzougui and Gonzalo Navarro. Alphabet-independent compressed text indexing. ACM Transactions on Algorithms, 10 (4):23:1–23:??, August 2014. ISSN 1549-6325 (print), 1549-6333 (electronic).

Belazzougui:2015:OLU

[BN15] Djamal Belazzougui and Gonzalo Navarro. Optimal lower and upper bounds for representing sequences. ACM Transactions on Algorithms, 11(4):31:1–31:??, June 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bar-Noy:2008:DCF

[BNCS08] Amotz Bar-Noy, Panagiotis Cheilaris, and Shakhar Smorodinsky. Deterministic conflict-free coloring for intervals: From offline to online. ACM Transactions on Algorithms, 4(4):44:1–44:18, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bar-Noy:2009:TMR

[BNGK^{+}09] Amotz Bar-Noy, Sudipto Guha, Yoav Katz, Joseph (Seffi) Naor, Baruch Schieber, and Hadas Shachnai. Throughput maximization of real-time scheduling with batching. ACM Transactions on Algorithms, 5(2):18:1–18:??, March 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bar-Noy:2007:WSR

[BNLT07] Amotz Bar-Noy, Richard E. Ladner, and Tami Tamir. Windows scheduling as a restricted version of bin packing. ACM Transactions on Algorithms, 3(3):28:1–28:??, August 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bun:2019:HHS

[BNS19] Mark Bun, Jelani Nelson, and Uri Stemmer. Heavy hitters and the structure of local privacy. ACM Transactions on Algorithms, 15(4):1–40, October 2019. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3344722.

Bar-On:2020:TBO

[BODD^{+}20] Achiya Bar-On, Itai Dinur, Orr Dunkelman, Rani Hod, Nathan Keller, Eyal Ronen, and Adi Shamir. Tight bounds on online checkpointing algorithms. ACM Transactions on Algorithms, 16 (3):31:1–31:22, June 2020. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3379543.

Bonnet:2022:VSU

[Bon22] Édouard Bonnet. 4 vs 7 sparse undirected unweighted diameter is SETH-hard at time n^{4/3}. ACM Transactions on Algorithms, 18(2):11:1–11:14, April 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3494540.

Berry:2009:LTA

[BPGN09] Vincent Berry, Christophe Paul, Sylvain Guillemot, and François Nicolas. Linear time 3-approximation for the MAST problem. ACM Transactions on Algorithms, 5(2):23:1–23:??, March 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Barbay:2018:ACS

[BPL18] Jérémy Barbay and Pablo Pérez-Lantero. Adaptive computation of the swap-insert correction distance. ACM Transactions on Algorithms, 14(4):1–16, October 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3232057.

Byrka:2017:IAM

[BPR^{+}17] Jaroslaw Byrka, Thomas Pensyl, Bartosz Rybicki, Aravind Srinivasan, and Khoa Trinh. An improved approximation for k-median and positive correlation in budgeted optimization. ACM Transactions on Algorithms, 13(2):23:1–23:??, May 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Berman:2014:AAM

[BR14] Piotr Berman and Sofya Raskhodnikova. Approximation algorithms for min-max generalization problems. ACM Transactions on Algorithms, 11(1):5:1–5:??, August 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Blasius:2016:SPO

[BR16] Thomas Bläsius and Ignaz Rutter. Simultaneous PQ-ordering with applications to constrained embedding problems. ACM Transactions on Algorithms, 12(2):16:1–16:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bressan:2023:ENO

[Bre23] Marco Bressan. Efficient and near-optimal algorithms for sampling small connected subgraphs. ACM Transactions on Algorithms, 19(3):26:1–26:??, July 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/

doi/10.1145/3596495.

Binkele-Raible:2012:KPN

[BRFF^{+}12] Daniel Binkele-Raible, Henning Fernau, Fedor V. Fomin, Daniel Lokshtanov, Saket Saurabh, and Yngve Villanger. Kernel(s) for problems with no kernel: On out-trees with many leaves. ACM Transactions on Algorithms, 8(4):38:1–38:??, September 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Blasius:2016:OOG

[BRW16] Thomas Bläsius, Ignaz Rutter, and Dorothea Wagner. Optimal orthogonal graph drawing with convex bend costs. ACM Transactions on Algorithms, 12(3):33:1–33:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Baswana:2006:ADO

[BS06a] Surender Baswana and Sandeep Sen. Approximate distance oracles for unweighted graphs in expected O(n^{2}) time. ACM Transactions on Algorithms, 2(4):557–577, October 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Berend:2006:CMP

[BS06b] Daniel Berend and Amir Sapir. The cyclic multi-peg Tower of Hanoi. ACM Transactions on Algorithms, 2(3):297–317, July 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bobier:2010:FAG

[BS10] Bruce Bobier and Joe Sawada. A fast algorithm to generate open meandric systems and meanders. ACM Transactions on Algorithms, 6(2):42:1–42:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Boissonnat:2018:ERF

[BS18] Jean-Daniel Boissonnat and Karthik C. S. An efficient representation for filtrations of simplicial complexes. ACM Transactions on Algorithms, 14(4):1–21, October 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3229146.

Brubach:2020:AAC

[BSSX20] Brian Brubach, Karthik A. Sankararaman, Aravind Srinivasan, and Pan Xu. Algorithms to approximate column-sparse packing problems. ACM Transactions on Algorithms, 16(1):10:1– 10:32, January 2020. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3355400.

Berend:2008:CDG

[BST08] Daniel Berend, Steven S. Skiena, and Yochai Twitto. Combinatorial dominance guarantees for problems with infeasible solutions. ACM Transactions on Algorithms, 5(1):8:1–8:??, November 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Borradaile:2015:MSC

[BSWN15] Glencora Borradaile, Piotr Sankowski, and Christian Wulff-Nilsen. Min st-cut oracle for planar graphs with near-linear preprocessing time. ACM Transactions on Algorithms, 11(3):16:1– 16:??, January 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Beretta:2024:BSE

[BT24] Lorenzo Beretta and Jakub Tetek. Better sum estimation via weighted sampling. ACM Transactions on Algorithms, 20 (3):27:1–27:??, July 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3650030.

Bhore:2025:OES

[BT25] Sujoy Bhore and Csaba D. Tóth. Online Euclidean spanners. ACM Transactions on Algorithms, 21(1):5:1–5:??, January 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Buchsbaum:2008:GE

[Buc08] Adam L. Buchsbaum. Guest editorial. ACM Transactions on Algorithms, 4(2):16:1–16:??, May 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Bodwin:2021:BDP

[BW21] Greg Bodwin and Virginia Vassilevska Williams. Better distance preservers and additive spanners. ACM Transactions on Algorithms, 17(4):36:1–36:24, October 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3490147.

Cabello:2010:FSC

[Cab10] Sergio Cabello. Finding shortest contractible and shortest separating cycles in embedded graphs. ACM Transactions on Algorithms, 6(2):24:1–24:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cabello:2019:SAD

[Cab19] Sergio Cabello. Subquadratic algorithms for the diameter and the sum of pairwise distances in planar graphs. ACM Transactions on Algorithms, 15(2):1–38, May 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3218821.

Cabello:2022:CIG

[Cab22] Sergio Cabello. Computing the inverse geodesic length in planar graphs and graphs of bounded treewidth. ACM Transactions on Algorithms, 18(2):14:1–14:26, April 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3501303.

Carlsson:2010:FEC

[CAY10] John Gunnar Carlsson, Benjamin Armbruster, and Yinyu Ye. Finding equitable convex partitions of points in a polygon efficiently. ACM Transactions on Algorithms, 6(4):72:1–72:??, August 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Charron-Bost:2015:TCL

[CBFWW15] Bernadette Charron-Bost, Matthias Függer, Jennifer L. Welch, and Josef Widder. Time complexity of link reversal routing. ACM Transactions on Algorithms, 11(3):18:1–18:??, January 2015. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chakraborty:2025:NEB

[CC25] Diptarka Chakraborty and Keerti Choudhary. New extremal bounds for reachability and strong-connectivity preservers under failures. ACM Transactions on Algorithms, 21(2):21:1–21:??, April 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cohen:2016:TBS

[CCDL16] Edith Cohen, Graham Cormode, Nick Duffield, and Carsten Lund. On the tradeoff between stability and fit. ACM Transactions on Algorithms, 13(1):7:1–7:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Caceres:2024:WHH

[CCG^{+}24] Manuel Cáceres, Massimo Cairo, Andreas Grigorjew, Shahbaz Khan, Brendan Mumey, Romeo Rizzi, Alexandru I. Tomescu, and Lucia Williams. Width helps and hinders splitting flows. ACM Transactions on Algorithms, 20(2):13:1–13:??, April 2024. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3641820.

Chitnis:2015:DSF

[CCHM15] Rajesh Chitnis, Marek Cygan, Mohammataghi Hajiaghayi, and Dániel Marx. Directed subset feedback vertex set is fixedparameter tractable. ACM Transactions on Algorithms, 11(4): 28:1–28:??, June 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chekuri:2012:SMP

[CCHP12] Chandra Chekuri, Kenneth L. Clarkson, and Sariel Har-Peled. On the set multicover problem in geometric settings. ACM Transactions on Algorithms, 9(1):9:1–9:??, December 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chee:2019:DCW

[CCKN19] Yeow Meng Chee, Johan Chrisnata, Han Mao Kiah, and Tuan Thanh Nguyen. Deciding the confusability of words under tandem repeats in linear time. ACM Transactions on Algorithms, 15(3):1–22, July 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3338514.

Calinescu:2011:IAA

[CCKR11] Gruia Calinescu, Amit Chakrabarti, Howard Karloff, and Yuval Rabani. An improved approximation algorithm for resource allocation. ACM Transactions on Algorithms, 7(4):48:1–48:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2009:OTE

[CCL^{+}09] Ho-Leung Chan, Joseph Wun-Tat Chan, Tak-Wah Lam, Lap-Kei Lee, Kin-Sum Mak, and Prudence W. H. Wong. Optimizing throughput and energy in online deadline scheduling. ACM Transactions on Algorithms, 6(1):10:1–10:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Campos:2025:NML

[CCLS25] Victor Campos, Jonas Costa, Raul Lopes, and Ignasi Sau. New Menger-like dualities in digraphs and applications to half-integral linkages. ACM Transactions on Algorithms, 21(3):26:1–26:28, July 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chakrabarti:2010:NOA

[CCM10] Amit Chakrabarti, Graham Cormode, and Andrew Mcgregor. A near-optimal algorithm for estimating the entropy of a stream. ACM Transactions on Algorithms, 6(3):51:1–51:??, June 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chakrabarti:2014:ADS

[CCMT14] Amit Chakrabarti, Graham Cormode, Andrew Mcgregor, and Justin Thaler. Annotations in data streams. ACM Transactions on Algorithms, 11(1):7:1–7:??, August 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2018:ANW

[CCW18] T.-H. Hubert Chan, Fei Chen, and Xiaowei Wu. Analyzing nodeweighted oblivious matching problem via continuous LP with jump discontinuity. ACM Transactions on Algorithms, 14(2): 12:1–12:??, June 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cohavi:2017:FSS

[CD17] Keren Cohavi and Shahar Dobzinski. Faster and simpler sketches of valuation functions. ACM Transactions on Algorithms, 13(3): 30:1–30:??, August 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Conradi:2024:CSF

[CD24] Jacobus Conradi and Anne Driemel. On computing the kshortcut Fréchet distance. ACM Transactions on Algorithms, 20 (4):29:1–29:??, October 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3663762.

Chalopin:2015:MSP

[CDD^{+}15] Jérémie Chalopin, Shantanu Das, Yann Disser, Matús Mihalák, and Peter Widmayer. Mapping simple polygons: The power of telling convex from reflex. ACM Transactions on Algorithms, 11 (4):33:1–33:??, June 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cabello:2010:FOT

[CDEM10] Sergio Cabello, Matt Devos, Jeff Erickson, and Bojan Mohar. Finding one tight cycle. ACM Transactions on Algorithms, 6(4):

61:1–61:??, August 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Condon:2009:ADA

[CDHW09] Anne Condon, Amol Deshpande, Lisa Hellerstein, and Ning Wu. Algorithms for distributional and adversarial pipelined filter ordering problems. ACM Transactions on Algorithms, 5(2):24:1– 24:??, March 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Collette:2012:ETP

[CDI^{+}12] Sébastien Collette, Vida Dujmović, John Iacono, Stefan Langerman, and Pat Morin. Entropy, triangulation, and point location in planar subdivisions. ACM Transactions on Algorithms, 8(3): 29:1–29:??, July 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chang:2023:NOT

[CDJ23] Yi-Jun Chang, Ran Duan, and Shunhua Jiang. Near-optimal time-energy tradeoffs for deterministic leader election. ACM Transactions on Algorithms, 19(4):33:1–33:??, October 2023. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3614429.

Chakrabarty:2017:PTP

[CDJS17] Deeparnab Chakrabarty, Kashyap Dixit, Madhav Jha, and C. Seshadhri. Property testing on product distributions: Optimal testers for bounded derivative properties. ACM Transactions on Algorithms, 13(2):20:1–20:??, May 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chlamtac:2020:ASD

[CDKL20] Eden Chlamtác, Michael Dinitz, Guy Kortsarz, and Bundit Laekhanukit. Approximating spanners and directed Steiner forest: Upper and lower bounds. ACM Transactions on Algorithms, 16(3):33:1–33:31, June 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3381451.

Cygan:2016:PHC

[CDL^{+}16] Marek Cygan, Holger Dell, Daniel Lokshtanov, Dániel Marx, Jesper Nederlof, Yoshio Okamoto, Ramamohan Paturi, Saket Saurabh, and Magnus Wahlström. On problems as hard as CNF-SAT. ACM Transactions on Algorithms, 12(3):41:1–41:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chen:2025:PTT

[CDL^{+}25] Xi Chen, Anindya De, Chin Ho Lee, Rocco Servedio, and Sandip Sinha. Polynomial-time trace reconstruction in the smoothed complexity model. ACM Transactions on Algorithms, 21(3):30:1– 30:28, July 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chang:2022:TCS

[CdM22] Hsien-Chih Chang and Arnaud de Mesmay. Tightening curves on surfaces monotonically with applications. ACM Transactions on Algorithms, 18(4):36:1–36:??, October 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3558097.

Coudert:2019:FPF

[CDP19] David Coudert, Guillaume Ducoffe, and Alexandru Popa. Fully polynomial FPT algorithms for some classes of bounded cliquewidth graphs. ACM Transactions on Algorithms, 15(3):1–57, July 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3310228.

Czumaj:2021:GSD

[CDP21] Artur Czumaj, Peter Davies, and Merav Parter. Graph sparsification for derandomizing massively parallel computation with low space. ACM Transactions on Algorithms, 17(2):16:1–16:27, June 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3451992.

Chang:2011:BSA

[CEGK11] Jessica Chang, Thomas Erlebach, Renars Gailis, and Samir Khuller. Broadcast scheduling: Algorithms and complexity. ACM Transactions on Algorithms, 7(4):47:1–47:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chekuri:2011:SCP

[CEGS11] Chandra Chekuri, Guy Even, Anupam Gupta, and Danny Segev. Set connectivity problems in undirected graphs and the directed Steiner network problem. ACM Transactions on Algorithms, 7(2): 18:1–18:??, March 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Christiansen:2021:OTD

[CEK^{+}21] Anders Roy Christiansen, Mikko Berggren Ettienne, Tomasz Kociumaka, Gonzalo Navarro, and Nicola Prezza. Optimal-time dictionary-compressed indexes. ACM Transactions on Algorithms, 17(1):8:1–8:39, January 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/

doi/10.1145/3426473.

Chekuri:2021:NWN

[CEV21] Chandra Chekuri, Alina Ene, and Ali Vakilian. Node-weighted network design in planar and minor-closed families of graphs. ACM Transactions on Algorithms, 17(2):14:1–14:25, June 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3447959.

Cechlarova:2005:GSR

[CF05] Katarína Cechlárová and Tamás Fleiner. On a generalization of the stable roommates problem. ACM Transactions on Algorithms, 1(1):143–156, July 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cseh:2020:CCC

[CF20] Ágnes Cseh and Tamás Fleiner. The complexity of cake cutting with unequal shares. ACM Transactions on Algorithms, 16(3): 29:1–29:21, June 2020. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3380742.

Chern:2007:PCR

[CFH07] Hua-Huai Chern, Michael Fuchs, and Hsien-Kuei Hwang. Phase changes in random point quadtrees. ACM Transactions on Algorithms, 3(2):12:1–12:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cohen:2008:LGG

[CFI^{+}08] Reuven Cohen, Pierre Fraigniaud, David Ilcinkas, Amos Korman, and David Peleg. Label-guided graph exploration by a finite automaton. ACM Transactions on Algorithms, 4(4):42:1–42:??, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Carmo:2007:QPI

[CFK^{+}07] Renato Carmo, Tomás Feder, Yoshiharu Kohayakawa, Eduardo Laber, Rajeev Motwani, Liadan O’Callaghan, Rina Panigrahy, and Dilys Thomas. Querying priced information in databases: The conjunctive case. ACM Transactions on Algorithms, 3(1):??, February 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ciriani:2007:DSS

[CFLM07] Valentina Ciriani, Paolo Ferragina, Fabrizio Luccio, and S. Muthukrishnan. A data structure for a sequence of string accesses in external memory. ACM Transactions on Algorithms, 3(1):??, February 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chitnis:2021:PAA

[CFM21] Rajesh Chitnis, Andreas Emil Feldmann, and Pasin Manurangsi. Parameterized approximation algorithms for bidirected Steiner network problems. ACM Transactions on Algorithms, 17(2):12:1– 12:68, June 2021. CODEN ???? ISSN 1549-6325 (print), 1549-

6333 (electronic). URL https://dl.acm.org/doi/10.1145/
3447584.

Coppersmith:2010:OWN

[CFR10] Don Coppersmith, Lisa K. Fleischer, and Atri Rurda. Ordering by weighted number of wins gives a good ranking for weighted tournaments. ACM Transactions on Algorithms, 6(3):55:1–55:??, June 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cao:2022:EAT

[CFRY22] Nairen Cao, Jeremy T. Fineman, Katina Russell, and Eugene Yang. I/O-efficient algorithms for topological sort and related problems. ACM Transactions on Algorithms, 18(1):5:1–5:24, January 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3418356.

Cygan:2017:TKB

[CGH17] Marek Cygan, Fabrizio Grandoni, and Danny Hermelin. Tight kernel bounds for problems on graphs with small degeneracy.

ACM Transactions on Algorithms, 13(3):43:1–43:??, August 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cabello:2011:GCF

[CGK^{+}11] Sergio Cabello, Panos Giannopoulos, Christian Knauer, Dániel Marx, and Günter Rote. Geometric clustering: Fixed-parameter tractability and lower bounds with respect to the dimension. ACM Transactions on Algorithms, 7(4):43:1–43:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chakrabarty:2020:NUC

[CGK20] Deeparnab Chakrabarty, Prachi Goyal, and Ravishankar Krishnaswamy. The non-uniform k-center problem. ACM Transactions on Algorithms, 16(4):46:1–46:19, September 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/10.1145/3392720.

Chrobak:2022:SAO

[CGMY22] Marek Chrobak, Mordecai Golin, J. Ian Munro, and Neal E. Young. A simple algorithm for optimal search trees with two-way comparisons. ACM Transactions on Algorithms, 18(1):2:1–2:11, January 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3477910.

Chan:2016:HRD

[CGMZ16] T.-H. Hubert Chan, Anupam Gupta, Bruce M. Maggs, and Shuheng Zhou. On hierarchical routing in doubling metrics. ACM Transactions on Algorithms, 12(4):55:1–55:??, September 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chuzhoy:2008:ASN

[CGNS08] Julia Chuzhoy, Anupam Gupta, Joseph (Seffi) Naor, and Amitabh Sinha. On the approximability of some network design problems. ACM Transactions on Algorithms, 4(2):23:1–23:??, May 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chawla:2008:ENT

[CGR08] Shuchi Chawla, Anupam Gupta, and Harald Räcke. Embeddings of negative-type metrics and an improved approximation to generalized sparsest cut. ACM Transactions on Algorithms, 4(2): 22:1–22:??, May 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2010:CBT

[Cha10a] Timothy M. Chan. Comparison-based time-space lower bounds for selection. ACM Transactions on Algorithms, 6(2):26:1–26:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2010:BSP

[Cha10b] Timothy M. Chan. On the bichromatic k-set problem. ACM Transactions on Algorithms, 6(4):62:1–62:??, August 2010. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2012:APS

[Cha12] Timothy M. Chan. All-pairs shortest paths for unweighted undirected graphs in o(mn) time. ACM Transactions on Algorithms, 8(4):34:1–34:??, September 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2013:PPS

[Cha13] Timothy M. Chan. Persistent predecessor search and orthogonal point location on the word RAM. ACM Transactions on Algorithms, 9(3):22:1–22:??, June 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

CHAN:2018:IDA

[CHA18] Timothy M. CHAN. Improved deterministic algorithms for linear programming in low dimensions. ACM Transactions on Algorithms, 14(3):30:1–30:??, July 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3155312.

Chan:2020:MLF

[Cha20] Timothy M. Chan. More logarithmic-factor speedups for 3SUM, (median,+)-convolution, and some geometric 3SUM-hard problems. ACM Transactions on Algorithms, 16(1):7:1–7:23, January 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3363541.

Censor-Hillel:2017:TBV

[CHGG^{+}17] Keren Censor-Hillel, Mohsen Ghaffari, George Giakkoupis, Bernhard Haeupler, and Fabian Kuhn. Tight bounds on vertex connectivity under sampling. ACM Transactions on Algorithms, 13 (2):19:1–19:??, May 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2018:OSM

[CHJ^{+}18] T.-H. Hubert Chan, Zhiyi Huang, Shaofeng H.-C. Jiang, Ning Kang, and Zhihao Gavin Tang. Online submodular maximization with free disposal. ACM Transactions on Algorithms, 14 (4):1–29, October 2018. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3242770.

Chang:2020:DEC

[CHL^{+}20] Yi-Jun Chang, Qizheng He, Wenzheng Li, Seth Pettie, and Jara Uitto. Distributed edge coloring and a special case of the constructive Lovász local lemma. ACM Transactions on Algorithms, 16(1):8:1–8:51, January 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3365004.

Chan:2007:CID

[CHLS07] Ho-Leung Chan, Wing-Kai Hon, Tak-Wah Lam, and Kunihiko Sadakane. Compressed indexes for dynamic text collections. ACM Transactions on Algorithms, 3(2):21:1–21:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cole:2014:TDP

[CHLT14] Richard Cole, Carmit Hazay, Moshe Lewenstein, and Dekel Tsur. Two-dimensional parameterized matching. ACM Transactions on Algorithms, 11(2):12:1–12:??, October 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cheraghchi:2017:NOD

[CI17] Mahdi Cheraghchi and Piotr Indyk. Nearly optimal deterministic algorithm for sparse Walsh–Hadamard transform. ACM Transactions on Algorithms, 13(3):34:1–34:??, August 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2018:RCD

[CJ18] T.-H. Hubert Chan and Shaofeng H.-C. Jiang. Reducing curse of dimensionality: Improved PTAS for TSP (with neighborhoods) in doubling metrics. ACM Transactions on Algorithms, 14(1): 9:1–9:??, January 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2020:UPP

[CJJ20] T.-H. Hubert Chan, Haotian Jiang, and Shaofeng H.-C. Jiang. A unified PTAS for prize collecting TSP and Steiner tree problem in doubling metrics. ACM Transactions on Algorithms, 16(2): 24:1–24:23, April 2020. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3378571.

Czumaj:2024:SAG

[CJKV24] Artur Czumaj, Shaofeng H.-C. Jiang, Robert Krauthgamer, and Pavel Veselý. Streaming algorithms for geometric Steiner forest. ACM Transactions on Algorithms, 20(4):28:1–28:??, October 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3663666.

Cheng:2017:FSS

[CJL17] Siu-Wing Cheng, Jiongxin Jin, and Man-Kit Lau. A fast and simple surface reconstruction algorithm. ACM Transactions on Algorithms, 13(2):27:1–27:??, May 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chrobak:2007:IOA

[CJST07] Marek Chrobak, Wojciech Jawor, Jiří Sgall, and Tomáš Tichý. Improved online algorithms for buffer management in QoS switches. ACM Transactions on Algorithms, 3(4):50:1–50:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chekuri:2007:EDP

[CK07] Chandra Chekuri and Sanjeev Khanna. Edge-disjoint paths revisited. ACM Transactions on Algorithms, 3(4):46:1–46:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Caragiannis:2010:TLA

[CKK10a] Ioannis Caragiannis, Christos Kaklamanis, and Panagiotis Kanellopoulos. Taxes for linear atomic congestion games. ACM Transactions on Algorithms, 7(1):13:1–13:??, November 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Christodoulou:2010:MDF

[CKK10b] George Christodoulou, Elias Koutsoupias, and Annamária Kovács. Mechanism design for fractional scheduling on unrelated machines. ACM Transactions on Algorithms, 6(2):38:1–38:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Carr:2009:CCN

[CKL^{+}09] Robert D. Carr, Goran Konjevod, Greg Little, Venkatesh Natarajan, and Ojas Parekh. Compacting cuts: a new linear formulation for minimum cut. ACM Transactions on Algorithms, 5(3): 27:1–27:??, July 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Coester:2021:ISP

[CKL21a] Christian Coester, Elias Koutsoupias, and Philip Lazos. The infinite server problem. ACM Transactions on Algorithms, 17 (3):20:1–20:23, August 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3456632.

Cygan:2021:RCM

[CKL^{+}21b] Marek Cygan, Paweł Komosa, Daniel Lokshtanov, Marcin Pilipczuk, Michał Pilipczuk, Saket Saurabh, and Magnus Wahlström. Randomized contractions meet lean decompositions. ACM Transactions on Algorithms, 17(1):6:1–6:30, January 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3426738.

Chalermsook:2024:ASC

[CKM^{+}24] Parinya Chalermsook, Matthias Kaul, Matthias Mnich, Joachim Spoerhase, Sumedha Uniyal, and Daniel Vaz. Approximating sparsest cut in low-treewidth graphs via combinatorial diameter. ACM Transactions on Algorithms, 20(1):6:1–6:??, January 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3632623.

Chekuri:2012:IAO

[CKP12] Chandra Chekuri, Nitish Korula, and Martin Pál. Improved algorithms for orienteering and related problems. ACM Transactions on Algorithms, 8(3):23:1–23:??, July 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chang:2019:ESE

[CKP^{+}19] Yi-Jun Chang, Tsvi Kopelowitz, Seth Pettie, Ruosong Wang, and Wei Zhan. Exponential separations in the energy complexity of leader election. ACM Transactions on Algorithms, 15(4): 1–31, October 2019. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3341111.

Chlebus:2012:AQM

[CKR12] Bogdan S. Chlebus, Dariusz R. Kowalski, and Mariusz A. Rokicki. Adversarial queuing on the multiple access channel. ACM Transactions on Algorithms, 8(1):5:1–5:??, January 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chrobak:2005:GAM

[CKS05] Marek Chrobak, Petr Kolman, and Jiří Sgall. The greedy algorithm for the minimum common string partition problem. ACM Transactions on Algorithms, 1(2):350–366, October 2005. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chen:2009:OCF

[CKS09] Ke Chen, Haim Kaplan, and Micha Sharir. Online conflict-free coloring for halfplanes, congruent disks, and axis-parallel rectangles. ACM Transactions on Algorithms, 5(2):16:1–16:??, March 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cygan:2019:ITT

[CKS19] Marek Cygan, Łukasz Kowalik, and Arkadiusz Socała. Improving TSP tours using dynamic programming over tree decompositions. ACM Transactions on Algorithms, 15(4):1–19, October 2019. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3341730.

Cheng:2022:DDS

[CL22] Siu-Wing Cheng and Man-Kit Lau. Dynamic distributionsensitive point location. ACM Transactions on Algorithms, 18 (1):3:1–3:63, January 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3487403.

Clarkson:2010:CSG

[Cla10] Kenneth L. Clarkson. Coresets, sparse greedy approximation, and the Frank–Wolfe algorithm. ACM Transactions on Algorithms, 6(4):63:1–63:??, August 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chen:2012:IEC

[CLL^{+}12] Jianer Chen, Yang Liu, Songjian Lu, Sing-Hoi Sze, and Fenghui Zhang. Iterative expansion and color coding: An improved algorithm for 3D-matching. ACM Transactions on Algorithms, 8(1): 6:1–6:??, January 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cheung:2014:AAL

[CLL14] Ho Yee Cheung, Lap Chi Lau, and Kai Man Leung. Algebraic algorithms for linear matroid parity problems. ACM Transactions on Algorithms, 10(3):10:1–10:??, June 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chen:2008:MCI

[CLLJ08] Xin Chen, Lan Liu, Zheng Liu, and Tao Jiang. On the minimum common integer partition problem. ACM Transactions on Algorithms, 5(1):12:1–12:??, November 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cheriyan:2014:ARS

[CLNV14] Joseph Cheriyan, Bundit Laekhanukit, Guyslain Naves, and Adrian Vetta. Approximating rooted Steiner networks. ACM Transactions on Algorithms, 11(2):8:1–8:??, October 2014. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2022:NDE

[CLS^{+}22] Pak Hay Chan, Lap Chi Lau, Aaron Schild, Sam Chiu-Wai Wong, and Hong Zhou. Network design for s − t effective resistance. ACM Transactions on Algorithms, 18(3):22:1–22:??, July 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3522588.

Cormode:2007:SED

[CM07] Graham Cormode and S. Muthukrishnan. The string edit distance matching problem with moves. ACM Transactions on Algorithms, 3(1):??, February 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cao:2015:IDF

[CM15] Yixin Cao and Dániel Marx. Interval deletion is fixed-parameter tractable. ACM Transactions on Algorithms, 11(3):21:1–21:??, January 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cairo:2019:ONA

[CMA^{+}19] Massimo Cairo, Paul Medvedev, Nidia Obscura Acosta, Romeo Rizzi, and Alexandru I. Tomescu. An optimal O(nm) algorithm for enumerating all walks common to all closed edge-covering walks of a graph. ACM Transactions on Algorithms, 15(4): 1–17, October 2019. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3341731.

Charikar:2009:NOA

[CMM09] Moses Charikar, Konstantin Makarychev, and Yury Makarychev. Near-optimal algorithms for maximum constraint satisfaction problems. ACM Transactions on Algorithms, 5(3):32:1–32:??, July 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cardinal:2025:CGP

[CMM25] Jean Cardinal, Arturo Merino, and Torsten Mütze. Combinatorial generation via permutation languages. IV. Elimination trees. ACM Transactions on Algorithms, 21(1):13:1–13:??, January 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chandran:2008:IAO

[CMO^{+}08] Nishanth Chandran, Ryan Moriarty, Rafail Ostrovsky, Omkant Pandey, Mohammad Ali Safari, and Amit Sahai. Improved algorithms for optimal embeddings. ACM Transactions on Algorithms, 4(4):45:1–45:14, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2018:SSR

[CMR18] Timothy M. Chan, J. Ian Munro, and Venkatesh Raman. Selection and sorting in the “Restore” model. ACM Transactions on Algorithms, 14(2):11:1–11:??, June 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chekuri:2007:MDF

[CMS07] Chandra Chekuri, Marcelo Mydlarz, and F. Bruce Shepherd. Multicommodity demand flow in a tree and packing integer programs.

ACM Transactions on Algorithms, 3(3):27:1–27:??, August 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Charalampopoulos:2022:EDO

[CMT22] Panagiotis Charalampopoulos, Shay Mozes, and Benjamin Tebeka. Exact distance oracles for planar graphs with failing vertices. ACM Transactions on Algorithms, 18(2):18:1–18:23, April 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3511541.

Cheng:2016:FAC

[CMV16] Siu-Wing Cheng, Liam Mencel, and Antoine Vigneron. A faster algorithm for computing straight skeletons. ACM Transactions on Algorithms, 12(3):44:1–44:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chuzhoy:2016:AAH

[CMVZ16] Julia Chuzhoy, Yury Makarychev, Aravindan Vijayaraghavan, and Yuan Zhou. Approximation algorithms and hardness of the k-route cut problem. ACM Transactions on Algorithms, 12(1): 2:1–2:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cygan:2019:PEM

[CMWW19] Marek Cygan, Marcin Mucha, Karol Wegrzycki, and Michał Włodarczyk. On problems equivalent to (min, +)-convolution. ACM Transactions on Algorithms, 15(1):1–25, January 2019. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3293465.

Cormode:2011:ADF

[CMY11] Graham Cormode, S. Muthukrishnan, and Ke Yi. Algorithms for distributed functional monitoring. ACM Transactions on Algorithms, 7(2):21:1–21:??, March 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2014:FCC

[CN14] T.-H. Hubert Chan and Li Ning. Fast convergence for consensus in dynamic networks. ACM Transactions on Algorithms, 10(3): 15:1–15:??, June 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chakrabarty:2019:GCP

[CN19] Deeparnab Chakrabarty and Maryam Negahbani. Generalized center problems with outliers. ACM Transactions on Algorithms, 15(3):1–14, July 2019. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3338513.

Coppersmith:2011:OOG

[CNP^{+}11] Don Coppersmith, Tomasz Nowicki, Giuseppe Paleologo, Charles Tresser, and Chai Wah Wu. The optimality of the online greedy algorithm in carpool and chairman assignment problems. ACM Transactions on Algorithms, 7(3):37:1–37:??, July 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cygan:2022:SCP

[CNP^{+}22] Marek Cygan, Jesper Nederlof, Marcin Pilipczuk, Michał Pilipczuk, Johan M. M. Van Rooij, and Jakub Onufry Wojtaszczyk. Solving connectivity problems parameterized by treewidth in single exponential time. ACM Transactions on Algorithms, 18(2):17:1–17:31, April 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/

doi/10.1145/3506707.

Cohen:2018:SSF

[Coh18] Edith Cohen. Stream sampling framework and application for frequency cap statistics. ACM Transactions on Algorithms, 14 (4):1–40, October 2018. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3234338.

Cohen:2021:E

[Coh21] Edith Cohen. Editorial. ACM Transactions on Algorithms, 17(3):19e:1, August 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3462270.

Cygan:2012:EFE

[CP12] Marek Cygan and Marcin Pilipczuk. Even faster exact bandwidth. ACM Transactions on Algorithms, 8(1):8:1–8:??, January 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chapuy:2025:SSW

[CP25] Guillaume Chapuy and Guillem Perarnau. Short synchronizing words for random automata. ACM Transactions on Algorithms, 21(4):41:1–41:55, October 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Czyzowicz:2012:HMA

[CPL12] Jurek Czyzowicz, Andrzej Pelc, and Arnaud Labourel. How to meet asynchronously (almost) everywhere. ACM Transactions on Algorithms, 8(4):37:1–37:??, September 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chakaravarthy:2011:DTE

[CPR^{+}11] Venkatesan T. Chakaravarthy, Vinayaka Pandit, Sambuddha Roy, Pranjal Awasthi, and Mukesh K. Mohania. Decision trees for entity identification: Approximation algorithms and hardness results. ACM Transactions on Algorithms, 7(2):15:1–15:??, March 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chowdhury:2018:COB

[CR18] Rezaul A. Chowdhury and Vijaya Ramachandran. Cacheoblivious buffer heap and cache-efficient computation of shortest paths in graphs. ACM Transactions on Algorithms, 14(1): 1:1–1:??, January 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Cohen:2009:TDM

[CRR09] Rami Cohen, Dror Rawitz, and Danny Raz. Time-dependent multi-scheduling of multicast. ACM Transactions on Algorithms, 6(1):14:1–14:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chawla:2013:FSI

[CRR13] Shuchi Chawla, Prasad Raghavendra, and Dana Randall. Foreword to the Special Issue on SODA’11. ACM Transactions on Algorithms, 9(3):20:1–20:??, June 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chen:2022:LBC

[CRSS22] Xi Chen, Tim Randolph, Rocco A. Servedio, and Timothy Sun. A lower bound on cycle-finding in sparse digraphs. ACM Transactions on Algorithms, 18(4):31:1–31:??, October 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https:

//dl.acm.org/doi/10.1145/3417979.

Cairo:2024:GAP

[CRTZ24] Massimo Cairo, Romeo Rizzi, Alexandru I. Tomescu, and Elia C. Zirondelli. Genome assembly, from practice to theory: Safe, complete and linear-time. ACM Transactions on Algorithms, 20(1):4:1–4:??, January 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3632176.

Chung:2011:CDK

[CRV11] Kai-Min Chung, Omer Reingold, and Salil Vadhan. S-T connectivity on digraphs with a known stationary distribution. ACM Transactions on Algorithms, 7(3):30:1–30:??, July 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Carbonnel:2020:PWM

[CRZ20] Clément Carbonnel, Miguel Romero, and Stanislav Zivný. Pointwidth and Max-CSPs. ACM Transactions on Algorithms, 16(4): 54:1–54:28, September 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3409447.

Cheriyan:2007:PED

[CS07] Joseph Cheriyan and Mohammad R. Salavatipour. Packing element-disjoint Steiner trees. ACM Transactions on Algorithms, 3(4):47:1–47:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Czumaj:2008:TEM

[CS08] Artur Czumaj and Christian Sohler. Testing Euclidean minimum spanning trees in the plane. ACM Transactions on Algorithms, 4 (3):31:1–31:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Choi:2011:CPM

[CS11] Yongwook Choi and Wojciech Szpankowski. Constrained pattern matching. ACM Transactions on Algorithms, 7(2):25:1–25:??, March 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chudnovsky:2021:FSO

[CSS21] Maria Chudnovsky, Alex Scott, and Paul Seymour. Finding a shortest odd hole. ACM Transactions on Algorithms, 17 (2):13:1–13:21, June 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3447869.

Chen:2012:CRS

[CSTW12] Wei Chen, Christian Sommer, Shang-Hua Teng, and Yajun Wang. A compact routing scheme and approximate distance oracle for power-law graphs. ACM Transactions on Algorithms, 9(1):4:1– 4:??, December 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Czumaj:2007:TBW

[CV07] Artur Czumaj and Berthold Vöcking. Tight bounds for worst-case equilibria. ACM Transactions on Algorithms, 3(1):??, February 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chiplunkar:2020:RMA

[CV20] Ashish Chiplunkar and Sundar Vishwanathan. Randomized memoryless algorithms for the weighted and the generalized kserver problems. ACM Transactions on Algorithms, 16(1):14:1– 14:28, January 2020. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3365002.

Cook:2010:GFD

[CW10] Atlas F. Cook IV and Carola Wenk. Geodesic Fréchet distance inside a simple polygon. ACM Transactions on Algorithms, 7(1): 9:1–9:??, November 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chen:2015:CSP

[CW15] Danny Z. Chen and Haitao Wang. Computing shortest paths among curved obstacles in the plane. ACM Transactions on Algorithms, 11(4):26:1–26:??, June 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2016:AAO

[CW16] Timothy M. Chan and Bryan T. Wilkinson. Adaptive and approximate orthogonal range counting. ACM Transactions on Algorithms, 12(4):45:1–45:??, September 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chan:2021:DAO

[CW21] Timothy M. Chan and R. Ryan Williams. Deterministic APSP, orthogonal vectors, and more: Quickly derandomizing Razborov– Smolensky. ACM Transactions on Algorithms, 17(1):2:1–2:14, January 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3402926.

Chechik:2018:NOL

[CWN18] Shiri Chechik and Christian Wulff-Nilsen. Near-optimal light spanners. ACM Transactions on Algorithms, 14(3):33:1–33:??, July 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3199607.

Chrobak:2025:CTW

[CY25] Marek Chrobak and Neal E. Young. Classification via two-way comparisons. ACM Transactions on Algorithms, 21(2):17:1–17:??, April 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Chen:2018:PGI

[CZ18] Lin Chen and Guochuan Zhang. Packing groups of items into multiple knapsacks. ACM Transactions on Algorithms, 14(4): 1–24, October 2018. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3233524.

Chan:2024:HPL

[CZ24] Timothy M. Chan and Da Wei Zheng. Hopcroft’s problem, log* shaving, two-dimensional fractional cascading, and decision trees. ACM Transactions on Algorithms, 20(3):24:1–24:??, July 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3591357.

Daskalakis:2013:CAN

[Das13] Constantinos Daskalakis. On the complexity of approximating a Nash equilibrium. ACM Transactions on Algorithms, 9(3):23:1– 23:??, June 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

deBerg:2021:FGC

[dBBJW21] Mark de Berg, Kevin Buchin, Bart M. P. Jansen, and Gerhard Woeginger. Fine-grained complexity analysis of two classic TSP variants. ACM Transactions on Algorithms, 17(1):5:1–5:29, January 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3414845.

Carvalho:2005:VAE

[DC05] Marcelo H. De Carvalho and Joseph Cheriyan. An O(V E) algorithm for ear decompositions of matching-covered graphs. ACM Transactions on Algorithms, 1(2):324–337, October 2005. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Darwish:2021:MAN

[DEK21] Omar Darwish, Amr Elmasry, and Jyrki Katajainen. Memoryadjustable navigation piles with applications to sorting and convex hulls. ACM Transactions on Algorithms, 17(2):18:1–18:19, June 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3452938.

Demaine:2005:FPA

[DFHT05] Erik D. Demaine, Fedor V. Fomin, Mohammadtaghi Hajiaghayi, and Dimitrios M. Thilikos. Fixed-parameter algorithms for (k, r)- center in planar graphs and map graphs. ACM Transactions on Algorithms, 1(1):33–47, July 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Dutting:2025:FDS

[DFL^{+}25] Paul Dütting, Federico Fusco, Silvio Lattanzi, Ashkan Norouzi-Fard, and Morteza Zadimoghaddam. Fully dynamic submodular maximization over matroids. ACM Transactions on Algorithms, 21(1):11:1–11:??, January 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Deligkas:2023:PTA

[DFM23] Argyrios Deligkas, Michail Fasoulakis, and Evangelos Markakis. A polynomial-time algorithm for 1/3-approximate Nash equilibria in bimatrix games. ACM Transactions on Algorithms, 19 (4):31:1–31:??, October 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3606697.

Demetrescu:2009:TSP

[DFR09] Camil Demetrescu, Irene Finocchi, and Andrea Ribichini. Trading off space for passes in graph streaming problems. ACM Transactions on Algorithms, 6(1):6:1–6:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Doerr:2014:QRS

[DFS14] Benjamin Doerr, Tobias Friedrich, and Thomas Sauerwald. Quasirandom rumor spreading. ACM Transactions on Algorithms, 11(2):9:1–9:??, October 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Dyer:2020:RWS

[DGG^{+}20] Martin E. Dyer, Andreas Galanis, Leslie Ann Goldberg, Mark Jerrum, and Eric Vigoda. Random walks on small world networks. ACM Transactions on Algorithms, 16(3):37:1–37:33, June 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3382208.

DeBerg:2018:FAC

[DGM18] Mark De Berg, Joachim Gudmundsson, and Mehran Mehr. Faster algorithms for computing plurality points. ACM Transactions on Algorithms, 14(3):36:1–36:??, July 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3186990.

Deshpande:2012:PPF

[DH12] Amol Deshpande and Lisa Hellerstein. Parallel pipelined filter ordering with precedence constraints. ACM Transactions on Algorithms, 8(4):41:1–41:??, September 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Devanur:2018:PDG

[DH18] Nikhil R. Devanur and Zhiyi Huang. Primal dual gives almost optimal energy-efficient online algorithms. ACM Transactions on Algorithms, 14(1):5:1–5:??, January 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Demaine:2014:NWS

[DHK14] Erik D. Demaine, Mohammadtaghi Hajiaghayi, and Philip N. Klein. Node-weighted Steiner tree and group Steiner tree in planar graphs. ACM Transactions on Algorithms, 10(3):13:1–13:??, June 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Deshpande:2016:AAS

[DHK16] Amol Deshpande, Lisa Hellerstein, and Devorah Kletenik. Approximation algorithms for stochastic submodular set cover with applications to Boolean function evaluation and min-knapsack. ACM Transactions on Algorithms, 12(3):42:1–42:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Demaine:2009:MM

[DHM^{+}09] Erik D. Demaine, Mohammadtaghi Hajiaghayi, Hamid Mahini, Amin S. Sayedi-Roshkhar, Shayan Oveisgharan, and Morteza Zadimoghaddam. Minimizing movement. ACM Transactions on Algorithms, 5(3):30:1–30:??, July 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Dell:2014:ETC

[DHM^{+}14a] Holger Dell, Thore Husfeldt, Dániel Marx, Nina Taslaman, and Martin Wahlén. Exponential time complexity of the permanent and the Tutte polynomial. ACM Transactions on Algorithms, 10 (4):21:1–21:??, August 2014. ISSN 1549-6325 (print), 1549-6333 (electronic).

Demaine:2014:MMF

[DHM14b] Erik D. Demaine, Mohammadtaghi Hajiaghayi, and Dániel Marx. Minimizing movement: Fixed-parameter tractability. ACM Transactions on Algorithms, 11(2):14:1–14:??, October 2014. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Demaine:2012:PAN

[DHMZ12] Erik D. Demaine, Mohammadtaghi Hajiaghayi, Hamid Mahini, and Morteza Zadimoghaddam. The price of anarchy in network creation games. ACM Transactions on Algorithms, 8(2):13:1– 13:??, April 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Driemel:2016:ECV

[DHPR16] Anne Driemel, Sariel Har-Peled, and Benjamin Raichel. On the expected complexity of Voronoi diagrams on terrains. ACM Transactions on Algorithms, 12(3):37:1–37:??, June 2016. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

DeCarliSilva:2016:SSP

[DHS16] Marcel K. De Carli Silva, Nicholas J. A. Harvey, and Cristiane M. Sato. Sparse sums of positive semidefinite matrices. ACM Transactions on Algorithms, 12(1):9:1–9:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Demetrescu:2006:EAD

[DI06] Camil Demetrescu and Giuseppe F. Italiano. Experimental analysis of dynamic all pairs shortest path algorithms. ACM Transactions on Algorithms, 2(4):578–601, October 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Dallant:2025:CLB

[DI25] Justin Dallant and John Iacono. Conditional lower bounds for dynamic geometric measure problems. ACM Transactions on Algorithms, 21(2):23:1–23:??, April 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Demaine:2007:RDS

[DIL07] Erik D. Demaine, John Iacono, and Stefan Langerman. Retroactive data structures. ACM Transactions on Algorithms, 3(2): 13:1–13:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Djidjev:2010:FAC

[Dji10] Hristo N. Djidjev. A faster algorithm for computing the girth of planar and bounded genus graphs. ACM Transactions on Algorithms, 7(1):3:1–3:??, November 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Dabrowski:2025:ACS

[DJO^{+}25] Konrad K. Dabrowski, Peter Jonsson, Sebastian Ordyniak, George Osipov, and Magnus Wahlström. Almost consistent systems of linear equations. ACM Transactions on Algorithms, 21 (4):44:1–44:55, October 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Diedrich:2012:TAA

[DJP^{+}12] Florian Diedrich, Klaus Jansen, Lars Prädel, Ulrich M. Schwarz, and Ola Svensson. Tight approximation algorithms for scheduling with fixed jobs and nonavailability. ACM Transactions on Algorithms, 8(3):27:1–27:??, July 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Drmota:2012:PAC

[DK12] Michael Drmota and Reinhard Kutzelnigg. A precise analysis of Cuckoo hashing. ACM Transactions on Algorithms, 8(2):11:1– 11:36, April 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Duncan:2006:OCG

[DKK06] Christian A. Duncan, Stephen G. Kobourov, and V. S. Anil Kumar. Optimal constrained graph exploration. ACM Transactions on Algorithms, 2(3):380–402, July 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Dinitz:2017:IAA

[DKN17] Michael Dinitz, Guy Kortsarz, and Zeev Nutov. Improved approximation algorithm for Steiner k-forest with nearly uniform weights. ACM Transactions on Algorithms, 13(3):40:1–40:??, August 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Dinitz:2016:LCI

[DKR16] Michael Dinitz, Guy Kortsarz, and Ran Raz. Label cover instances with large girth and the hardness of approximating basic k-spanner. ACM Transactions on Algorithms, 12(2):25:1–25:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Dvorak:2011:TCT

[DKT11] Zdenek Dvorák, Ken-Ichi Kawarabayashi, and Robin Thomas. Three-coloring triangle-free planar graphs in linear time. ACM Transactions on Algorithms, 7(4):41:1–41:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Dom:2014:KLB

[DLS14] Michael Dom, Daniel Lokshtanov, and Saket Saurabh. Kernelization lower bounds through colors and IDs. ACM Transactions on Algorithms, 11(2):13:1–13:??, October 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Duch:2009:URK

[DM09] Amalia Duch and Conrado Martínez. Updating relaxed K-d trees. ACM Transactions on Algorithms, 6(1):4:1–4:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Damerow:2012:SAL

[DMM^{+}12] Valentina Damerow, Bodo Manthey, Friedhelm Meyer Auf Der Heide, Harald Räcke, Christian Scheideler, Christian Sohler, and Till Tantau. Smoothed analysis of left-to-right maxima with applications. ACM Transactions on Algorithms, 8(3):30:1–30:??, July 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Demaine:2009:ODA

[DMRW09] Erik D. Demaine, Shay Mozes, Benjamin Rossman, and Oren Weimann. An optimal decomposition algorithm for tree edit distance. ACM Transactions on Algorithms, 6(1):2:1–2:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Dadush:2024:IAS

[DMT24] Daniel Dadush, Martin Milanic, and Tami Tamir. Introduction: ACM-SIAM Symposium on Discrete Algorithms (SODA) 2022 special issue. ACM Transactions on Algorithms, 20(3):21:1–21:??, July 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3655622.

Drmota:2006:RFA

[DP06] Michael Drmota and Helmut Prodinger. The register function for t-ary trees. ACM Transactions on Algorithms, 2(3):318–334, July 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Dieudonne:2014:DNE

[DP14] Yoann Dieudonné and Andrzej Pelc. Deterministic network exploration by anonymous silent agents with local traffic reports. ACM Transactions on Algorithms, 11(2):10:1–10:??, October 2014. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Dieudonne:2014:GDM

[DPP14] Yoann Dieudonné, Andrzej Pelc, and David Peleg. Gathering despite mischief. ACM Transactions on Algorithms, 11(1):1:1– 1:??, August 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Duan:2018:SAW

[DPS18] Ran Duan, Seth Pettie, and Hsin-Hao Su. Scaling algorithms for weighted matching in general graphs. ACM Transactions on Algorithms, 14(1):8:1–8:??, January 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Dinitz:2008:OAS

[DS08] Yefim Dinitz and Shay Solomon. Optimality of an algorithm solving the Bottleneck Tower of Hanoi problem. ACM Transactions on Algorithms, 4(3):25:1–25:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

DeVerdiere:2011:SVD

[DS11] Éric Colin De Verdière and Alexander Schrijver. Shortest vertexdisjoint two-face paths in planar graphs. ACM Transactions on Algorithms, 7(2):19:1–19:??, March 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Disser:2019:SAN

[DS19] Yann Disser and Martin Skutella. The simplex algorithm is NP-Mighty. ACM Transactions on Algorithms, 15(1):1–19, January 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3280847.

Dumitrescu:2016:TSP

[DT16] Adrian Dumitrescu and Csaba D. Tóth. The traveling salesman problem for lines, balls, and planes. ACM Transactions on Algorithms, 12(3):43:1–43:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Drescher:2010:AAM

[DV10] Matthew Drescher and Adrian Vetta. An approximation algorithm for the maximum leaf spanning arborescence problem. ACM Transactions on Algorithms, 6(3):46:1–46:??, June 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Even-Dar:2007:CTN

[EDKM07] Eyal Even-Dar, Alex Kesselman, and Yishay Mansour. Convergence time to Nash equilibrium in load balancing. ACM Transactions on Algorithms, 3(3):32:1–32:??, August 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ebrahimi:2012:CAW

[EF12] Javad B. Ebrahimi and Christina Fragouli. Combinatiorial algorithms for wireless information flow. ACM Transactions on Algorithms, 9(1):8:1–8:??, December 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ehsani:2015:BBN

[EFF^{+}15] Shayan Ehsani, Saber Shokat Fadaee, Mohammadamin Fazli, Abbas Mehrabian, Sina Sadeghian Sadeghabad, Mohammadali Safari, and Morteza Saghafian. A bounded budget network creation game. ACM Transactions on Algorithms, 11(4):34:1–34:??, June 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Even:2009:AAA

[EFKN09] Guy Even, Jon Feldman, Guy Kortsarz, and Zeev Nutov. A 1.8 approximation algorithm for augmenting edge-connectivity of a graph from 1 to 2. ACM Transactions on Algorithms, 5(2):21:1– 21:??, March 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Efrat:2016:IAA

[EFM^{+}16] Alon Efrat, Sándor P. Fekete, Joseph S. B. Mitchell, Valentin Polishchuk, and Jukka Suomela. Improved approximation algorithms for relay placement. ACM Transactions on Algorithms, 12(2):20:1–20:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Esfandiari:2018:SAE

[EHL^{+}18] Hossein Esfandiari, Mohammadtaghi Hajiaghayi, Vahid Liaghat, Morteza Monemizadeh, and Krzysztof Onak. Streaming algorithms for estimating the matching size in planar graphs and beyond. ACM Transactions on Algorithms, 14(4):1–23, October 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3230819.

Emek:2016:SCI

[EHR16] Yuval Emek, Magnús M. Halldórsson, and Adi Rosén. Spaceconstrained interval selection. ACM Transactions on Algorithms, 12(4):51:1–51:??, September 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Elmasry:2008:MPQ

[EJK08] Amr Elmasry, Claus Jensen, and Jyrki Katajainen. Multipartite priority queues. ACM Transactions on Algorithms, 5(1):14:1– 14:??, November 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Evans:2006:OSV

[EK06] William Evans and David Kirkpatrick. Optimally scheduling video-on-demand to minimize delay when sender and receiver bandwidth may differ. ACM Transactions on Algorithms, 2(4): 661–678, October 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Elkin:2007:IAR

[EK07] Michael Elkin and Guy Kortsarz. An improved algorithm for radio broadcast. ACM Transactions on Algorithms, 3(1):??, February 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Eiben:2020:CPP

[EK20] Eduard Eiben and Iyad Kanj. A colored path problem and its applications. ACM Transactions on Algorithms, 16(4):47:1–47:48, September 2020. CODEN ???? ISSN 1549-6325 (print), 1549-

6333 (electronic). URL https://dl.acm.org/doi/10.1145/
3396573.

Even:2005:NDP

[EKS05] Guy Even, Guy Kortsarz, and Wolfgang Slany. On network design problems: fixed cost flows and the covering Steiner problem. ACM Transactions on Algorithms, 1(1):74–101, July 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Elkin:2005:CAS

[Elk05] Michael Elkin. Computing almost shortest paths. ACM Transactions on Algorithms, 1(2):283–323, October 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Elkin:2011:SFD

[Elk11] Michael Elkin. Streaming and fully dynamic centralized algorithms for constructing and maintaining sparse spanners. ACM Transactions on Algorithms, 7(2):20:1–20:??, March 2011. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Elmasry:2017:TOS

[Elm17] Amr Elmasry. Toward optimal self-adjusting heaps. ACM Transactions on Algorithms, 13(4):55:1–55:??, December 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Even:2021:SRA

[ELMR21] Guy Even, Reut Levi, Moti Medina, and Adi Rosén. Sublinear random access generators for preferential attachment graphs. ACM Transactions on Algorithms, 17(4):28:1–28:26, October 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3464958.

Ene:2025:IAS

[ELMS25] Alina Ene, Troy Lee, Piotr Micek, and Sushant Sachdeva. Introduction: ACM–SIAM Symposium on Discrete Algorithms (SODA) 2021 special issue. ACM Transactions on Algorithms, 21 (3):29:1–29:2, July 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Even:2008:ACR

[ELR^{+}08] Guy Even, Retsef Levi, Dror Rawitz, Baruch Schieber, Shimon (Moni) Shahar, and Maxim Sviridenko. Algorithms for capacitated rectangle stabbing and lot sizing with joint set-up costs. ACM Transactions on Algorithms, 4(3):34:1–34:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Eden:2020:TBA

[ELR20] Talya Eden, Reut Levi, and Dana Ron. Testing bounded arboricity. ACM Transactions on Algorithms, 16(2):18:1–18:22, April 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3381418.

Ergun:2010:PTS

[EMS10] Funda Ergun, S. Muthukrishnan, and Cenk Sahinalp. Periodicity testing with sublinear samples and space. ACM Transactions on Algorithms, 6(2):43:1–43:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Eberle:2023:OTM

[EMS23] Franziska Eberle, Nicole Megow, and Kevin Schewior. Online throughput maximization on unrelated machines: Commitment is no burden. ACM Transactions on Algorithms, 19(1):10:1–10:??, January 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3569582.

Equi:2023:CSM

[EMTG23] Massimo Equi, Veli Mäkinen, Alexandru I. Tomescu, and Roberto Grossi. On the complexity of string matching for graphs. ACM Transactions on Algorithms, 19(3):21:1–21:??, July 2023. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3588334.

Elkin:2019:EAC

[EN19] Michael Elkin and Ofer Neiman. Efficient algorithms for constructing very sparse spanners and emulators. ACM Transactions on Algorithms, 15(1):1–29, January 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3274651.

Edmonds:2005:MAL

[EP05] Jeff Edmonds and Kirk Pruhs. A maiden analysis of longest wait first. ACM Transactions on Algorithms, 1(1):14–32, July 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Edmonds:2011:CCR

[EP11] Jeff Edmonds and Kirk Pruhs. Cake cutting really is not a piece of cake. ACM Transactions on Algorithms, 7(4):51:1–51:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Edmonds:2012:SSP

[EP12] Jeff Edmonds and Kirk Pruhs. Scalably scheduling processes with arbitrary speedup curves. ACM Transactions on Algorithms, 8 (3):28:1–28:??, July 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Elkin:2016:LSL

[EP16] Michael Elkin and Seth Pettie. A linear-size logarithmic stretch path-reporting distance oracle for general graphs. ACM Transactions on Algorithms, 12(4):50:1–50:??, September 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Eppstein:2006:QAM

[Epp06] David Eppstein. Quasiconvex analysis of multivariate recurrence equations for backtracking algorithms. ACM Transactions on Algorithms, 2(4):492–509, October 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Eppstein:2007:FSI

[Epp07] David Eppstein. Foreword to special issue on SODA 2002. ACM Transactions on Algorithms, 3(1):??, February 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Eppstein:2009:AMI

[Epp09a] David Eppstein. All maximal independent sets and dynamic dominance for sparse graphs. ACM Transactions on Algorithms, 5(4): 38:1–38:??, October 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Eppstein:2009:STS

[Epp09b] David Eppstein. Squarepants in a tree: Sum of subtree clustering and hyperbolic pants decomposition. ACM Transactions on Algorithms, 5(3):29:1–29:??, July 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Eppstein:2009:TBG

[Epp09c] David Eppstein. Testing bipartiteness of geometric intersection graphs. ACM Transactions on Algorithms, 5(2):15:1–15:??, March 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Eppstein:2018:PCP

[Epp18] David Eppstein. The parametric closure problem. ACM Transactions on Algorithms, 14(1):2:1–2:??, January 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Emek:2010:NLT

[EPR10] Yuval Emek, David Peleg, and Liam Roditty. A near-linear-time algorithm for computing replacement paths in planar directed graphs. ACM Transactions on Algorithms, 6(4):64:1–64:??, August 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Eisenbrand:2013:PDP

[EPR13] Friedrich Eisenbrand, Dömötör Pálvölgyi, and Thomas Rothvoß. Bin packing via discrepancy of permutations. ACM Transactions on Algorithms, 9(3):24:1–24:??, June 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Emek:2016:SSS

[ER16] Yuval Emek and Adi Rosén. Semi-streaming set cover. ACM Transactions on Algorithms, 13(1):6:1–6:??, December 2016. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Etscheid:2017:SAL

[ER17] Michael Etscheid and Heiko Röglin. Smoothed analysis of local search for the maximum-cut problem. ACM Transactions on Algorithms, 13(2):25:1–25:??, May 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Englert:2016:SAO

[ERV16] Matthias Englert, Heiko Röglin, and Berthold Vöcking. Smoothed analysis of the 2-opt algorithm for the general TSP. ACM Transactions on Algorithms, 13(1):10:1–10:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Elkin:2016:FCL

[ES16] Michael Elkin and Shay Solomon. Fast constructions of lightweight spanners for general graphs. ACM Transactions on Algorithms, 12(3):29:1–29:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Epstein:2006:OSS

[EV06a] Leah Epstein and Rob Van Stee. Online scheduling of splittable tasks. ACM Transactions on Algorithms, 2(1):79–94, January 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Epstein:2006:SU

[EV06b] Leah Epstein and Rob Van Stee. This side up! ACM Transactions on Algorithms, 2(2):228–243, April 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Epstein:2010:OUC

[EV10] Leah Epstein and Rob Van Stee. On the online unit clustering problem. ACM Transactions on Algorithms, 7(1):7:1–7:??, November 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Eisenbrand:2020:PRF

[EW20] Friedrich Eisenbrand and Robert Weismantel. Proximity results and faster algorithms for integer programming using the Steinitz lemma. ACM Transactions on Algorithms, 16(1):5:1– 5:14, January 2020. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3340322.

Fernandez-Baca:2009:PAU

[FBV09] David Fernández-Baca and Balaji Venkatachalam. Parametric analysis for ungapped Markov models of evolution. ACM Transactions on Algorithms, 5(4):44:1–44:??, October 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Farach-Colton:2009:BHO

[FCFM09] Martin Farach-Colton, Rohan J. Fernandes, and Miguel A. Mosteiro. Bootstrapping a hop-optimal network in the weak sensor model. ACM Transactions on Algorithms, 5(4):37:1–37:??, October 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Feldman:2017:MSS

[Fel17] Moran Feldman. Maximizing symmetric submodular functions. ACM Transactions on Algorithms, 13(3):39:1–39:??, August 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Filtser:2023:SSD

[FF23] Arnold Filtser and Omrit Filtser. Static and streaming data structures for Fréchet distance queries. ACM Transactions on Algorithms, 19(4):39:1–39:??, October 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3610227.

Fanelli:2012:SCC

[FFM12] Angelo Fanelli, Michele Flammini, and Luca Moscardelli. The speed of convergence in congestion games under best-response dynamics. ACM Transactions on Algorithms, 8(3):25:1–25:??, July 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Franceschini:2008:NSB

[FG08] Gianni Franceschini and Roberto Grossi. No sorting? better searching! ACM Transactions on Algorithms, 4(1):2:1–2:??, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Faour:2025:LDR

[FGG^{+}25] Salwa Faour, Mohsen Ghaffari, Christoph Grunau, Fabian Kuhn, and Václav Rozhoň. Local distributed rounding: Generalized to MIS, matching, set cover, and beyond. ACM Transactions on Algorithms, 21(4):42:1–42:48, October 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Foschini:2006:WIE

[FGGV06] Luca Foschini, Roberto Grossi, Ankur Gupta, and Jeffrey Scott Vitter. When indexing equals compression: Experiments with compressing suffix arrays and applications. ACM Transactions on Algorithms, 2(4):611–639, October 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Finocchi:2009:RD

[FGI09] Irene Finocchi, Fabrizio Grandoni, and Giuseppe F. Italiano. Resilient dictionaries. ACM Transactions on Algorithms, 6(1):1:1– 1:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Fleischer:2016:SEA

[FGK^{+}16] Lisa Fleischer, Rahul Garg, Sanjiv Kapoor, Rohit Khandekar, and Amin Saberi. A simple and efficient algorithm for computing market equilibria. ACM Transactions on Algorithms, 12(3):34:1– 34:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Fomin:2024:SCM

[FGK^{+}24a] Fedor V. Fomin, Petr A. Golovach, Tuukka Korhonen, Daniel Lokshtanov, and Giannos Stamoulis. Shortest cycles with monotone submodular costs. ACM Transactions on Algorithms, 20 (1):2:1–2:??, January 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3626824.

Fomin:2024:FPT

[FGK^{+}24b] Fedor V. Fomin, Petr A. Golovach, Tuukka Korhonen, Kirill Simonov, and Giannos Stamoulis. Fixed-parameter tractability of maximum colored path and beyond. ACM Transactions on Algorithms, 20(4):36:1–36:??, October 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3674835.

Fried:2024:IAD

[FGK^{+}24c] Dvir Fried, Shay Golan, Tomasz Kociumaka, Tsvi Kopelowitz, Ely Porat, and Tatiana Starikovskaya. An improved algorithm for the k-Dyck edit distance problem. ACM Transactions on Algorithms, 20(3):26:1–26:??, July 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/

doi/10.1145/3627539.

Fomin:2019:CWI

[FGL^{+}19] Fedor V. Fomin, Petr A. Golovach, Daniel Lokshtanov, Saket Saurabh, and Meirav Zehavi. Clique-width III: Hamiltonian cycle and the odd case of graph coloring. ACM Transactions on Algorithms, 15(1):1–27, January 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3280824.

Fomin:2020:ASL

[FGL^{+}20] Fedor V. Fomin, Petr A. Golovach, Daniel Lokshtanov, Fahad Panolan, and Saket Saurabh. Approximation schemes for lowrank binary matrix approximation problems. ACM Transactions on Algorithms, 16(1):12:1–12:39, January 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3365653.

Fomin:2019:SCR

[FGLS19] Fedor V. Fomin, Petr A. Golovach, Daniel Lokshtanov, and Saket Saurabh. Spanning circuits in regular matroids. ACM Transactions on Algorithms, 15(4):1–38, October 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3355629.

Fomin:2008:CBM

[FGPS08] Fedor V. Fomin, Fabrizio Grandoni, Artem V. Pyatkin, and Alexey A. Stepanov. Combinatorial bounds via measure and conquer: Bounding minimal dominating sets and applications. ACM Transactions on Algorithms, 5(1):9:1–9:??, November 2008. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Focke:2025:ACA

[FGRZ25] Jacob Focke, Leslie Ann Goldberg, Marc Roth, and Stanislav Zivný. Approximately counting answers to conjunctive queries with disequalities and negations. ACM Transactions on Algorithms, 21(1):8:1–8:??, January 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Feng:2022:RMS

[FGYZ22] Weiming Feng, Heng Guo, Yitong Yin, and Chihao Zhang. Rapid mixing from spectral independence beyond the Boolean domain. ACM Transactions on Algorithms, 18(3):28:1–28:??, July 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3531008.

Focke:2021:CAC

[FGZ21] Jacob Focke, Leslie Ann Goldberg, and Stanislav Zivný. The complexity of approximately counting retractions to square-free graphs. ACM Transactions on Algorithms, 17(3):22:1–22:51, August 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3458040.

Fakcharoenphol:2007:TRP

[FHR07] Jittat Fakcharoenphol, Chris Harrelson, and Satish Rao. The ktraveling repairmen problem. ACM Transactions on Algorithms, 3(4):40:1–40:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Filtser:2024:SSP

[Fil24] Arnold Filtser. Scattering and sparse partitions, and their applications. ACM Transactions on Algorithms, 20(4):30:1–30:??, October 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3672562.

Filtser:2025:FCP

[Fil25] Arnold Filtser. A face cover perspective to l_{1} embeddings of planar graphs. ACM Transactions on Algorithms, 21(1):4:1–4:??, January 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Feigenbaum:2006:SMC

[FIM^{+}06] Joan Feigenbaum, Yuval Ishai, Tal Malkin, Kobbi Nissim, Martin J. Strauss, and Rebecca N. Wright. Secure multiparty computation of approximations. ACM Transactions on Algorithms, 2 (3):435–472, July 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Feigenbaum:2014:APF

[FJS14] Joan Feigenbaum, Aaron D. Jaggard, and Michael Schapira. Approximate privacy: Foundations and quantification. ACM Transactions on Algorithms, 10(3):11:1–11:??, June 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Friggstad:2019:ASC

[FKRS19] Zachary Friggstad, Kamyar Khodamoradi, Mohsen Rezapour, and Mohammad R. Salavatipour. Approximation schemes for clustering with outliers. ACM Transactions on Algorithms, 15 (2):1–26, May 2019. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3301446.

Fotakis:2008:ACG

[FKS08] Dimitris Fotakis, Spyros Kontogiannis, and Paul Spirakis. Atomic congestion games among coalitions. ACM Transactions on Algorithms, 4(4):52:1–52:??, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Fu:2011:DAH

[FKW11] Bin Fu, Ming-Yang Kao, and Lusheng Wang. Discovering almost any hidden motif from multiple sequences. ACM Transactions on Algorithms, 7(2):26:1–26:??, March 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Feldmann:2025:PAS

[FL25] Andreas Emil Feldmann and Michael Lampis. Parameterized algorithms for Steiner forest in bounded width graphs. ACM Transactions on Algorithms, 21(4):47:1–47:26, October 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Fomin:2020:SAR

[FLK^{+}20] Fedor V. Fomin, Daniel Lokshtanov, Sudeshna Kolay, Fahad Panolan, and Saket Saurabh. Subexponential algorithms for rectilinear Steiner tree and arborescence problems. ACM Transactions on Algorithms, 16(2):21:1–21:37, April 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3381420.

Fomin:2019:SKI

[FLL^{+}19] Fedor V. Fomin, Tien-Nam Le, Daniel Lokshtanov, Saket Saurabh, Stéphan Thomassé, and Meirav Zehavi. Subquadratic kernels for implicit 3-hitting set and 3-set packing problems. ACM Transactions on Algorithms, 15(1):1–44, January 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https:

//dl.acm.org/doi/abs/10.1145/3293466.

Fischer:2012:QCT

[FLM^{+}12] Eldar Fischer, Oded Lachish, Arie Matsliah, Ilan Newman, and Orly Yahalom. On the query complexity of testing orientations for being Eulerian. ACM Transactions on Algorithms, 8(2):15:1– 15:??, April 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Fakcharoenphol:2014:FAS

[FLN14] Jittat Fakcharoenphol, Bundit Laekhanukit, and Danupon Nanongkai. Faster algorithms for semi-matching problems. ACM Transactions on Algorithms, 10(3):14:1–14:??, June 2014. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Frigo:2012:COA

[FLPR12] Matteo Frigo, Charles E. Leiserson, Harald Prokop, and Sridhar Ramachandran. Cache-oblivious algorithms. ACM Transactions on Algorithms, 8(1):4:1–4:??, January 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Fomin:2017:RFP

[FLPS17] Fedor V. Fomin, Daniel Lokshtanov, Fahad Panolan, and Saket Saurabh. Representative families of product families. ACM Transactions on Algorithms, 13(3):36:1–36:??, August 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Fomin:2018:FPT

[FLS^{+}18] Fedor V. Fomin, Daniel Lokshtanov, Saket Saurabh, MichaL Pilipczuk, and Marcin Wrochna. Fully polynomial-time parameterized computations for graphs and matrices of low treewidth.

ACM Transactions on Algorithms, 14(3):34:1–34:??, July 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/ft_gateway.cfm?id=3186898.

Fomin:2018:KCD

[FLST18] Fedor V. Fomin, Daniel Lokshtanov, Saket Saurabh, and Dimitrios M. Thilikos. Kernels for (connected) dominating set on graphs with excluded topological minors. ACM Transactions on Algorithms, 14(1):6:1–6:??, January 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Friggstad:2025:ADS

[FM25] Zachary Friggstad and Ramin Mousavi. A O(log, k)-approximation for directed Steiner tree in planar graphs. ACM Transactions on Algorithms, 21(4):45:1–45:14, October 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Focke:2025:TCB

[FMM^{+}25] Jacob Focke, Dániel Marx, Fionn Mc Inerney, Daniel Neuen, Govind S. Sankar, Philipp Schepper, and Philip Wellnitz. Tight complexity bounds for counting generalized dominating sets in bounded-treewidth graphs — Part I: Algorithmic results. ACM Transactions on Algorithms, 21(3):27:1–27:45, July 2025. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ferragina:2007:CRS

[FMMN07] Paolo Ferragina, Giovanni Manzini, Veli Mäkinen, and Gonzalo Navarro. Compressed representations of sequences and full-text indexes. ACM Transactions on Algorithms, 3(2):20:1–20:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Focke:2024:CLH

[FMR24] Jacob Focke, Dániel Marx, and Paweł Rzazewski. Counting list homomorphisms from graphs of bounded treewidth: Tight complexity bounds. ACM Transactions on Algorithms, 20(2):11:1– 11:??, April 2024. CODEN ???? ISSN 1549-6325 (print), 1549-

6333 (electronic). URL https://dl.acm.org/doi/10.1145/
3640814.

Feldman:2010:DSS

[FMS^{+}10] Jon Feldman, S. Muthukrishnan, Anastasios Sidiropoulos, Cliff Stein, and Zoya Svitkina. On distributing symmetric streaming computations. ACM Transactions on Algorithms, 6(4):66:1–66:??, August 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Flammini:2010:BSP

[FN10] Michele Flammini and Gaia Nicosia. On the bicriteria kserver problem. ACM Transactions on Algorithms, 7(1):6:1–6:??, November 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Fischer:2020:TAP

[FN20] Manuela Fischer and Andreas Noever. Tight analysis of parallel randomized greedy MIS. ACM Transactions on Algorithms, 16(1):6:1–6:13, January 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3326165.

Fotakis:2011:MFL

[Fot11] Dimitris Fotakis. Memoryless facility location in one pass. ACM Transactions on Algorithms, 7(4):49:1–49:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Faigle:2010:TPG

[FP10] Ulrich Faigle and Britta Peis. Two-phase greedy algorithms for some classes of combinatorial linear programs. ACM Transactions on Algorithms, 6(4):65:1–65:??, August 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Fraigniaud:2013:DIE

[FP13] Pierre Fraigniaud and Andrzej Pelc. Delays induce an exponential memory gap for rendezvous in trees. ACM Transactions on Algorithms, 9(2):17:1–17:??, March 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Fearnley:2022:FAF

[FPS22] John Fearnley, Dömötör Pálvölgyi, and Rahul Savani. A faster algorithm for finding Tarski fixed points. ACM Transactions on Algorithms, 18(3):23:1–23:??, July 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3524044.

Fox:2023:MCM

[FPZ23] Kyle Fox, Debmalya Panigrahi, and Fred Zhang. Minimum cut and minimum k-cut in hypergraphs via branching contractions. ACM Transactions on Algorithms, 19(2):13:1–13:??, April 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3570162.

Fattal:2010:ADM

[FR10] Shahar Fattal and Dana Ron. Approximating the distance to monotonicity in high dimensions. ACM Transactions on Algorithms, 6(3):52:1–52:??, June 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Friggstad:2011:MMM

[FS11] Zachary Friggstad and Mohammad R. Salavatipour. Minimizing movement in mobile facility location problems. ACM Transactions on Algorithms, 7(3):28:1–28:??, July 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Fusy:2008:DOT

[FSP08] Éric Fusy, Gilles Schaeffer, and Dominique Poulalhon. Dissections, orientations, and trees with applications to optimal mesh encoding and random sampling. ACM Transactions on Algorithms, 4(2):19:1–19:??, May 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Fischer:2020:DSS

[FTK20] Johannes Fischer, Tomohiro I, and Dominik Köppl. Deterministic sparse suffix sorting in the restore model. ACM Transactions on Algorithms, 16(4):50:1–50:53, September 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/10.1145/3398681.

Fujito:2012:HTM

[Fuj12] Toshihiro Fujito. How to trim a MST: a 2-approximation algorithm for minimum cost-tree cover. ACM Transactions on Algorithms, 8(2):16:1–16:??, April 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Fukunaga:2017:SCP

[Fuk17] Takuro Fukunaga. Spider covers for prize-collecting network activation problem. ACM Transactions on Algorithms, 13(4):49:1– 49:??, December 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ferragina:2010:CPI

[FV10] Paolo Ferragina and Rossano Venturini. The compressed permuterm index. ACM Transactions on Algorithms, 7(1):10:1– 10:??, November 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ferragina:2016:CCO

[FV16] Paolo Ferragina and Rossano Venturini. Compressed cacheoblivious string B-tree. ACM Transactions on Algorithms, 12 (4):52:1–52:??, September 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ferraioli:2019:MLD

[FV19] Diodato Ferraioli and Carmine Ventre. Metastability of the logit dynamics for asymptotically well-behaved potential games. ACM Transactions on Algorithms, 15(2):1–42, May 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3301315.

Feng:2007:FAS

[FZ07] Jianxing Feng and Daming Zhu. Faster algorithms for sorting by transpositions and sorting by block interchanges. ACM Transactions on Algorithms, 3(3):25:1–25:14, August 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gabow:2005:EF

[Gab05] Harold N. Gabow. Editor’s foreword. ACM Transactions on Algorithms, 1(1):1, July 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gabow:2009:FSI

[Gab09] Harold Gabow. Foreword to special issue on SODA 2007. ACM Transactions on Algorithms, 5(3):25:1–25:??, July 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gabow:2016:MPA

[Gab16] Harold N. Gabow. The minset-poset approach to representations of graph connectivity. ACM Transactions on Algorithms, 12(2): 24:1–24:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gabow:2017:DSN

[Gab17] Harold N. Gabow. A data structure for nearest common ancestors with linking. ACM Transactions on Algorithms, 13(4):45:1–45:??, December 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gabow:2025:MCM

[Gab25] Harold N. Gabow. Maximum cardinality f-matching in time O(n^{2/3}m). ACM Transactions on Algorithms, 21(1):9:1–9:??, January 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gawrychowski:2013:OPM

[Gaw13] Pawel Gawrychowski. Optimal pattern matching in LZW compressed strings. ACM Transactions on Algorithms, 9(3):25:1– 25:??, June 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gabow:2007:ISS

[GBFC07] H. N. Gabow, Michael A. Bender, and Martin Farach-Colton. Introduction to SODA 2002 and 2003 special issue. ACM Transactions on Algorithms, 3(4):36:1–36:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gonzalez-Gutierrez:2010:ACT

[GGG10] Arturo Gonzalez-Gutierrez and Teofilo F. Gonzalez. Approximating corridors and tours via restriction and relaxation techniques. ACM Transactions on Algorithms, 6(3):56:1–56:??, June 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Galvez:2021:AGK

[GGI^{+}21] Waldo Gálvez, Fabrizio Grandoni, Salvatore Ingala, Sandy Heydrich, Arindam Khan, and Andreas Wiese. Approximating geometric knapsack via L-packings. ACM Transactions on Algorithms, 17(4):33:1–33:67, October 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3473713.

Goel:2010:HPE

[GGM10] Ashish Goel, Sudipto Guha, and Kamesh Munagala. How to probe for an extreme value. ACM Transactions on Algorithms, 7 (1):12:1–12:??, November 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gramm:2006:PMA

[GGN06] Jens Gramm, Jiong Guo, and Rolf Niedermeier. Pattern matching for arc-annotated sequences. ACM Transactions on Algorithms, 2(1):44–65, January 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Graf:2022:AWI

[GHH22] Alessandra Graf, David G. Harris, and Penny Haxell. Algorithms for weighted independent transversals and strong colouring. ACM Transactions on Algorithms, 18(1):1:1–1:16, January 2022. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3474057.

Ganian:2024:FGC

[GHK^{+}24] Robert Ganian, Thekla Hamm, Viktoriia Korchemna, Karolina Okrasa, and Kirill Simonov. The fine-grained complexity of graph homomorphism parameterized by clique-width. ACM Transactions on Algorithms, 20(3):19:1–19:??, July 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/10.1145/3652514.

Gandhi:2006:IRD

[GHKS06] Rajiv Gandhi, Magnús M. Halldórsson, Guy Kortsarz, and Hadas Shachnai. Improved results for data migration and open shop scheduling. ACM Transactions on Algorithms, 2(1):116–129, January 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). See corrigendum [GHKS13].

Gandhi:2008:IBS

[GHKS08] Rajiv Gandhi, Magnús M. Halldórsson, Guy Kortsarz, and Hadas Shachnai. Improved bounds for scheduling conflicting jobs with minsum criteria. ACM Transactions on Algorithms, 4(1):11:1– 11:??, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gandhi:2013:CIR

[GHKS13] Rajiv Gandhi, Magnús M. Halldórsson, Guy Kortsarz, and Hadas Shachnai. Corrigendum: Improved results for data migration and open shop scheduling. ACM Transactions on Algorithms, 9(4):

34:1–34:??, September 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). See [GHKS06].

Graham:2016:AFN

[GHLL16] Ronald Graham, Linus Hamilton, Ariel Levavi, and Po-Shen Loh. Anarchy is free in network creation. ACM Transactions on Algorithms, 12(2):15:1–15:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gottesburen:2024:SHQ

[GHM^{+}24] Lars Gottesbüren, Tobias Heuer, Nikolai Maas, Peter Sanders, and Sebastian Schlag. Scalable high-quality hypergraph partitioning. ACM Transactions on Algorithms, 20(1):9:1–9:??, January 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3626527.

Gupta:2010:DRF

[GHNR10a] Anupam Gupta, Mohammadtaghi Hajiaghayi, Viswanath Nagarajan, and R. Ravi. Dial a Ride from k-forest. ACM Transactions on Algorithms, 6(2):41:1–41:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gupta:2010:DRK

[GHNR10b] Anupam Gupta, Mohammadtaghi Hajiaghayi, Viswanath Nagarajan, and R. Ravi. Dial a Ride from k-forest. ACM Transactions on Algorithms, 6(2):41:1–41:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Grabner:2005:ALC

[GHPT05] Peter J. Grabner, Clemens Heuberger, Helmut Prodinger, and Jörg M. Thuswaldner. Analysis of linear combination algorithms in cryptography. ACM Transactions on Algorithms, 1(1):123– 142, July 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Goranci:2018:IEM

[GHT18] Gramoz Goranci, Monika Henzinger, and Mikkel Thorup. Incremental exact min-cut in polylogarithmic amortized update time. ACM Transactions on Algorithms, 14(2):17:1–17:??, June 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gao:2019:CFO

[GIKW19] Jiawei Gao, Russell Impagliazzo, Antonina Kolokolova, and Ryan Williams. Completeness for first-order properties on sparse structures with algorithmic applications. ACM Transactions on Algorithms, 15(2):1–35, May 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3196275.

Georgiadis:2016:ECD

[GILP16] Loukas Georgiadis, Giuseppe F. Italiano, Luigi Laura, and Nikos Parotsidis. 2-edge connectivity in directed graphs. ACM Transactions on Algorithms, 13(1):9:1–9:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Grossi:2017:AOE

[GIN^{+}17] Roberto Grossi, John Iacono, Gonzalo Navarro, Rajeev Raman, and S. Rao Satti. Asymptotically optimal encodings of range data structures for selection and top-k queries. ACM Transactions on Algorithms, 13(2):28:1–28:??, May 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Giesen:2012:APC

[GJL12] Joachim Giesen, Martin Jaggi, and Sören Laue. Approximating parameterized convex optimization problems. ACM Transactions on Algorithms, 9(1):10:1–10:??, December 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Giannopoulou:2017:UKC

[GJLS17] Archontia C. Giannopoulou, Bart M. P. Jansen, Daniel Lokshtanov, and Saket Saurabh. Uniform kernelization complexity of hitting forbidden minors. ACM Transactions on Algorithms, 13 (3):35:1–35:??, August 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Groenland:2023:APG

[GJNW23] Carla Groenland, Gwenaël Joret, Wojciech Nadara, and Bartosz Walczak. Approximating pathwidth for graphs of small treewidth. ACM Transactions on Algorithms, 19(2):16:1–16:??, April 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3576044.

Giora:2009:ODV

[GK09] Yoav Giora and Haim Kaplan. Optimal dynamic vertical ray shooting in rectilinear planar subdivisions. ACM Transactions on Algorithms, 5(3):28:1–28:??, July 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Golubchik:2009:AAD

[GKK^{+}09] Leana Golubchik, Sanjeev Khanna, Samir Khuller, Ramakrishna Thurimella, and An Zhu. Approximation algorithms for data placement on parallel disks. ACM Transactions on Algorithms, 5(4):34:1–34:??, October 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Goel:2010:PMU

[GKK10] Ashish Goel, Michael Kapralov, and Sanjeev Khanna. Perfect matchings via uniform sampling in regular bipartite graphs. ACM Transactions on Algorithms, 6(2):27:1–27:??, March 2010. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Garg:2023:ANS

[GKK23] Jugal Garg, Pooja Kulkarni, and Rucha Kulkarni. Approximating Nash social welfare under submodular valuations through (un)Matchings. ACM Transactions on Algorithms, 19(4):36:1– 36:??, October 2023. CODEN ???? ISSN 1549-6325 (print), 1549-

6333 (electronic). URL https://dl.acm.org/doi/10.1145/
3613452.

Gaspers:2009:ETA

[GKLT09] Serge Gaspers, Dieter Kratsch, Mathieu Liedloff, and Ioan Todinca. Exponential time algorithms for the minimum dominating set problem on some graph classes. ACM Transactions on Algorithms, 6(1):9:1–9:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Golovnev:2016:FIS

[GKM16] Alexander Golovnev, Alexander S. Kulikov, and Ivan Mihajlin. Families with infants: Speeding up algorithms for NP-hard problems using FFT. ACM Transactions on Algorithms, 12(3):35:1– 35:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Grandoni:2008:DWV

[GKP08] Fabrizio Grandoni, Jochen Könemann, and Alessandro Panconesi. Distributed weighted vertex cover via maximal matchings. ACM Transactions on Algorithms, 5(1):6:1–6:12, November 2008. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Grzesik:2022:PTA

[GKPP22] Andrzej Grzesik, Tereza Klimosová, Marcin Pilipczuk, and Michał Pilipczuk. Polynomial-time algorithm for maximum weight independent set on P_{6}-free graphs. ACM Transactions on Algorithms, 18(1):4:1–4:57, January 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3414473.

Georgiadis:2011:DSM

[GKS^{+}11] Loukas Georgiadis, Haim Kaplan, Nira Shafrir, Robert E. Tarjan, and Renato F. Werneck. Data structures for mergeable trees. ACM Transactions on Algorithms, 7(2):14:1–14:??, March 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Guo:2021:ZHP

[GLLZ21] Heng Guo, Chao Liao, Pinyan Lu, and Chihao Zhang. Zeros of Holant problems: Locations and algorithms. ACM Transactions on Algorithms, 17(1):4:1–4:25, January 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3418056.

Gudmundsson:2008:ADO

[GLNS08] Joachim Gudmundsson, Christos Levcopoulos, Giri Narasimhan, and Michiel Smid. Approximate distance oracles for geometric spanners. ACM Transactions on Algorithms, 4(1):10:1–10:??, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gonzalez:2006:MTC

[GLP06] Teofilo F. Gonzalez, Joseph Y.-T. Leung, and Michael Pinedo. Minimizing total completion time on uniform machines with deadline constraints. ACM Transactions on Algorithms, 2(1):95–115, January 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Guerraoui:2008:CMA

[GLPP08] Rachid Guerraoui, Ron R. Levy, Bastian Pochon, and Jim Pugh. The collective memory of amnesic processes. ACM Transactions on Algorithms, 4(1):12:1–12:??, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gilbert:2017:ASR

[GLPS17] Anna C. Gilbert, Yi Li, Ely Porat, and Martin J. Strauss. Forall sparse recovery in near-optimal time. ACM Transactions on Algorithms, 13(3):32:1–32:??, August 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gao:2010:CLH

[GLS10] Jie Gao, Michael Langberg, and Leonard J. Schulman. Clustering lines in high-dimensional space: Classification of incomplete data. ACM Transactions on Algorithms, 7(1):8:1–8:??, November 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gishboliner:2023:CHC

[GLSY23] Lior Gishboliner, Yevgeny Levanzov, Asaf Shapira, and Raphael Yuster. Counting homomorphic cycles in degenerate graphs. ACM Transactions on Algorithms, 19(1):2:1–2:??, January 2023. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3560820.

Guha:2012:AUR

[GM12] Sudipto Guha and Kamesh Munagala. Adaptive uncertainty resolution in Bayesian combinatorial optimization problems. ACM Transactions on Algorithms, 8(1):1:1–1:??, January 2012. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Grohe:2014:CSF

[GM14] Martin Grohe and Dániel Marx. Constraint solving via fractional edge covers. ACM Transactions on Algorithms, 11(1):4:1–4:??, August 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Goel:2005:AMF

[GMP05] Ashish Goel, Adam Meyerson, and Serge Plotkin. Approximate majorization and fair online load balancing. ACM Transactions on Algorithms, 1(2):338–349, October 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Glacet:2017:TVI

[GMP17] Christian Glacet, Avery Miller, and Andrzej Pelc. Time vs. information tradeoffs for leader election in anonymous trees. ACM Transactions on Algorithms, 13(3):31:1–31:??, August 2017. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ganesh:2023:RAT

[GMP23a] Arun Ganesh, Bruce M. Maggs, and Debmalya Panigrahi. Robust algorithms for TSP and Steiner tree. ACM Transactions on Algorithms, 19(2):12:1–12:??, April 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3570957.

Ganesh:2023:UAC

[GMP23b] Arun Ganesh, Bruce M. Maggs, and Debmalya Panigrahi. Universal algorithms for clustering problems. ACM Transactions on Algorithms, 19(2):15:1–15:??, April 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3572840.

Golebiewski:2019:EOC

[GMS19] Zbigniew Gołebiewski, Abram Magner, and Wojciech Szpankowski. Entropy and optimal compression of some general plane trees. ACM Transactions on Algorithms, 15(1):1– 23, January 2019. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3275444.

Gairing:2011:RSF

[GMT11] Martin Gairing, Burkhard Monien, and Karsten Tiemann. Routing (un-) splittable flow in games with player-specific affine latency functions. ACM Transactions on Algorithms, 7(3):31:1– 31:??, July 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Guha:2009:SEE

[GMV09] Sudipto Guha, Andrew McGregor, and Suresh Venkatasubramanian. Sublinear estimation of entropy and information distances. ACM Transactions on Algorithms, 5(4):35:1–35:??, October 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gawrychowski:2020:SMQ

[GMW20] Paweł Gawrychowski, Shay Mozes, and Oren Weimann. Submatrix maximum queries in Monge and partial Monge matrices are equivalent to predecessor search. ACM Transactions on Algorithms, 16(2):16:1–16:24, April 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3381416.

Gabow:2008:FLD

[GN08] Harold N. Gabow and Shuxin Nie. Finding a long directed cycle. ACM Transactions on Algorithms, 4(1):7:1–7:??, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Goel:2014:PBP

[GN14] Ashish Goel and Hamid Nazerzadeh. Price-based protocols for fair resource allocation: Convergence time analysis and extension to Leontief utilities. ACM Transactions on Algorithms, 10(2): 5:1–5:??, February 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gortz:2015:MMM

[GNR15] Inge Li Gørtz, Viswanath Nagarajan, and R. Ravi. Minimum makespan multi-vehicle dial-a-ride. ACM Transactions on Algorithms, 11(3):23:1–23:??, January 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gupta:2016:RMO

[GNR16] Anupam Gupta, Viswanath Nagarajan, and R. Ravi. Robust and MaxMin optimization under matroid and knapsack uncertainty sets. ACM Transactions on Algorithms, 12(1):10:1–10:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Grohe:2020:IIT

[GNSW20] Martin Grohe, Daniel Neuen, Pascal Schweitzer, and Daniel Wiebking. An improved isomorphism test for bounded-tree-width graphs. ACM Transactions on Algorithms, 16(3):34:1–34:31, June 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3382082.

Goldwasser:2008:ONS

[GP08] Michael H. Goldwasser and Mark Pedigo. Online nonpreemptive scheduling of equal-length jobs on two identical machines. ACM Transactions on Algorithms, 5(1):2:1–2:18, November 2008. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gorain:2019:DGE

[GP19] Barun Gorain and Andrzej Pelc. Deterministic graph exploration with advice. ACM Transactions on Algorithms, 15(1): 1–17, January 2019. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3280823.

Gerke:2015:MML

[GPSS15] Stefanie Gerke, Konstantinos Panagiotou, Justus Schwartz, and Angelika Steger. Maximizing the minimum load for random processing times. ACM Transactions on Algorithms, 11(3):17:1– 17:??, January 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gordon:2010:CWT

[GR10] Eyal Gordon and Adi Rosén. Competitive weighted throughput analysis of greedy protocols on DAGs. ACM Transactions on Algorithms, 6(3):50:1–50:??, June 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Geary:2006:SOT

[GRR06] Richard F. Geary, Rajeev Raman, and Venkatesh Raman. Succinct ordinal trees with level-ancestor queries. ACM Transactions on Algorithms, 2(4):510–534, October 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ganian:2017:DAT

[GRS17] Robert Ganian, M. S. Ramanujan, and Stefan Szeider. Discovering archipelagos of tractability for constraint satisfaction and counting. ACM Transactions on Algorithms, 13(2):29:1–29:??, May 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Guruswami:2016:BUS

[GRSW16] Venkatesan Guruswami, Prasad Raghavendra, Rishi Saket, and Yi Wu. Bypassing UGC from some optimal geometric inapproximability results. ACM Transactions on Algorithms, 12(1):6:1– 6:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gamzu:2009:IOA

[GS09] Iftah Gamzu and Danny Segev. Improved online algorithms for the sorting buffer problem on line metrics. ACM Transactions on Algorithms, 6(1):15:1–15:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gaspers:2017:SMC

[GS17] Serge Gaspers and Gregory B. Sorkin. Separate, measure and conquer: Faster polynomial-space algorithms for Max 2-CSP and counting dominating sets. ACM Transactions on Algorithms, 13 (4):44:1–44:??, December 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gold:2018:DTW

[GS18] Omer Gold and Micha Sharir. Dynamic time warping and geometric edit distance: Breaking the quadratic barrier. ACM Transactions on Algorithms, 14(4):1–17, October 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3230734.

Golovach:2023:HTM

[GST23] Petr A. Golovach, Giannos Stamoulis, and Dimitrios M. Thilikos. Hitting topological minor models in planar graphs is fixed parameter tractable. ACM Transactions on Algorithms, 19(3):23:1–23:??, July 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3583688.

Galvez:2020:SEO

[GSV20] Waldo Gálvez, José A. Soto, and José Verschae. Symmetry exploitation for online machine covering with bounded migration. ACM Transactions on Algorithms, 16(4):43:1–43:22, September 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3397535.

Gudmundsson:2024:MMQ

[GSW24] Joachim Gudmundsson, Martin P. Seybold, and Sampson Wong. Map matching queries on realistic input graphs under the Fréchet distance. ACM Transactions on Algorithms, 20(2):14:1–14:??, April 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3643683.

Gu:2008:OBD

[GT08] Qian-Ping Gu and Hisao Tamaki. Optimal branch-decomposition of planar graphs in O(n^{3}) time. ACM Transactions on Algorithms, 4(3):30:1–30:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Georgiadis:2016:ADT

[GT16a] Loukas Georgiadis and Robert E. Tarjan. Addendum to “Dominator Tree Certification and Divergent Spanning Trees”. ACM Transactions on Algorithms, 12(4):56:1–56:??, September 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Georgiadis:2016:DTC

[GT16b] Loukas Georgiadis and Robert E. Tarjan. Dominator tree certification and divergent spanning trees. ACM Transactions on Algorithms, 12(1):11:1–11:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Gupta:2007:CPD

[GW07] Himanshu Gupta and Rephael Wenger. Constructing pairwise disjoint paths with few links. ACM Transactions on Algorithms, 3(3):26:1–26:??, August 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Grandoni:2020:FRP

[GW20] Fabrizio Grandoni and Virginia Vassilevska Williams. Faster replacement paths and distance sensitivity oracles. ACM Transactions on Algorithms, 16(1):15:1–15:25, January 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https:

//dl.acm.org/doi/abs/10.1145/3365835.

Gudmundsson:2022:IDM

[GW22] Joachim Gudmundsson and Sampson Wong. Improving the dilation of a metric graph by adding edges. ACM Transactions on Algorithms, 18(3):20:1–20:??, July 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3517807.

Gutin:2021:RSP

[GWZ21] Gregory Gutin, Magnus Wahlström, and Meirav Zehavi. r-simple k-path and related problems parameterized by k/r. ACM Transactions on Algorithms, 17(1):10:1–10:64, January 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https:

//dl.acm.org/doi/10.1145/3439721.

Halldorsson:2012:WSP

[Hal12] Magnús M. Halldórsson. Wireless scheduling with power control. ACM Transactions on Algorithms, 9(1):7:1–7:??, December 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Han:2007:OPS

[Han07] Yijie Han. Optimal parallel selection. ACM Transactions on Algorithms, 3(4):38:1–38:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Harris:2016:LMT

[Har16] David G. Harris. Lopsidependency in the Moser–Tardos framework: Beyond the Lopsided Lovász Local Lemma. ACM Transactions on Algorithms, 13(1):17:1–17:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Harris:2018:DPA

[Har18] David G. Harris. Deterministic parallel algorithms for fooling polylogarithmic juntas and the Lovász local lemma. ACM Transactions on Algorithms, 14(4):1–24, October 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3230651.

Harris:2019:DCB

[Har19] David G. Harris. Derandomized concentration bounds for polynomials, and hypergraph maximal independent set. ACM Transactions on Algorithms, 15(3):1–29, July 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3326171.

Harris:2021:ORO

[Har21] David G. Harris. Oblivious resampling oracles and parallel algorithms for the lopsided Lovász local lemma. ACM Transactions on Algorithms, 17(1):1:1–1:32, January 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3392035.

Han:2011:NUB

[HCT^{+}11] Xin Han, Francis Y. L. Chin, Hing-Fung Ting, Guochuan Zhang, and Yong Zhang. A new upper bound 2.5545 on 2D Online Bin Packing. ACM Transactions on Algorithms, 7(4):50:1–50:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hegde:2006:FSE

[Heg06] Rajneesh Hegde. Finding 3-shredders efficiently. ACM Transactions on Algorithms, 2(1):14–43, January 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Haeupler:2017:PAC

[HH17] Bernhard Haeupler and David G. Harris. Parallel algorithms and concentration bounds for the Lovász local lemma via witness DAGs. ACM Transactions on Algorithms, 13(4):53:1–53:??, December 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hajiaghayi:2016:CFA

[HHL^{+}16] Mohammadtaghi Hajiaghayi, Wei Hu, Jian Li, Shi Li, and Barna Saha. A constant factor approximation algorithm for faulttolerant k-median. ACM Transactions on Algorithms, 12(3):36:1– 36:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hujdurovic:2018:PPB

[HHM^{+}18] Ademir Hujdurović, Edin Husić, Martin Milanić, Romeo Rizzi, and Alexandru I. Tomescu. Perfect phylogenies via branchings in acyclic digraphs and a generalization of Dilworth’s theorem. ACM Transactions on Algorithms, 14(2):20:1–20:??, June 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Halldorsson:2007:IAR

[HIMY07] Magnús M. Halldórsson, Kazuo Iwama, Shuichi Miyazaki, and Hiroki Yanagisawa. Improved approximation results for the stable marriage problem. ACM Transactions on Algorithms, 3(3):30:1– 30:??, August 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hirai:2019:TCB

[HIMŽ19] Hiroshi Hirai, Yuni Iwamasa, Kazuo Murota, and Stanislav Živný. A tractable class of binary VCSPs via M-convex intersection. ACM Transactions on Algorithms, 15(3):1–41, July 2019. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3329862.

Hirai:2019:DDA

[Hir19] Hiroshi Hirai. A dual descent algorithm for node-capacitated multiflow problems and its applications. ACM Transactions on Algorithms, 15(1):1–24, January 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3291531.

Hohn:2015:PSR

[HJ15] Wiebke Höhn and Tobias Jacobs. On the performance of Smith’s rule in single-machine scheduling with nonlinear cost. ACM Transactions on Algorithms, 11(4):25:1–25:??, June 2015. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hwang:2017:EAS

[HJT17] Hsien-Kuei Hwang, Svante Janson, and Tsung-Hsi Tsai. Exact and asymptotic solutions of a divide-and-conquer recurrence dividing at half: Theory and applications. ACM Transactions on Algorithms, 13(4):47:1–47:??, December 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hoefer:2022:IAS

[HK22] Martin Hoefer and Tsvi Kopelowitz. Introduction to the ACM-SIAM Symposium on Discrete Algorithms (SODA) 2019 special issue. ACM Transactions on Algorithms, 18(1):4e:1–4e:2, January 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3508460.

Harris:2025:PEG

[HK25] David G. Harris and Vladimir Kolmogorov. Parameter estimation for Gibbs distributions. ACM Transactions on Algorithms, 21(1): 3:1–3:??, January 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hajiaghayi:2016:AAM

[HKKK16] Mohammad Taghi Hajiaghayi, Rohit Khandekar, Mohammad Reza Khani, and Guy Kortsarz. Approximation algorithms for movement repairmen. ACM Transactions on Algorithms, 12 (4):54:1–54:??, September 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hajiaghayi:2012:PCS

[HKKN12] Mohammadtaghi Hajiaghayi, Rohit Khandekar, Guy Kortsarz, and Zeev Nutov. Prize-collecting Steiner network problems. ACM Transactions on Algorithms, 9(1):2:1–2:??, December 2012. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Haeupler:2012:ICD

[HKM^{+}12] Bernhard Haeupler, Telikepalli Kavitha, Rogers Mathew, Siddhartha Sen, and Robert E. Tarjan. Incremental Cycle Detection, Topological Ordering, and Strong Component Maintenance. ACM Transactions on Algorithms, 8(1):3:1–3:??, January 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Henzinger:2017:STM

[HKN17] Monika Henzinger, Sebastian Krinninger, and Danupon Nanongkai. Sublinear-time maintenance of breadth-first spanning trees in partially dynamic networks. ACM Transactions on Algorithms, 13 (4):51:1–51:??, December 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hopkins:2018:IGD

[HKP^{+}18] Samuel B. Hopkins, Pravesh Kothari, Aaron Henry Potechin, Prasad Raghavendra, and Tselil Schramm. On the integrality gap of degree-4 sum of squares for planted clique. ACM Transactions on Algorithms, 14(3):28:1–28:??, July 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/ft_gateway.cfm?id=3178538.

Hajiaghayi:2007:ORN

[HKRL07] Mohammad Taghi Hajiaghayi, Robert D. Kleinberg, Harald Räcke, and Tom Leighton. Oblivious routing on node-capacitated and directed graphs. ACM Transactions on Algorithms, 3(4): 51:1–51:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Halldorsson:2011:SEC

[HKS11] Magnús M. Halldórsson, Guy Kortsarz, and Maxim Sviridenko. Sum edge coloring of multigraphs via configuration LP. ACM Transactions on Algorithms, 7(2):22:1–22:21, March 2011. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hansen:2017:HH

[HKTZ17] Thomas Dueholm Hansen, Haim Kaplan, Robert E. Tarjan, and Uri Zwick. Hollow heaps. ACM Transactions on Algorithms, 13(3):42:1–42:??, August 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hassin:2006:MGV

[HL06a] Refael Hassin and Asaf Levin. The minimum generalized vertex cover problem. ACM Transactions on Algorithms, 2(1):66–78, January 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Huo:2006:MMF

[HL06b] Yumei Huo and Joseph Y.-T. Leung. Minimizing mean flow time for UET tasks. ACM Transactions on Algorithms, 2(2):244–262, April 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hochbaum:2013:AAM

[HL13] Dorit S. Hochbaum and Asaf Levin. Approximation algorithms for a minimization variant of the order-preserving submatrices and for biclustering problems. ACM Transactions on Algorithms, 9(2):19:1–19:??, March 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hazay:2007:APM

[HLS07] Carmit Hazay, Moshe Lewenstein, and Dina Sokol. Approximate parameterized matching. ACM Transactions on Algorithms, 3(3): 29:1–29:??, August 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hassin:2009:AMQ

[HLS09] Refael Hassin, Asaf Levin, and Maxim Sviridenko. Approximating the minimum quadratic assignment problems. ACM Transactions on Algorithms, 6(1):18:1–18:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hershberger:2007:FSS

[HMS07] John Hershberger, Matthew Maxel, and Subhash Suri. Finding the k shortest simple paths: a new algorithm and its implementation. ACM Transactions on Algorithms, 3(4):45:1–45:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

He:2012:SOT

[HMS12] Meng He, J. Ian Munro, and Srinivasa Rao Satti. Succinct ordinal trees based on tree covering. ACM Transactions on Algorithms, 8(4):42:1–42:??, September 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hermelin:2019:DWS

[HMVW19] Danny Hermelin, Matthias Mnich, Erik Jan Van Leeuwen, and Gerhard Woeginger. Domination when the stars are out. ACM Transactions on Algorithms, 15(2):1–90, May 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3301445.

He:2016:DSP

[HMZ16] Meng He, J. Ian Munro, and Gelin Zhou. Data structures for path queries. ACM Transactions on Algorithms, 12(4):53:1–53:??, September 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Henzinger:2022:CTD

[HP22] Monika Henzinger and Pan Peng. Constant-time dynamic (Δ + 1)-coloring. ACM Transactions on Algorithms, 18(2):16:1–16:21, April 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3501403.

Har-Peled:2021:JCP

[HPJ21] Sariel Har-Peled and Mitchell Jones. Journey to the center of the point set. ACM Transactions on Algorithms, 17(1):9:1–9:21, January 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3431285.

Har-Peled:2023:RSM

[HPMO23] Sariel Har-Peled, Manor Mendel, and Dániel Oláh. Reliable spanners for metric spaces. ACM Transactions on Algorithms, 19(1):7:1–7:??, January 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3563356.

Har-Peled:2014:FDR

[HPR14] Sariel Har-Peled and Benjamin Raichel. The Fréchet distance revisited and extended. ACM Transactions on Algorithms, 10(1): 3:1–3:??, January 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Harris:2019:LMC

[HPST19] David G. Harris, Thomas Pensyl, Aravind Srinivasan, and Khoa Trinh. A lottery model for center-type problems with outliers. ACM Transactions on Algorithms, 15(3):1–25, July 2019. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3311953.

Hassin:2006:RST

[HS06] Refael Hassin and Danny Segev. Robust subgraphs for trees and paths. ACM Transactions on Algorithms, 2(2):263–281, April 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hay:2009:JRM

[HS09] David Hay and Gabriel Scalosub. Jitter regulation for multiple streams. ACM Transactions on Algorithms, 6(1):12:1–12:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Harris:2017:AEA

[HS17] David G. Harris and Aravind Srinivasan. Algorithmic and enumerative aspects of the Moser–Tardos distribution. ACM Transactions on Algorithms, 13(3):33:1–33:??, August 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Held:2018:BAC

[HS18] Stephan Held and Sophie Theresa Spirkl. Binary adder circuits of asymptotically minimum depth, linear size, and fan-out two. ACM Transactions on Algorithms, 14(1):4:1–4:??, January 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hershberger:2007:DSS

[HSB07] John Hershberger, Subhash Suri, and Amit Bhosle. On the difficulty of some shortest path problems. ACM Transactions on Algorithms, 3(1):??, February 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hayward:2007:IAW

[HSS07] Ryan B. Hayward, Jeremy P. Spinrad, and R. Sritharan. Improved algorithms for weakly chordal graphs. ACM Transactions on Algorithms, 3(2):14:1–14:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Haslegrave:2022:TDB

[HSS22] John Haslegrave, Thomas Sauerwald, and John Sylvester. Time dependent biased random walks. ACM Transactions on Algorithms, 18(2):12:1–12:30, April 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/

doi/10.1145/3498848.

Haeupler:2015:RBT

[HST15] Bernhard Haeupler, Siddhartha Sen, and Robert E. Tarjan. Rankbalanced trees. ACM Transactions on Algorithms, 11(4):30:1– 30:??, June 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hajiaghayi:2010:FSI

[HT10] Mohammad T. Hajiaghayi and Shang-Hua Teng. Foreword to special issue on SODA 2008. ACM Transactions on Algorithms, 6(4):58:1–58:??, August 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Halldorsson:2021:SBO

[HT21] Magnús M. Halldórsson and Tigran Tonoyan. Sparse backbone and optimal distributed SINR algorithms. ACM Transactions on Algorithms, 17(2):17:1–17:34, June 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3452937.

Huang:2019:OVW

[HTWZ19] Zhiyi Huang, Zhihao Gavin Tang, Xiaowei Wu, and Yuhao Zhang. Online vertex-weighted bipartite matching: Beating 1 − 1/e with random arrivals. ACM Transactions on Algorithms, 15(3):1–15, July 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3326169.

Hsu:2020:NAF

[HU20] Chloe Ching-Yun Hsu and Chris Umans. A new algorithm for fast generalized DFTs. ACM Transactions on Algorithms, 16(1): 4:1–4:20, January 2020. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3301313.

Hunkenschroder:2019:AAM

[HVV19] Christoph Hunkenschröder, Santosh Vempala, and Adrian Vetta. A 4/3-approximation algorithm for the minimum 2-edge connected subgraph problem. ACM Transactions on Algorithms, 15 (4):1–28, October 2019. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3341599.

Heydrich:2019:FAS

[HW19] Sandy Heydrich and Andreas Wiese. Faster approximation schemes for the two-dimensional knapsack problem. ACM Transactions on Algorithms, 15(4):1–28, October 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3338512.

Ibarra:2008:FDA

[Iba08] Louis Ibarra. Fully dynamic algorithms for chordal graphs and split graphs. ACM Transactions on Algorithms, 4(4):40:1–40:??, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ito:2009:EIC

[II09] Hiro Ito and Kazuo Iwama. Enumeration of isolated cliques and pseudo-cliques. ACM Transactions on Algorithms, 5(4):40:1– 40:??, October 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ito:2023:MEF

[IIK^{+}23] Takehiro Ito, Yuni Iwamasa, Naonori Kakimura, Naoyuki Kamiyama, Yusuke Kobayashi, Shun-Ichi Maezawa, Yuta Nozaki, Yoshio Okamoto, and Kenta Ozeki. Monotone edge flips to an orientation of maximum edge-connectivity à la Nash–Williams. ACM Transactions on Algorithms, 19(1):6:1–6:??, January 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3561302.

Ito:2025:RPC

[IIK^{+}25] Takehiro Ito, Yuni Iwamasa, Naonori Kakimura, Yusuke Kobayashi, Shun-Ichi Maezawa, Yuta Nozaki, Yoshio Okamoto, and Kenta Ozeki. Rerouting planar curves and disjoint paths. ACM Transactions on Algorithms, 21(2):20:1–20:??, April 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Irving:2006:RMM

[IKM^{+}06] Robert W. Irving, Telikepalli Kavitha, Kurt Mehlhorn, Dimitrios Michail, and Katarzyna E. Paluch. Rank-maximal matchings.

ACM Transactions on Algorithms, 2(4):602–610, October 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kawarabayashi:2025:AIM

[iKMNZ25] Ken ichi Kawarabayashi, Bojan Mohar, Roman Nedela, and Peter Zeman. Automorphisms and isomorphisms of maps in linear time. ACM Transactions on Algorithms, 21(1):6:1–6:??, January 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Im:2012:OSA

[IM12] Sungjin Im and Benjamin Moseley. An online scalable algorithm for average flow time in broadcast scheduling. ACM Transactions on Algorithms, 8(4):39:1–39:??, September 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Immorlica:2008:LCM

[IMM08] Nicole Immorlica, Mohammad Mahdian, and Vahab S. Mirrokni. Limitations of cross-monotonic cost-sharing schemes. ACM Transactions on Algorithms, 4(2):24:1–24:??, May 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Iwama:2010:AAS

[IMY10] Kazuo Iwama, Shuichi Miyazaki, and Hiroki Yanagisawa. Approximation algorithms for the sex-equal stable marriage problem. ACM Transactions on Algorithms, 7(1):2:1–2:??, November 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Indyk:2007:NNP

[IN07] Piotr Indyk and Assaf Naor. Nearest-neighbor-preserving embeddings. ACM Transactions on Algorithms, 3(3):31:1–31:??, August 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Im:2016:MLS

[INV16] Sungjin Im, Viswanath Nagarajan, and Ruben Van Der Zwaan. Minimum latency submodular cover. ACM Transactions on Algorithms, 13(1):13:1–13:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Izsak:2011:CPM

[IP11] Alexander Izsak and Nicholas Pippenger. Carry propagation in multiplication by constants. ACM Transactions on Algorithms, 7 (4):54:1–54:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Irani:2007:APS

[ISG07] Sandy Irani, Sandeep Shukla, and Rajesh Gupta. Algorithms for power savings. ACM Transactions on Algorithms, 3(4):41:1– 41:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Jacobs:2011:CFA

[Jac11] Tobias Jacobs. Constant factor approximations for the hotlink assignment problem. ACM Transactions on Algorithms, 7(2): 16:1–16:??, March 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Janson:2005:IDL

[Jan05] Svante Janson. Individual displacements for linear probing hashing with different insertion policies. ACM Transactions on Algorithms, 1(2):177–213, October 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Jin:2022:GSI

[JCCW22] Kai Jin, Siu-Wing Cheng, Man-Kwun Chiu, and Man Ting Wong. A generalization of self-improving algorithms. ACM Transactions on Algorithms, 18(3):29:1–29:??, July 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3531227.

Jez:2015:FFC

[Jez15] Artur Jez. Faster fully compressed pattern matching by recompression. ACM Transactions on Algorithms, 11(3):20:1–20:??, January 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Jez:2012:OSP

[JLSS12] Lukasz Jez, Fei Li, Jay Sethuraman, and Clifford Stein. Online scheduling of packets with agreeable deadlines. ACM Transactions on Algorithms, 9(1):5:1–5:??, December 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Jansen:2019:ASM

[JMR19] Klaus Jansen, Marten Maack, and Malin Rau. Approximation schemes for machine scheduling with resource (in-)dependent processing times. ACM Transactions on Algorithms, 15(3):1–28, July 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3302250.

Joseph:2022:ESL

[JMR22] Matthew Joseph, Jieming Mao, and Aaron Roth. Exponential separations in local privacy. ACM Transactions on Algorithms, 18(4):32:1–32:??, October 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3459095.

Jin:2024:QSU

[JN24] Ce Jin and Jakob Nogler. Quantum speed-ups for string synchronizing sets, longest common substring, and k-mismatch matching. ACM Transactions on Algorithms, 20(4):32:1–32:??, October 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3672395.

Johnson:2005:NCC

[Joh05] David S. Johnson. The NP-completeness column. ACM Transactions on Algorithms, 1(1):160–176, July 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Johnson:2006:NCC

[Joh06] David S. Johnson. The NP-completeness column: The many limits on approximation. ACM Transactions on Algorithms, 2(3): 473–489, July 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Johnson:2007:NCC

[Joh07] David S. Johnson. The NP-completeness column: Finding needles in haystacks. ACM Transactions on Algorithms, 3(2):24:1–24:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Jiang:2022:OAW

[JPS22] Zhihao Jiang, Debmalya Panigrahi, and Kevin Sun. Online algorithms for weighted paging with predictions. ACM Transactions on Algorithms, 18(4):39:1–39:??, October 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/10.1145/3548774.

Jothi:2005:AAC

[JR05] Raja Jothi and Balaji Raghavachari. Approximation algorithms for the capacitated minimum spanning tree problem and its variants in network design. ACM Transactions on Algorithms, 1(2): 265–282, October 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Janson:2007:PFS

[JS07] Svante Janson and Wojciech Szpankowski. Partial fillup and search time in LC tries. ACM Transactions on Algorithms, 3 (4):44:1–44:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Jayaprakash:2023:ASC

[JS23] Aditya Jayaprakash and Mohammad R. Salavatipour. Approximation schemes for capacitated vehicle routing on graphs of bounded treewidth, bounded doubling, or highway dimension. ACM Transactions on Algorithms, 19(2):20:1–20:??, April 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3582500.

Jambulapati:2025:UUF

[JS25] Arun Jambulapati and Aaron Sidford. Ultrasparse ultrasparsifiers and faster Laplacian system solvers. ACM Transactions on Algorithms, 21(3):31:1–31:49, July 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Juge:2024:ASS

[Jug24] Vincent Jugé. Adaptive shivers sort: an alternative sorting algorithm. ACM Transactions on Algorithms, 20(4):31:1–31:55, October 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3664195.

Jayram:2013:OBJ

[JW13] T. S. Jayram and David P. Woodruff. Optimal bounds for Johnson–Lindenstrauss Transforms and streaming problems with subconstant error. ACM Transactions on Algorithms, 9(3):26:1– 26:??, June 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Jayaram:2023:TOM

[JW23] Rajesh Jayaram and David P. Woodruff. Towards optimal moment estimation in streaming and distributed models. ACM Transactions on Algorithms, 19(3):27:1–27:??, July 2023. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3596494.

Jansen:2006:AAS

[JZ06] Klaus Jansen and Hu Zhang. An approximation algorithm for scheduling malleable tasks under general precedence constraints. ACM Transactions on Algorithms, 2(3):416–434, July 2006. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kamath:2022:ISI

[KADK22] Gautam Kamath, Sepehr Assadi, Anne Driemel, and Janardhan Kulkarni. Introduction to the special issue on ACM-SIAM Symposium on Discrete Algorithms (SODA) 2020. ACM Transactions on Algorithms, 18(4):30:1–30:??, October 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/10.1145/3561912.

Karakostas:2008:FAS

[Kar08] George Karakostas. Faster approximation schemes for fractional multicommodity flow problems. ACM Transactions on Algorithms, 4(1):13:1–13:17, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Karakostas:2009:BAR

[Kar09] George Karakostas. A better approximation ratio for the vertex cover problem. ACM Transactions on Algorithms, 5(4):41:1– 41:??, October 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kauers:2007:ADZ

[Kau07] Manuel Kauers. An algorithm for deciding zero equivalence of nested polynomially recurrent sequences. ACM Transactions on Algorithms, 3(2):18:1–18:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kavitha:2024:PMO

[Kav24] Telikepalli Kavitha. Popular matchings with one-sided bias. ACM Transactions on Algorithms, 20(1):8:1–8:??, January 2024. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3638764.

Katriel:2006:OTO

[KB06] Irit Katriel and Hans L. Bodlaender. Online topological ordering. ACM Transactions on Algorithms, 2(3):364–379, July 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kisfaludi-Bak:2020:NET

[KBNvL20] Sándor Kisfaludi-Bak, Jesper Nederlof, and Erik Jan van Leeuwen. Nearly ETH-tight algorithms for planar Steiner tree with terminals on few faces. ACM Transactions on Algorithms, 16 (3):28:1–28:30, June 2020. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3371389.

Khuller:2005:PC

[Khu05] Samir Khuller. Problems column. ACM Transactions on Algorithms, 1(1):157–159, July 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Khuller:2006:PC

[Khu06] Samir Khuller. Problems column. ACM Transactions on Algorithms, 2(1):130–134, January 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Khuller:2007:PC

[Khu07] Samir Khuller. Problems column. ACM Transactions on Algorithms, 3(3):35:1–35:??, August 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kowalik:2006:OBL

[KK06] Lukasz Kowalik and Maciej Kurowski. Oracles for bounded-length shortest paths in planar graphs. ACM Transactions on Algorithms, 2(3):335–363, July 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kawarabayashi:2013:LAA

[KK13] Ken-Ichi Kawarabayashi and Yusuke Kobayashi. An O(log n)- approximation algorithm for the edge-disjoint paths problem in Eulerian planar graphs. ACM Transactions on Algorithms, 9(2): 16:1–16:??, March 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kawarabayashi:2016:IAA

[KK16] Ken-Ichi Kawarabayashi and Yusuke Kobayashi. An improved approximation algorithm for the edge-disjoint paths problem with congestion two. ACM Transactions on Algorithms, 13(1):5:1–5:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kapron:2010:FAB

[KKK^{+}10] Bruce M. Kapron, David Kempe, Valerie King, Jared Saia, and Vishal Sanwalani. Fast asynchronous Byzantine agreement and leader election with full information. ACM Transactions on Algorithms, 6(4):68:1–68:??, August 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Karppa:2018:FSA

[KKK18] Matti Karppa, Petteri Kaski, and Jukka Kohonen. A faster subquadratic algorithm for finding outlier correlations. ACM Transactions on Algorithms, 14(3):31:1–31:??, July 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/ft_gateway.cfm?id=3174804.

Kranakis:2011:RRL

[KKM11] Evangelos Kranakis, Danny Krizanc, and Pat Morin. Randomized rendezvous with limited memory. ACM Transactions on Algorithms, 7(3):34:1–34:??, July 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kim:2024:FAI

[KKPW24] Eun Jung Kim, Stefan Kratsch, Marcin Pilipczuk, and Magnus Wahlström. Flow-augmentation II: Undirected graphs. ACM Transactions on Algorithms, 20(2):12:1–12:??, April 2024. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3641105.

Kazda:2019:EDM

[KKR19] Alexandr Kazda, Vladimir Kolmogorov, and Michal Rolínek. Even delta-matroids and the complexity of planar Boolean CSPs. ACM Transactions on Algorithms, 15(2):1–33, May 2019. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3230649.

Kociumaka:2020:LTA

[KKR^{+}20] Tomasz Kociumaka, Marcin Kubica, Jakub Radoszewski, Wojciech Rytter, and Tomasz Waleń. A linear-time algorithm for seeds computation. ACM Transactions on Algorithms, 16(2): 27:1–27:23, April 2020. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3386369.

Karakostas:2012:FMT

[KKW12] George Karakostas, Stavros G. Kolliopoulos, and Jing Wang. An FPTAS for the minimum total weighted tardiness problem with a fixed number of distinct due dates. ACM Transactions on Algorithms, 8(4):40:1–40:??, September 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Korsh:2006:LGC

[KL06] James Korsh and Paul Lafollette. A loopless Gray code for rooted trees. ACM Transactions on Algorithms, 2(2):135–152, April 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Khan:2025:TAA

[KLM^{+}25] Arindam Khan, Aditya Lonkar, Arnab Maiti, Amatya Sharma, and Andreas Wiese. Tight approximation algorithms for 2D guillotine strip packing. ACM Transactions on Algorithms, 21(3): 28:1–28:30, July 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kim:2016:LKS

[KLP^{+}16a] Eun Jung Kim, Alexander Langer, Christophe Paul, Felix Reidl, Peter Rossmanith, Ignasi Sau, and Somnath Sikdar. Linear kernels and single-exponential algorithms via protrusion decompositions. ACM Transactions on Algorithms, 12(2):21:1–21:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Koutis:2016:FSS

[KLP16b] Ioannis Koutis, Alex Levin, and Richard Peng. Faster spectral sparsification and numerical algorithms for SDD matrices. ACM Transactions on Algorithms, 12(2):17:1–17:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kantor:2023:MPS

[KLPP23] Erez Kantor, Zvi Lotker, Merav Parter, and David Peleg. The minimum principle of SINR: a useful discretization tool for wireless communication. ACM Transactions on Algorithms, 19 (1):1:1–1:??, January 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3477144.

Krokhin:2012:HLW

[KM12] Andrei Krokhin and Dániel Marx. On the hardness of losing weight. ACM Transactions on Algorithms, 8(2):19:1–19:??, April 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Khuller:2011:FFG

[KMM11] Samir Khuller, Azarakhsh Malekian, and Julián Mestre. To fill or not to fill: The gas station problem. ACM Transactions on Algorithms, 7(3):36:1–36:??, July 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kavitha:2007:SSM

[KMMP07] Telikepalli Kavitha, Kurt Mehlhorn, Dimitrios Michail, and Katarzyna E. Paluch. Strongly stable matchings in time O(nm) and extension to the hospitals-residents problem. ACM Transactions on Algorithms, 3(2):15:1–15:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kaplan:2017:SMQ

[KMNS17] Haim Kaplan, Shay Mozes, Yahav Nussbaum, and Micha Sharir. Submatrix maximum queries in Monge matrices and partial Monge matrices, and their applications. ACM Transactions on Algorithms, 13(2):26:1–26:??, May 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kumar:2016:DAE

[KMPS16] V. S. Anil Kumar, Madhav V. Marathe, Srinivasan Parthasarathy, and Aravind Srinivasan. Distributed algorithms for end-to-end packet scheduling in wireless ad hoc networks. ACM Transactions on Algorithms, 12(3):28:1–28:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kavitha:2025:PAT

[KMSY25] Telikepalli Kavitha, Kazuhisa Makino, Ildikó Schlotter, and Yu Yokoi. Popular arborescences and their matroid generalization. ACM Transactions on Algorithms, 21(2):22:1–22:??, April 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kanade:2023:CTG

[KMTS23] Varun Kanade, Frederik Mallmann-Trenn, and Thomas Sauerwald. On coalescence time in graphs: When is coalescing as fast as meeting? ACM Transactions on Algorithms, 19(2):18:1–18:??, April 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3576900.

Klein:2010:SPD

[KMW10] Philip N. Klein, Shay Mozes, and Oren Weimann. Shortest paths in directed planar graphs with negative lengths: a linear-space O(n log^{2} n)-time algorithm. ACM Transactions on Algorithms, 6(2):30:1–30:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kannan:2018:GRV

[KMZ18] Sampath Kannan, Claire Mathieu, and Hang Zhou. Graph reconstruction and verification. ACM Transactions on Algorithms, 14 (4):1–30, October 2018. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3199606.

Kortsarz:2016:SAA

[KN16] Guy Kortsarz and Zeev Nutov. A simplified 1.5-approximation algorithm for augmenting edge-connectivity of a graph from 1 to 2. ACM Transactions on Algorithms, 12(2):23:1–23:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Krivelevich:2007:AAH

[KNS^{+}07] Michael Krivelevich, Zeev Nutov, Mohammad R. Salavatipour, Jacques Verstraete Yuster, and Raphael Yuster. Approximation algorithms and hardness results for cycle packing problems. ACM Transactions on Algorithms, 3(4):48:1–48:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kolluri:2008:PGM

[Kol08] Ravikrishna Kolluri. Provably good moving least squares. ACM Transactions on Algorithms, 4(2):18:1–18:??, May 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kolmogorov:2025:SPA

[Kol25] Vladimir Kolmogorov. A simpler and parallelizable O(\sqrt{log n})- approximation algorithm for sparsest cut. ACM Transactions on Algorithms, 21(4):46:1–46:22, October 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Korman:2010:LSV

[Kor10] Amos Korman. Labeling schemes for vertex connectivity. ACM Transactions on Algorithms, 6(2):39:1–39:??, March 2010. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Korman:2008:DRS

[KP08] Amos Korman and David Peleg. Dynamic routing schemes for graphs with low local density. ACM Transactions on Algorithms, 4(4):41:1–41:??, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kratsch:2016:PLC

[KPR16] Stefan Kratsch, Geevarghese Philip, and Saurabh Ray. Point line cover: The easy kernel is essentially tight. ACM Transactions on Algorithms, 12(3):40:1–40:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Konrad:2016:ASM

[KR16] Christian Konrad and Adi Rosén. Approximating semi-matchings in streaming and in two-party communication. ACM Transactions on Algorithms, 12(3):32:1–32:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Korman:2018:DDP

[KR18] Amos Korman and Yoav Rodeh. The dependent doors problem: an investigation into sequential decisions without feedback. ACM Transactions on Algorithms, 14(4):1–23, October 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https:

//dl.acm.org/doi/abs/10.1145/3218819.

Klemz:2019:OLP

[KR19] Boris Klemz and Günter Rote. Ordered level planarity and its relationship to geodesic planarity, bi-monotonicity, and variations of level planarity. ACM Transactions on Algorithms, 15 (4):1–25, October 2019. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3359587.

Klein:2024:CTC

[KR24] Kim-Manuel Klein and Janina Reuter. Collapsing the tower-on the complexity of multistage stochastic IPs. ACM Transactions on Algorithms, 20(3):22:1–22:??, July 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3604554.

Kratsch:2014:CNC

[Kra14] Stefan Kratsch. Co-nondeterminism in compositions: a kernelization lower bound for a Ramsey-type problem. ACM Transactions on Algorithms, 10(4):19:1–19:??, August 2014. ISSN 1549-6325 (print), 1549-6333 (electronic).

Kreutzer:2019:PKW

[KRS19] Stephan Kreutzer, Roman Rabinovich, and Sebastian Siebertz. Polynomial kernels and wideness properties of nowhere dense graph classes. ACM Transactions on Algorithms, 15(2):1–19, May 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3274652.

Konjevod:2016:SFC

[KRX16] Goran Konjevod, Andréa W. Richa, and Donglin Xia. Scale-free compact routing schemes in networks of low doubling dimension. ACM Transactions on Algorithms, 12(3):27:1–27:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kowalski:2008:WAD

[KS08] Dariusz R. Kowalski and Alexander A. Shvartsman. Writingall deterministically and optimally using a nontrivial number of asynchronous processors. ACM Transactions on Algorithms, 4 (3):33:1–33:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Krishnaswamy:2016:IMU

[KS16] Ravishankar Krishnaswamy and Maxim Sviridenko. Inapproximability of the multilevel uncapacitated facility location problem. ACM Transactions on Algorithms, 13(1):1:1–1:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kao:2009:RFD

[KSS09] Ming-Yang Kao, Manan Sanghi, and Robert Schweller. Randomized fast design of short DNA words. ACM Transactions on Algorithms, 5(4):43:1–43:??, October 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kaplan:2008:THT

[KT08] Haim Kaplan and Robert Endre Tarjan. Thin heaps, thick heaps. ACM Transactions on Algorithms, 4(1):3:1–3:??, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Krauthgamer:2018:CLB

[KT18] Robert Krauthgamer and Ohad Trabelsi. Conditional lower bounds for all-pairs max-flow. ACM Transactions on Algorithms, 14(4):1–15, October 2018. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3212510.

Kannan:2019:LEF

[KT19] Sampath Kannan and Kevin T. Tian. Locating errors in faulty formulas. ACM Transactions on Algorithms, 15(3):1–13, July 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3313776.

Kratsch:2014:CMR

[KW14] Stefan Kratsch and Magnus Wahlström. Compression via matroids: a randomized polynomial kernel for odd cycle transversal. ACM Transactions on Algorithms, 10(4):20:1–20:??, August 2014. ISSN 1549-6325 (print), 1549-6333 (electronic).

Koutis:2016:LAG

[KW16a] Ioannis Koutis and Ryan Williams. Limits and applications of group algebras for parameterized problems. ACM Transactions on Algorithms, 12(3):31:1–31:??, June 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Krauthgamer:2016:CTA

[KW16b] Robert Krauthgamer and Tal Wagner. Cheeger-type approximation for sparsest st-cut. ACM Transactions on Algorithms, 13 (1):14:1–14:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Koiliaris:2019:FPT

[KX19] Konstantinos Koiliaris and Chao Xu. Faster pseudopolynomial time algorithms for subset sum. ACM Transactions on Algorithms, 15(3):1–20, July 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3329863.

Kaltofen:2013:MBM

[KY13] Erich Kaltofen and George Yuhasz. On the matrix Berlekamp– Massey algorithm. ACM Transactions on Algorithms, 9(4):33:1– 33:??, September 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Lacki:2013:IDA

[Lac13] Jakub Lacki. Improved deterministic algorithms for decremental reachability and strongly connected components. ACM Transactions on Algorithms, 9(3):27:1–27:??, June 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Lafond:2023:RLP

[Laf23] Manuel Lafond. Recognizing k-leaf powers in polynomial time, for constant k. ACM Transactions on Algorithms, 19(4):35:1–35:??, October 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3614094.

Lau:2006:BRG

[Lau06] Lap Chi Lau. Bipartite roots of graphs. ACM Transactions on Algorithms, 2(2):178–208, April 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Lokshtanov:2021:ACP

[LBSZ21] Daniel Lokshtanov, Andreas BjÖrklund, Saket Saurabh, and Meirav Zehavi. Approximate counting of k-paths: Simpler, deterministic, and in polynomial space. ACM Transactions on Algorithms, 17(3):26:1–26:44, August 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3461477.

Liu:2019:DFJ

[LCS^{+}19] Zhengyang Liu, Xi Chen, Rocco A. Servedio, Ying Sheng, and Jinyu Xie. Distribution-free junta testing. ACM Transactions on Algorithms, 15(1):1–23, January 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3264434.

Li:2009:PTA

[LDX09] Guojun Li, Xiaotie Deng, and Ying Xu. A polynomial-time approximation scheme for embedding hypergraph in a cycle. ACM Transactions on Algorithms, 5(2):20:1–20:??, March 2009. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Levin:2009:GMC

[Lev09] Asaf Levin. A generalized minimum cost k-clustering. ACM Transactions on Algorithms, 5(4):36:1–36:??, October 2009. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Li:2017:UCM

[Li17] Shi Li. On uniform capacitated k-median beyond the natural LP relaxation. ACM Transactions on Algorithms, 13(2):22:1–22:??, May 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Lemire:2008:HBO

[LK08] Daniel Lemire and Owen Kaser. Hierarchical bin buffering: Online local moments for dynamic external memory arrays. ACM Transactions on Algorithms, 4(1):14:1–14:??, March 2008. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Li:2025:PIG

[LL25] Shi Li and Bundit Laekhanukit. Polynomial integrality gap of flow LP for directed Steiner tree. ACM Transactions on Algorithms, 21(1):2:1–2:??, January 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Laber:2011:IAH

[LM11] Eduardo Laber and Marco Molinaro. Improved approximations for the hotlink assignment problem. ACM Transactions on Algorithms, 7(3):39:1–39:??, July 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Lokshtanov:2019:CIS

[LM19] Daniel Lokshtanov and Amer E. Mouawad. The complexity of independent set reconfiguration on bipartite graphs. ACM Transactions on Algorithms, 15(1):1–19, January 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3280825.

Lokshtanov:2021:AFV

[LMM^{+}21] Daniel Lokshtanov, Pranabendu Misra, Joydeep Mukherjee, Fahad Panolan, Geevarghese Philip, and Saket Saurabh. 2approximating feedback vertex set in tournaments. ACM Transactions on Algorithms, 17(2):11:1–11:14, June 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/10.1145/3446969.

Lubbecke:2016:NAO

[LMMW16] Elisabeth Lübbecke, Olaf Maurer, Nicole Megow, and Andreas Wiese. A new approach to online scheduling: Approximating the optimal competitive ratio. ACM Transactions on Algorithms, 13 (1):15:1–15:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Lokshtanov:2018:DTL

[LMPS18] Daniel Lokshtanov, Pranabendu Misra, Fahad Panolan, and Saket Saurabh. Deterministic truncation of linear matroids. ACM Transactions on Algorithms, 14(2):14:1–14:??, June 2018. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Lokshtanov:2018:KAG

[LMS18] Daniel Lokshtanov, Dániel Marx, and Saket Saurabh. Known algorithms on graphs of bounded treewidth are probably optimal. ACM Transactions on Algorithms, 14(2):13:1–13:??, June 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Li:2022:DFV

[LN22] Jason Li and Jesper Nederlof. Detecting feedback vertex sets of size k in O^{⋆}(2.7k) time. ACM Transactions on Algorithms, 18 (4):34:1–34:??, October 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3504027.

Lokshtanov:2014:FPA

[LNR^{+}14] Daniel Lokshtanov, N. S. Narayanaswamy, Venkatesh Raman, M. S. Ramanujan, and Saket Saurabh. Faster parameterized algorithms using linear programming. ACM Transactions on Algorithms, 11(2):15:1–15:??, October 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Lopez-Ortiz:2006:F

[LOM06] Alejandro López-Ortiz and J. Ian Munro. Foreword. ACM Transactions on Algorithms, 2(4):491, October 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Lokshtanov:2020:CSI

[LPS^{+}20] Daniel Lokshtanov, Fahad Panolan, Saket Saurabh, Roohani Sharma, and Meirav Zehavi. Covering small independent sets and separators with applications to parameterized algorithms. ACM Transactions on Algorithms, 16(3):32:1–32:31, June 2020. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3379698.

Li:2024:CEP

[LPS24] Shaohua Li, Marcin Pilipczuk, and Manuel Sorge. Cluster editing parameterized above modification-disjoint P_{3}-packings. ACM Transactions on Algorithms, 20(1):3:1–3:??, January 2024. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3626526.

Lokshtanov:2018:IED

[LPV18] Daniel Lokshtanov, Marcin Pilipczuk, and Erik Jan Van Leeuwen. Independence and efficient domination on P_{6}-free graphs. ACM Transactions on Algorithms, 14(1):3:1–3:??, January 2018. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Lee:2020:ISI

[LPW20] Yin Tat Lee, Marcin Pilipczuk, and David Woodruff. Introduction to the special issue on SODA’18. ACM Transactions on Algorithms, 16(1):1:1–1:2, January 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3368307.

Levi:2015:QPT

[LR15] Reut Levi and Dana Ron. A quasi-polynomial time partition oracle for graphs with an excluded minor. ACM Transactions on Algorithms, 11(3):24:1–24:??, January 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Lokshtanov:2018:LTP

[LRS18] Daniel Lokshtanov, M. S. Ramanujan, and Saket Saurabh. Linear time parameterized algorithms for subset feedback vertex set. ACM Transactions on Algorithms, 14(1):7:1–7:??, January 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Lee:2020:MMO

[LS20] Euiwoong Lee and Sahil Singla. Maximum matching in the online batch-arrival model. ACM Transactions on Algorithms, 16 (4):49:1–49:31, September 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3399676.

Los:2024:IDT

[LS24] Dimitrios Los and Thomas Sauerwald. An improved drift theorem for balanced allocations. ACM Transactions on Algorithms, 20 (4):35:1–35:??, October 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3673900.

Le:2024:GSE

[LT24] Hung Le and Cuong Than. Greedy spanners in Euclidean spaces admit sublinear separators. ACM Transactions on Algorithms, 20(3):23:1–23:??, July 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3590771.

Lu:2008:BPS

[LY08] Hsueh-I Lu and Chia-Chi Yeh. Balanced parentheses strike back. ACM Transactions on Algorithms, 4(3):28:1–28:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Leverrier:2025:EDC

[LZ25] Anthony Leverrier and Gilles Zémor. Efficient decoding up to a constant fraction of the code length for asymptotically good quantum codes. ACM Transactions on Algorithms, 21(4):35:1– 35:34, October 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Moroz:2016:CDB

[MA16] Guillaume Moroz and Boris Aronov. Computing the distance between piecewise-linear bivariate functions. ACM Transactions on Algorithms, 12(1):3:1–3:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Manthey:2012:AMT

[Man12] Bodo Manthey. On approximating multicriteria TSP. ACM Transactions on Algorithms, 8(2):17:1–17:??, April 2012. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Marx:2010:AFH

[Mar10] Dániel Marx. Approximating fractional hypertree width. ACM Transactions on Algorithms, 6(2):29:1–29:??, March 2010. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Mastrolilli:2021:CIM

[Mas21] Monaldo Mastrolilli. The complexity of the ideal membership problem for constrained problems over the Boolean domain. ACM Transactions on Algorithms, 17(4):32:1–32:29, October 2021. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3449350.

Mathieu:2010:FSI

[Mat10] Claire Mathieu. Foreword to special issue SODA 2009. ACM Transactions on Algorithms, 6(2):23:1–23:??, March 2010. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Mestre:2014:WPM

[Mes14] Julián Mestre. Weighted popular matchings. ACM Transactions on Algorithms, 10(1):2:1–2:??, January 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Mehlhorn:2009:MCB

[MM09] Kurt Mehlhorn and Dimitrios Michail. Minimum cycle bases: Faster and simpler. ACM Transactions on Algorithms, 6(1):8:1– 8:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Makarychev:2014:MQA

[MMS14] Konstantin Makarychev, Rajsekar Manokaran, and Maxim Sviridenko. Maximum quadratic assignment problem: Reduction from maximum label cover and LP-based approximation algorithm. ACM Transactions on Algorithms, 10(4):18:1–18:??, August 2014. ISSN 1549-6325 (print), 1549-6333 (electronic).

Makinen:2008:DEC

[MN08] Veli Mäkinen and Gonzalo Navarro. Dynamic entropy-compressed sequences and full-text indexes. ACM Transactions on Algorithms, 4(3):32:1–32:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

MUTZE:2018:ECM

[MN18] Torsten MÜTZE and Jerri Nummenpalo. Efficient computation of middle levels Gray codes. ACM Transactions on Algorithms, 14 (2):15:1–15:??, June 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Mahdian:2012:OOU

[MNS12] Mohammad Mahdian, Hamid Nazerzadeh, and Amin Saberi. Online Optimization with Uncertain Information. ACM Transactions on Algorithms, 8(1):2:1–2:??, January 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Marx:2013:FSS

[MOR13] Dáaniel Marx, Barry O’Sullivan, and Igor Razgon. Finding small separators in linear time via treewidth reduction. ACM Transactions on Algorithms, 9(4):30:1–30:??, September 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Marx:2022:OPA

[MP22] Dániel Marx and Michał Pilipczuk. Optimal parameterized algorithms for planar facility location problems using Voronoi diagrams. ACM Transactions on Algorithms, 18(2):13:1–13:64, April 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3483425.

Martins:2020:PIF

[MPQS20] Rodrigo S. V. Martins, Daniel Panario, Claudio Qureshi, and Eric Schmutz. Periods of iterations of functions with restricted preimage sizes. ACM Transactions on Algorithms, 16(3):30:1– 30:28, June 2020. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3378570.

Martinez:2010:ASS

[MPV10] Conrado Martínez, Daniel Panario, and Alfredo Viola. Adaptive sampling strategies for quickselects. ACM Transactions on Algorithms, 6(3):53:1–53:??, June 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Miller:2022:DLE

[MPY22] Avery Miller, Andrzej Pelc, and Ram Narayan Yadav. Deterministic leader election in anonymous radio networks. ACM Transactions on Algorithms, 18(3):25:1–25:??, July 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/10.1145/3527171.

Marko:2009:ADP

[MR09] Sharon Marko and Dana Ron. Approximating the distance to properties in bounded-degree and general sparse graphs. ACM Transactions on Algorithms, 5(2):22:1–22:??, March 2009. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Moore:2006:GQF

[MRR06] Cristopher Moore, Daniel Rockmore, and Alexander Russell. Generic quantum Fourier transforms. ACM Transactions on Algorithms, 2(4):707–723, October 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Mathieu:2024:CDS

[MRYY24] Claire Mathieu, Rajmohan Rajaraman, Neal E. Young, and Arman Yousefi. Competitive data-structure dynamization. ACM Transactions on Algorithms, 20(4):37:1–37:??, October 2024. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3672614.

Makarychev:2017:MPS

[MS17] Konstantin Makarychev and Maxim Sviridenko. Maximizing polynomials subject to assignment constraints. ACM Transactions on Algorithms, 13(4):54:1–54:??, December 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Moran:2011:PCR

[MSS11] Shlomo Moran, Sagi Snir, and Wing-Kin Sung. Partial convex recolorings of trees and galled networks: Tight upper and lower bounds. ACM Transactions on Algorithms, 7(4):42:1–42:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Makinen:2019:SDP

[MTK^{+}19] Veli Mäkinen, Alexandru I. Tomescu, Anna Kuosmanen, Topi Paavilainen, Travis Gagie, and Rayan Chikhi. Sparse dynamic programming on DAGs with small width. ACM Transactions on Algorithms, 15(2):1–21, May 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3301312.

Mendelson:2006:MPQ

[MTTZ06] Ran Mendelson, Robert E. Tarjan, Mikkel Thorup, and Uri Zwick. Melding priority queues. ACM Transactions on Algorithms, 2(4):535–556, October 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Madani:2010:DDM

[MTZ10] Omid Madani, Mikkel Thorup, and Uri Zwick. Discounted deterministic Markov decision processes and discounted all-pairs shortest paths. ACM Transactions on Algorithms, 6(2):33:1–33:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Massberg:2008:AAF

[MV08] Jens Maßberg and Jens Vygen. Approximation algorithms for a facility location problem with service capacities. ACM Transactions on Algorithms, 4(4):50:1–50:15, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Marx:2015:FPA

[MV15] Dániel Marx and László A. Végh. Fixed-parameter algorithms for minimum-cost edge-connectivity augmentation. ACM Transactions on Algorithms, 11(4):27:1–27:??, June 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Merino:2024:TDK

[MW24] Arturo Merino and Andreas Wiese. On the two-dimensional knapsack problem for convex polygons. ACM Transactions on Algorithms, 20(2):16:1–16:??, April 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/

doi/10.1145/3644390.

Mezei:2023:PSG

[MWsZ23] Balázs F. Mezei, Marcin Wrochna, and stanislav Zivný. PTAS for sparse general-valued CSPs. ACM Transactions on Algorithms, 19(2):14:1–14:??, April 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3569956.

Marx:2019:ISI

[MWY19] Dániel Marx, Virgi Vassilevska Williams, and Neal E. Young. Introduction to the special issue on SODA 2017. ACM Transactions on Algorithms, 15(2):1–2, May 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3319426.

Meyer:2012:ESP

[MZ12] Ulrich Meyer and Norbert Zeh. I/O-efficient shortest path algorithms for undirected graphs with random or bounded edge lengths. ACM Transactions on Algorithms, 8(3):22:1–22:??, July 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Mathieu:2023:PCV

[MZ23] Claire Mathieu and Hang Zhou. A PTAS for capacitated vehicle routing on trees. ACM Transactions on Algorithms, 19 (2):17:1–17:??, April 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3575799.

Navarro:2025:CMR

[Nav25] Gonzalo Navarro. Computing MEMs and relatives on repetitive text collections. ACM Transactions on Algorithms, 21(1):12:1– 12:??, January 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Narayanaswamy:2025:PRS

[ND25] N. S. Narayanaswamy and S. M. Dhannya. Perfect resolution of strong conflict-free colouring of interval hypergraphs. ACM Transactions on Algorithms, 21(1):10:1–10:??, January 2025. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Naor:2019:BFA

[NE19] Moni Naor and Yogev Eylon. Bloom filters in adversarial environments. ACM Transactions on Algorithms, 15(3):1–30, July 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3306193.

Neuen:2022:HIG

[Neu22] Daniel Neuen. Hypergraph isomorphism for groups with restricted composition factors. ACM Transactions on Algorithms, 18(3):27:1–27:??, July 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3527667.

Neuen:2024:ITG

[Neu24] Daniel Neuen. Isomorphism testing for graphs excluding small topological subgraphs. ACM Transactions on Algorithms, 20 (3):25:1–25:??, July 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3651986.

Nieberg:2008:ASW

[NHK08] Tim Nieberg, Johann Hurink, and Walter Kern. Approximation schemes for wireless networks. ACM Transactions on Algorithms, 4(4):49:1–49:??, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Newman:2018:ASS

[NRS18] Alantha Newman, Heiko Röglin, and Johanna Seif. The alternating stock size problem and the gasoline puzzle. ACM Transactions on Algorithms, 14(2):19:1–19:??, June 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Nguyen:2009:LDL

[NS09] Phong Q. Nguyen and Damien Stehlé. Low-dimensional lattice basis reduction revisited. ACM Transactions on Algorithms, 5(4): 46:1–46:??, October 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Naor:2010:DCA

[NS10] Joseph (Seffi) Naor and Roy Schwartz. The directed circular arrangement problem. ACM Transactions on Algorithms, 6(3): 47:1–47:??, June 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Navarro:2014:FFS

[NS14] Gonzalo Navarro and Kunihiko Sadakane. Fully functional static and dynamic succinct trees. ACM Transactions on Algorithms, 10 (3):16:1–16:??, June 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Neiman:2016:SDA

[NS16] Ofer Neiman and Shay Solomon. Simple deterministic algorithms for fully dynamic maximal matching. ACM Transactions on Algorithms, 12(1):7:1–7:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Nutov:2009:ACA

[Nut09] Zeev Nutov. Approximating connectivity augmentation problems. ACM Transactions on Algorithms, 6(1):5:1–5:19, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Nutov:2012:AMC

[Nut12] Zeev Nutov. Approximating minimum-cost connectivity problems via uncrossable bifamilies. ACM Transactions on Algorithms, 9 (1):1:1–1:??, December 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Naor:2007:NAP

[NW07] Moni Naor and Udi Wieder. Novel architectures for P2P applications: The continuous-discrete approach. ACM Transactions on Algorithms, 3(3):34:1–34:??, August 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Nakajima:2024:CSV

[NZ24] Tamio-Vesa Nakajima and Stanislav Zivný. On the complexity of symmetric vs. functional PCSPs. ACM Transactions on Algorithms, 20(4):33:1–33:??, October 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3673655.

Nong:2011:CIS

[NZC11] Ge Nong, Sen Zhang, and Wai Hong Chan. Computing the Inverse Sort Transform in linear time. ACM Transactions on Algorithms, 7(2):27:1–27:??, March 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Oudot:2010:GDT

[OGGW10] Steve Y. Oudot, Leonidas J. Guibas, Jie Gao, and Yue Wang. Geodesic Delaunay triangulations in bounded planar domains. ACM Transactions on Algorithms, 6(4):67:1–67:??, August 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Olbrich:2024:GNR

[OOB24] Jannik Olbrich, Enno Ohlebusch, and Thomas Büchler. Generic non-recursive suffix array construction. ACM Transactions on Algorithms, 20(2):18:1–18:??, April 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3641854.

Onak:2020:FDM

[OSSW20] Krzysztof Onak, Baruch Schieber, Shay Solomon, and Nicole Wein. Fully dynamic MIS in uniformly sparse graphs. ACM Transactions on Algorithms, 16(2):26:1–26:19, April 2020. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3378025.

Oum:2008:ARW

[Oum08] Sang-Il Oum. Approximating rank-width and clique-width quickly. ACM Transactions on Algorithms, 5(1):10:1–10:??, November 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Pagh:2018:CLS

[Pag18] Rasmus Pagh. CoveringLSH: Locality-sensitive hashing without false negatives. ACM Transactions on Algorithms, 14(3):29:1– 29:??, July 2018. CODEN ???? ISSN 1549-6325 (print), 1549-

6333 (electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3155300.

Pettie:2009:LDS

[Pet09] Seth Pettie. Low distortion spanners. ACM Transactions on Algorithms, 6(1):7:1–7:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Parter:2016:SFT

[PP16] Merav Parter and David Peleg. Sparse fault-tolerant BFS structures. ACM Transactions on Algorithms, 13(1):11:1–11:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Parter:2018:FTA

[PP18] Merav Parter and David Peleg. Fault-tolerant approximate BFS structures. ACM Transactions on Algorithms, 14(1):10:1–10:??, January 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Pilipczuk:2018:NSS

[PPSV18] Marcin Pilipczuk, Michał Pilipczuk, Piotr Sankowski, and Erik Jan Van Leeuwen. Network sparsification for Steiner problems on planar and bounded-genus graphs. ACM Transactions on Algorithms, 14(4):1–73, October 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3239560.

Pettie:2008:RMS

[PR08] Seth Pettie and Vijaya Ramachandran. Randomized minimum spanning tree algorithms using exponentially fewer random bits. ACM Transactions on Algorithms, 4(1):5:1–5:??, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Panconesi:2012:EPS

[PR12] Alessandro Panconesi and Jaikumar Radhakrishnan. Expansion properties of (secure) wireless networks. ACM Transactions on Algorithms, 8(3):21:1–21:??, July 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Prezza:2021:OSE

[Pre21] Nicola Prezza. Optimal substring equality queries with applications to sparse text indexing. ACM Transactions on Algorithms, 17(1):7:1–7:23, January 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3426870.

Philip:2012:PKD

[PRS12] Geevarghese Philip, Venkatesh Raman, and Somnath Sikdar. Polynomial kernels for dominating set in graphs of bounded degeneracy and beyond. ACM Transactions on Algorithms, 9(1): 11:1–11:??, December 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Pandurangan:2020:TMO

[PRS20] Gopal Pandurangan, Peter Robinson, and Michele Scquizzato. A time- and message-optimal distributed algorithm for minimum spanning trees. ACM Transactions on Algorithms, 16(1):13:1– 13:27, January 2020. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3365005.

Pemmaraju:2011:MCO

[PRV11] Sriram V. Pemmaraju, Rajiv Raman, and Kasturi Varadarajan. Max-coloring and online coloring with bandwidths on interval graphs. ACM Transactions on Algorithms, 7(3):35:1–35:??, July 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Panagiotou:2010:MBS

[PS10] Konstantinos Panagiotou and Angelika Steger. Maximal biconnected subgraphs of random planar graphs. ACM Transactions on Algorithms, 6(2):31:1–31:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Polacek:2016:QPL

[PS16] Lukáš Poláček and Ola Svensson. Quasi-polynomial local search for restricted max-min fair allocation. ACM Transactions on Algorithms, 12(2):13:1–13:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Paz:2019:AMW

[PS19] Ami Paz and Gregory Schwartzman. A (2 + ϵ)-approximation for maximum weight matching in the semi-streaming model. ACM Transactions on Algorithms, 15(2):1–15, May 2019. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3274668.

Peserico:2023:MLA

[PS23] Enoch Peserico and Michele Scquizzato. Matching on the line admits no o(\sqrt{log n})-competitive algorithm. ACM Transactions on Algorithms, 19(3):28:1–28:??, July 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3594873.

Panolan:2024:CDU

[PSZ24] Fahad Panolan, Saket Saurabh, and Meirav Zehavi. Contraction decomposition in unit disk graphs and algorithmic applications in parameterized complexity. ACM Transactions on Algorithms, 20(2):15:1–15:??, April 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3648594.

Pritchard:2011:FCS

[PT11] David Pritchard and Ramakrishna Thurimella. Fast computation of small cuts via cycle space sampling. ACM Transactions on Algorithms, 7(4):46:1–46:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Patrascu:2016:IRL

[PT16] Mihai Pătraşcu and Mikkel Thorup. On the k-independence required by linear probing and minwise independence. ACM Transactions on Algorithms, 12(1):8:1–8:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Pandurangan:2007:EBB

[PU07] Gopal Pandurangan and Eli Upfal. Entropy-based bounds for online algorithms. ACM Transactions on Algorithms, 3(1):??, February 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Pruhs:2008:GBR

[PUW08] Kirk Pruhs, Patchrawat Uthaisombut, and Gerhard Woeginger. Getting the best response for your erg. ACM Transactions on Algorithms, 4(3):38:1–38:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Pelleg:2024:ASC

[PZ24] Eden Pelleg and Stanislav Zivný. Additive sparsification of CSPs. ACM Transactions on Algorithms, 20(1):1:1–1:??, January 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3625824.

Rottenstreich:2020:CHM

[RKH20] Ori Rottenstreich, Haim Kaplan, and Avinatan Hassidim. Clustering in hypergraphs to minimize average edge service time. ACM Transactions on Algorithms, 16(3):40:1–40:28, June 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3386121.

Russo:2011:FCS

[RNO11] Luís M. S. Russo, Gonzalo Navarro, and Arlindo L. Oliveira. Fully compressed suffix trees. ACM Transactions on Algorithms, 7(4):53:1–53:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Roditty:2008:FSF

[Rod08] Liam Roditty. A faster and simpler fully dynamic transitive closure. ACM Transactions on Algorithms, 4(1):6:1–6:??, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Raman:2007:SID

[RRS07] Rajeev Raman, Venkatesh Raman, and Srinivasa Rao Satti. Succinct indexable dictionaries with applications to encoding k-ary trees, prefix sums and multisets. ACM Transactions on Algorithms, 3(4):43:1–43:??, November 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Rabani:2016:ESI

[RRSW16] Yuval Rabani, Andrea Richa, Jared Saia, and David P. Woodruff. Editorial to the special issue on SODA’12. ACM Transactions on Algorithms, 12(1):1:1–1:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Rabani:2009:BAT

[RS09] Yuval Rabani and Gabriel Scalosub. Bicriteria approximation tradeoff for the node-cost budget problem. ACM Transactions on Algorithms, 5(2):19:1–19:??, March 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Roditty:2011:APS

[RS11a] Liam Roditty and Asaf Shapira. All-pairs shortest paths with a sublinear additive error. ACM Transactions on Algorithms, 7 (4):45:1–45:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Rosen:2011:RVB

[RS11b] Adi Rosén and Gabriel Scalosub. Rate vs. buffer size — greedy information gathering on the line. ACM Transactions on Algorithms, 7(3):32:1–32:??, July 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ramanujan:2017:LTP

[RS17] M. S. Ramanujan and Saket Saurabh. Linear-time parameterized algorithms via skew-symmetric multicuts. ACM Transactions on Algorithms, 13(4):46:1–46:??, December 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Raman:2006:FFP

[RSS06] Venkatesh Raman, Saket Saurabh, and C. R. Subramanian. Faster fixed parameter tractable algorithms for finding feedback vertex sets. ACM Transactions on Algorithms, 2(3):403–415, July 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Rue:2014:DPG

[RST14] Juanjo Rué, Ignasi Sau, and Dimitrios M. Thilikos. Dynamic programming for graphs on surfaces. ACM Transactions on Algorithms, 10(2):8:1–8:??, February 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Roditty:2013:AG

[RT13] Liam Roditty and Roei Tov. Approximating the girth. ACM Transactions on Algorithms, 9(2):15:1–15:??, March 2013. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ron:2014:TPS

[RT14] Dana Ron and Gilad Tsur. Testing properties of sparse images. ACM Transactions on Algorithms, 10(4):17:1–17:??, August 2014. ISSN 1549-6325 (print), 1549-6333 (electronic).

Rahul:2022:GTB

[RT22] Saladi Rahul and Yufei Tao. Generic techniques for building topk structures. ACM Transactions on Algorithms, 18(4):38:1–38:??, October 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3546074.

Roditty:2008:RSR

[RTZ08] Iam Roditty, Mikkel Thorup, and Uri Zwick. Roundtrip spanners and roundtrip routing in directed graphs. ACM Transactions on Algorithms, 4(3):29:1–29:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ruzic:2008:UDD

[Ruž08] Milan Ružić. Uniform deterministic dictionaries. ACM Transactions on Algorithms, 4(1):1:1–1:??, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ruzic:2009:MDS

[Ruž09] Milan Ružić. Making deterministic signatures quickly. ACM Transactions on Algorithms, 5(3):26:1–26:??, July 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Reed:2009:LTA

[RW09] Bruce Reed and David R. Wood. A linear-time algorithm to find a separator in a graph excluding a minor. ACM Transactions on Algorithms, 5(4):39:1–39:??, October 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ruskey:2010:EUC

[RW10a] Frank Ruskey and Aaron Williams. An explicit universal cycle for the (n − 1)-permutations of an n-set. ACM Transactions on Algorithms, 6(3):45:1–45:12, June 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Rutter:2010:CLM

[RW10b] Ignaz Rutter and Alexander Wolff. Computing large matchings fast. ACM Transactions on Algorithms, 7(1):1:1–1:??, November 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Roditty:2012:RPS

[RZ12] Liam Roditty and Uri Zwick. Replacement paths and k simple shortest paths in unweighted directed graphs. ACM Transactions on Algorithms, 8(4):33:1–33:??, September 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Sawada:2006:GRF

[Saw06] Joe Sawada. Generating rooted and free plane trees. ACM Transactions on Algorithms, 2(1):1–13, January 2006. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Shi:2017:DPD

[SCRS17] Elaine Shi, T.-H. Hubert Chan, Eleanor Rieffel, and Dawn Song. Distributed private data analysis: Lower bounds and practical constructions. ACM Transactions on Algorithms, 13(4):50:1– 50:??, December 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Shallue:2016:TPT

[Sha16] Andrew Shallue. Tabulating pseudoprimes and tabulating liars. ACM Transactions on Algorithms, 13(1):4:1–4:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Segal-Halevi:2016:WMH

[SHHA16] Erel Segal-Halevi, Avinatan Hassidim, and Yonatan Aumann. Waste makes haste: Bounded time algorithms for envy-free cake cutting with free disposal. ACM Transactions on Algorithms, 13 (1):12:1–12:??, December 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Socala:2016:TLB

[Soc16] Arkadiusz Socala. Tight lower bound for the channel assignment problem. ACM Transactions on Algorithms, 12(4):48:1–48:??, September 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Solomon:2013:SES

[Sol13] Shay Solomon. Sparse Euclidean spanners with tiny diameter. ACM Transactions on Algorithms, 9(3):28:1–28:??, June 2013. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

S:2024:DRP

[SP24] Karthik C. S. and Merav Parter. Deterministic replacement path covering. ACM Transactions on Algorithms, 20(4):34:1–34:??, October 2024. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3673760.

Srinivasan:2019:E

[Sri19] Aravind Srinivasan. Editorial. ACM Transactions on Algorithms, 15(2):1, May 2019. CODEN ???? ISSN 1549-6325 (print),

1549-6333 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3325824.

Sanders:2008:AAS

[SS08a] Peter Sanders and David Steurer. An asymptotic approximation scheme for multigraph edge coloring. ACM Transactions on Algorithms, 4(2):21:1–21:??, May 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Swamy:2008:FTF

[SS08b] Chaitanya Swamy and David B. Shmoys. Fault-tolerant facility location. ACM Transactions on Algorithms, 4(4):51:1–51:??, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Scott:2009:PCS

[SS09] Alexander D. Scott and Gregory B. Sorkin. Polynomial constraint satisfaction problems, graph bisection, and the Ising partition function. ACM Transactions on Algorithms, 5(4):45:1–45:??, October 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Schmid:2018:CWP

[SS18] Andreas Schmid and Jens M. Schmidt. Computing 2-walks in polynomial time. ACM Transactions on Algorithms, 14(2):22:1– 22:??, June 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Sala:2025:ECP

[SSA25] Evan Sala, Joe Sawada, and Abbas Alhakim. Efficient constructions of the prefer-same and prefer-opposite de Bruijn sequences. ACM Transactions on Algorithms, 21(1):1:1–1:??, January 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Salvy:2011:PFF

[SSS^{+}11] Bruno Salvy, Bob Sedgewick, Michele Soria, Wojciech Szpankowski, and Brigitte Vallee. Philippe Flajolet, the father of analytic combinatorics. ACM Transactions on Algorithms, 7 (4):40:1–40:??, September 2011. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Sau:2022:KAM

[SST22] Ignasi Sau, Giannos Stamoulis, and Dimitrios M. Thilikos. kapices of minor-closed graph classes. II. Parameterized algorithms. ACM Transactions on Algorithms, 18(3):21:1–21:??, July 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3519028.

Shah:2023:RDR

[SSTV23] Rahul Shah, Cheng Sheng, Sharma Thankachan, and Jeffrey Vitter. Ranked document retrieval in external memory. ACM Transactions on Algorithms, 19(1):5:1–5:??, January 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https:

//dl.acm.org/doi/10.1145/3559763.

Suzuki:2008:DSP

[ST08] Akiko Suzuki and Takeshi Tokuyama. Dense subgraph problems with output-density conditions. ACM Transactions on Algorithms, 4(4):43:1–43:??, August 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Svitkina:2010:FLH

[ST10] Zoya Svitkina and Éva Tardos. Facility location with hierarchical facility costs. ACM Transactions on Algorithms, 6(2):37:1–37:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Sinnamon:2025:ESA

[ST25] Corwin Sinnamon and Robert E. Tarjan. Efficiency of selfadjusting heaps. ACM Transactions on Algorithms, 21(4):37:1– 37:39, October 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Sen:2016:DRB

[STK16] Siddhartha Sen, Robert E. Tarjan, and David Hong Kyun Kim. Deletion without rebalancing in binary search trees. ACM Transactions on Algorithms, 12(4):57:1–57:??, September 2016. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Svitkina:2010:LBF

[Svi10] Zoya Svitkina. Lower-bounded facility location. ACM Transactions on Algorithms, 6(4):69:1–69:??, August 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Sawada:2020:SST

[SW20a] Joe Sawada and Aaron Williams. Solving the sigma–tau problem. ACM Transactions on Algorithms, 16(1):11:1–11:17, January 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3359589.

Solomon:2020:IDG

[SW20b] Shay Solomon and Nicole Wein. Improved dynamic graph coloring. ACM Transactions on Algorithms, 16(3):41:1–41:24, June 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3392724.

Swamy:2012:ESS

[Swa12] Chaitanya Swamy. The effectiveness of Stackelberg strategies and tolls for network congestion games. ACM Transactions on Algorithms, 8(4):36:1–36:??, September 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Swamy:2016:IAA

[Swa16] Chaitanya Swamy. Improved approximation algorithms for matroid and knapsack median problems and applications. ACM Transactions on Algorithms, 12(4):49:1–49:??, September 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Sun:2021:QMT

[SWYZ21] Xiaoming Sun, David P. Woodruff, Guang Yang, and Jialin Zhang. Querying a matrix through matrix–vector products. ACM Transactions on Algorithms, 17(4):31:1–31:19, October 2021. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3470566.

Soltan:2020:DBC

[SYZ20] Saleh Soltan, Mihalis Yannakakis, and Gil Zussman. Doubly balanced connected graph partitioning. ACM Transactions on Algorithms, 16(2):20:1–20:24, April 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3381419.

Shalita:2010:EAG

[SZ10] Alon Shalita and Uri Zwick. Efficient algorithms for the 2gathering problem. ACM Transactions on Algorithms, 6(2):34:1– 34:??, March 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Stein:2020:SWY

[SZ20] Clifford Stein and Mingxian Zhong. Scheduling when you do not know the number of machines. ACM Transactions on Algorithms, 16(1):9:1–9:20, January 2020. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3340320.

Tao:2014:DRS

[Tao14] Yufei Tao. Dynamic ray stabbing. ACM Transactions on Algorithms, 11(2):11:1–11:??, October 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Terao:2025:FMP

[Ter25] Tatsuya Terao. Faster matroid partition algorithms. ACM Transactions on Algorithms, 21(2):14:1–14:??, April 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Thomasse:2010:KFV

[Tho10a] Stéphan Thomassé. A 4k^{2} kernel for feedback vertex set. ACM Transactions on Algorithms, 6(2):32:1–32:??, March 2010. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Thomasse:2010:KKF

[Tho10b] Stéphan Thomassé. A 4k^{2} kernel for feedback vertex set. ACM Transactions on Algorithms, 6(2):32:1–32:??, March 2010. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Tarjan:2021:ZT

[TLT21] Robert E. Tarjan, Caleb Levy, and Stephen Timmel. Zip trees. ACM Transactions on Algorithms, 17(4):34:1–34:12, October 2021. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3476830.

Torng:2008:SOU

[TM08] Eric Torng and Jason McCullough. SRPT optimally utilizes faster machines to minimize flow time. ACM Transactions on Algorithms, 5(1):1:1–1:??, November 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ta-Shma:2014:DRT

[TSZ14] Amnon Ta-Shma and Uri Zwick. Deterministic rendezvous, treasure hunts, and strongly universal exploration sequences. ACM Transactions on Algorithms, 10(3):12:1–12:??, June 2014. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Tamaki:2018:AGM

[TY18] Suguru Tamaki and Yuichi Yoshida. Approximation guarantees for the minimum linear arrangement problem by higher eigenvalues. ACM Transactions on Algorithms, 14(4):1–13, October 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3228342.

VeghVegh:2008:PDA

[VB08] László A. VéghVégh and András A. Benczúr. Primal-dual approach for directed vertex connectivity augmentation and generalizations. ACM Transactions on Algorithms, 4(2):20:1–20:??, May 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Vinkemeier:2005:LTA

[VH05] Doratha E. Drake Vinkemeier and Stefan Hougardy. A linear-time approximation algorithm for weighted matchings in graphs. ACM Transactions on Algorithms, 1(1):107–122, July 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Vigneron:2014:GOS

[Vig14] Antoine Vigneron. Geometric optimization and sums of algebraic functions. ACM Transactions on Algorithms, 10(1):4:1–4:??, January 2014. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Viola:2005:EDI

[Vio05] Alfredo Viola. Exact distribution of individual displacements in linear probing hashing. ACM Transactions on Algorithms, 1(2): 214–242, October 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Vishwanathan:2008:HIA

[Vis08] Sundar Vishwanathan. On hard instances of approximate vertex cover. ACM Transactions on Algorithms, 5(1):7:1–7:??, November 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Voronenko:2007:MMC

[VP07] Yevgen Voronenko and Markus Püschel. Multiplierless multiple constant multiplication. ACM Transactions on Algorithms, 3(2): 11:1–11:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Vassilevska:2010:FHS

[VWY10] Virginia Vassilevska, Ryan Williams, and Raphael Yuster. Finding heaviest H-subgraphs in real weighted graphs, with applications. ACM Transactions on Algorithms, 6(3):44:1–44:??, June 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Viola:2021:CBL

[VZ21] Caterina Viola and Stanislav Zivný. The combined basic LP and affine IP relaxation for promise VCSPs on infinite domains. ACM Transactions on Algorithms, 17(3):21:1–21:23, August 2021. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL

https://dl.acm.org/doi/10.1145/3458041.

Wahlstrom:2022:QMM

[Wah22] Magnus Wahlström. Quasipolynomial multicut-mimicking networks and kernels for multiway cut problems. ACM Transactions on Algorithms, 18(2):15:1–15:19, April 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.

org/doi/10.1145/3501304.

Walzer:2023:LTC

[Wal23] Stefan Walzer. Load thresholds for cuckoo hashing with overlapping blocks. ACM Transactions on Algorithms, 19(3):24:1–24:??, July 2023. CODEN ???? ISSN 1549-6325 (print), 1549-6333

(electronic). URL https://dl.acm.org/doi/10.1145/3589558.

Walzer:2025:PCO

[Wal25] Stefan Walzer. Peeling close to the orientability threshold spatial coupling in hashing-based data structures. ACM Transactions on Algorithms, 21(3):33:1–33:23, July 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Wang:2022:MFP

[Wan22] Yipu Wang. Max flows in planar graphs with vertex capacities. ACM Transactions on Algorithms, 18(1):9:1–9:27, January 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

URL https://dl.acm.org/doi/10.1145/3504032.

Wei:2022:OVA

[Wei22] Alexander Wei. Optimal Las Vegas approximate near neighbors in ℓ_{p}. ACM Transactions on Algorithms, 18(1):7:1–7:27, January 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3461777.

Williams:2010:NPW

[Wil10] Virginia Vassilevska Williams. Nondecreasing paths in a weighted graph or: How to optimally read a train schedule. ACM Transactions on Algorithms, 6(4):70:1–70:??, August 2010. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Wimmer:2016:ALP

[Wim16] Karl Wimmer. Agnostic learning in permutation-invariant domains. ACM Transactions on Algorithms, 12(4):46:1–46:??, September 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Wild:2015:ACD

[WNN15] Sebastian Wild, Markus E. Nebel, and Ralph Neininger. Average case and distributional analysis of dual-pivot Quicksort. ACM Transactions on Algorithms, 11(3):22:1–22:??, January 2015. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Wang:2022:TBO

[WW22] Ruosong Wang and David P. Woodruff. Tight bounds for ℓ_{1} oblivious subspace embeddings. ACM Transactions on Algorithms, 18(1):8:1–8:32, January 2022. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic). URL https://dl.acm.org/doi/

10.1145/3477537.

Weimann:2013:RPD

[WY13] Oren Weimann and Raphael Yuster. Replacement paths and distance sensitivity oracles via fast matrix multiplication. ACM Transactions on Algorithms, 9(2):14:1–14:??, March 2013. CO-DEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Weimann:2016:ADP

[WY16] Oren Weimann and Raphael Yuster. Approximating the diameter of planar graphs in near linear time. ACM Transactions on Algorithms, 12(1):12:1–12:??, February 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Wei:2018:TSB

[WY18] Zhewei Wei and Ke Yi. Tight space bounds for two-dimensional approximate range counting. ACM Transactions on Algorithms, 14(2):23:1–23:??, June 2018. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Woodruff:2025:OLW

[WY25] David P. Woodruff and Taisuke Yasuda. Online Lewis weight sampling. ACM Transactions on Algorithms, 21(4):39:1–39:50, October 2025. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ward:2016:MSF

[WZ16] Justin Ward and Stanislav Zivný. Maximizing k-submodular functions and beyond. ACM Transactions on Algorithms, 12 (4):47:1–47:??, September 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ye:2012:EG

[YB12] Yuli Ye and Allan Borodin. Elimination graphs. ACM Transactions on Algorithms, 8(2):14:1–14:??, April 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Yu:2008:IAM

[YLW08] Hung-I. Yu, Tzu-Chin Lin, and Biing-Feng Wang. Improved algorithms for the minmax-regret 1-center and 1-median problems. ACM Transactions on Algorithms, 4(3):36:1–36:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Yuster:2005:FSM

[YZ05] Raphael Yuster and Uri Zwick. Fast sparse matrix multiplication. ACM Transactions on Algorithms, 1(1):2–13, July 2005. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Yi:2012:MOT

[YZ12] Ke Yi and Qin Zhang. Multidimensional online tracking. ACM Transactions on Algorithms, 8(2):12:1–12:??, April 2012. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Zhang:2008:CCP

[ZO08] Cun-Quan Zhang and Yongbin Ou. Clustering, community partition and disjoint spanning trees. ACM Transactions on Algorithms, 4(3):35:1–35:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).